### v2

# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

Abstract
This paper constructs a full exceptional collection in the bounded derived category of coherent sheaves on a generalized flag scheme $G/P$ over $\mathbb{Z}$, by means of a $G$-linear semiorthogonal decomposition of the bounded derived category of noetherian representations of $P$. The ensuing semiorthogonal decompositions are compatible with the Bruhat order on cosets of the Weyl group of $P$ in the Weyl group of $G$. The construction builds upon the foundational results on $B$-modules from the works of Mathieu, Polo, and van der Kallen, and upon properties of the Steinberg basis of the $T$-equivariant ${\sf K}$-theory of $G/B$.
keywords:
flag variety, exceptional collection, $B$-module, equivariant $K$-theory
pacs:
MSC2020 Classification]14F08, 14L30, 14M15, 19L47, 20G05, 20G10
Acknowledgements
A.S. is grateful to Henning Krause for his unwavering support of this work and for his and Charles Vial’s joint invitation to the BIREP group of the University of Bielefeld for a long-term visit in 2023-2024. A.S. work was supported by the Deutsche Forschungsgemeinschaft (SFB-TRR 358/1 2023 - 491392403). He also acknowledges the hospitality of the Mathematical Institute of the University of Utrecht during his visit in May 2024.
W.v.d.K. thanks BIREP for the opportunity to visit Bielefeld in September 2023.
1 Introduction
1.1 The context
Let $\mathbb{G}$ be a split semi-simple simply connected Chevalley group scheme and $\mathbb{B}$ a split Borel subgroup scheme. Associated to it is the flag scheme $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$. This paper constructs full exceptional collections in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$. More generally, let $\mathbb{P}$ be a parabolic subgroup scheme containing $\mathbb{B}$. We also construct full exceptional collections in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{P}$ over $\mathbb{Z}$.
The problem of finding full exceptional collections on flag varieties and, more generally, on quotients of $\mathbb{G}$ modulo a parabolic subgroup ${\mathbb{P}}$ has a long history; we refer the reader to the paper [22] and the references therein. In this non-technical part of the introduction we emphasise some of the features that make the approach of the present paper different from many of the existing constructions.
Given a smooth proper Noetherian scheme over a field ${\sf k}$, a full exceptional collection in the bounded derived category of coherent sheaves $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ is, informally speaking, a way to break up $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ into elementary pieces, each equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{vect}\textrm{-}{\sf k}}\nolimits)$ which are glued to each other in a non-trivial way. Upon decategorification, a full exceptional collection gives rise to a basis of the Grothendieck group ${{\sf K}}^{0}(X)$. If one is interested only in the ${\sf K}$-theoretic information, for many schemes of interest it is easy to produce a basis of ${{\sf K}}^{0}(X)$: for instance, if a scheme $X$ has a ${\mathbb{G}}_{m}$-action with isolated fixed points then by Białynicki-Birula’s theorem the scheme $X$ has a stratification into locally closed subschemes such that the structure sheaves of their closures are a basis of ${{\sf K}}^{0}(X)$. Flag schemes $\mathbb{G}/\mathbb{B}$ furnish a classical example of such a stratification: in this case, the closures of algebraic cells of the Białynicki-Birula’s stratification are Schubert schemes via the Bruhat decomposition.
Given a scheme as above and assuming it has a natural geometric basis of the group ${{\sf K}}^{0}(X)$, one can ask whether such a basis can be lifted to a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. Speaking less loosely, a full exceptional collection must satisfy two conditions: it should generate the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ in a suitable sense and some cohomological vanishing should hold. Generation doesn’t pose a problem: for schemes as above, i.e. having an algebraic stratification into locally closed subschemes, which are isomorphic to affine spaces, the direct sum of structure sheaves of their closures is a classical generator of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. But there is no chance that for such a set of generators the cohomological vanishing (known as “semiorthogonality” condition) would hold. On the other hand, there is more to flag schemes that one would want to take into account: there is a partial order on Schubert schemes (the Bruhat order) and the order on the sought-for full exceptional collection is expected to be compatible with the Bruhat order on Schubert schemes. Thus, what one is after is finding a correct “lift” of the natural ${\sf K}$-theoretic basis to the derived category. In the present paper, we provide such a lift. As explained below, there are two major inputs into our construction: one is representation-theoretic and comes from the highest weight category structures on the category of rational representations of the Borel group scheme $\mathbb{B}$, and the other one is combinatorial and comes from the distinguished basis of the ($\mathbb{T}$-equivariant) ${\sf K}$-group of the flag scheme $\mathbb{G}/\mathbb{B}$, called the Steinberg basis.
There is a basic way to relate algebraic representations of $\mathbb{B}$ to $\mathbb{G}$-equivariant vector bundles on $\mathbb{G}/\mathbb{B}$ via the “associated sheaf” construction. Note that the latter construction was the starting point in [34] (see Section 1.3 below for more on that). The category ${\mathrm{rep}}(\mathbb{G})$ of rational modules over a reductive group scheme $\mathbb{G}$ has a highest weight category structure, [13], [17], and it is also such a structure that one expects to manifest itself through conjectural full exceptional collections. In this respect, a model example of such semiorthogonal decompositions/full exceptional collections is given by a theorem of Efimov’s [20] that solves the problem for the Grassmannians $\sf Gr_{k,n}$ in the best possible way. Precisely, one has (cf. also [8]):
Theorem 1.1 ([20, Theorem 1.8]).
Let $\sf Gr_{k,n}$ be the Grassmannian of $k$-dimensional vector subspaces of $n$-dimensional space, defined over $\mathbb{Z}$. There exists a tilting vector bundle on $\sf Gr_{k,n}$ such that its endomorphism algebra has two natural structures of a split quasi–hereditary algebra over $\mathbb{Z}$.
The above theorem 1.1 and, more generally, results of [20] and [8] rely considerably on representation theory of the full linear group $GL_{n}$ and on the highest weight category structure on ${\mathrm{rep}}(GL_{n})$. Remarkably, highest weight category structures have also been known to exist on the category ${\mathrm{rep}}(B)$ of rational $B$-modules for the Borel subgroup $B\subset G$ through the works [35], [41], [46] and [47]. It seems, however, that the category ${\mathrm{rep}}(B)$ as a highest weight category has not yet received the full attention that it undoubtedly deserves. The following quote from Stephen Donkin, [19, Remark 4.7] has served as a particular impetus for the first named author: “In fact ${\mathsf{k}}[{B}]$ is a quasi–hereditary coalgebra. This follows from van der Kallen’s paper [46]. This is a deep and sophisticated work which generalizes the above (restriction gives a full embedding of the category of $G$-modules into the category of $B$-modules) and whose consequences, to the best of my knowledge, have so far not been investigated or exploited.”
In the present paper, we demonstrate the force of highest weight category structures on the category ${\mathrm{Rep}}(\mathbb{B})$ and, in accordance with Donkin’s suggestion, develop one of its consequences constructing full exceptional collections on flag schemes $\mathbb{G}/\mathbb{B}$. In that context, the two main theorems are
Theorem 1.2 (Theorem 11.5).
The category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ has a $G$-linear semiorthogonal decomposition
$\mathcal{D}=\langle\sf X_{v}\rangle_{v\in W}$ (1)
with respect to the total order $\prec$ on the Weyl group $W$. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Theorem 1.3 (Theorem 12.1).
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Let $v,w$ denote elements of $W$. Then there are objects ${\mathcal{X}}_{v}\in\mathcal{D}$ such that
  1. 1.
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{v}[i])=\begin{cases}{\sf k}&\text{if $i=0$}\\ 0&\text{else.}\end{cases}$.
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{w}[i])=0$ for all $i$.
  3. 3.
    The triangulated hull of $\{{\mathcal{X}}_{v}\mid v\in W\}$ is $\mathcal{D}$.
For more details, see Section 11 and Section 12, respectively. And see section 13 for the treatment of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$, $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/P))$.
By its very design, highest weight category structure on an abelian category gives rise to two distinguished collections of objects (standard and costandard objects) that satisfy the ${\mathrm{Ext}}$-vanishing close to the one that is required for exceptional collections. It would therefore be reasonable to start off with one of such collections in ${\mathrm{Rep}}(\mathbb{B})$, convert them into a collection of equivariant vector bundles, and try to ensure the cohomological vanishing conditions in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits({\mathbb{G}}/{\mathbb{B}}))$. It turns out that this doesn’t work on the nose, but it does work with the input provided by the combinatorics of the Steinberg basis. Coming back to the problem of lifts of $\sf K$-theoretic classes to the derived category level, it is important to mention the classical works [5] and [16] that compute the singular cohomology and, respectively, the $\sf K$-theory of flag schemes $\mathbb{G}/\mathbb{B}$ in terms of natural operators called nowadays BGG-Demazure operators. These operators, in different guises, will appear throughout the paper: one of the highest weight category structures on ${\mathrm{rep}}(B)$ is defined in terms of those, see Section 2. We now proceed to the second part of the introduction in which we outline the main steps of the argument.
1.2 Outline of the proof
Evidence for the principal construction of this paper comes from semiorthogonal decompositions for flag varieties of rank two groups over ${\mathbb{Z}}[\frac{1}{6}]$ established in [42]. We provide more detail of the rank two case in Section 14 below.
For the purposes of introduction, assume for simplicity that $\mathbb{P}=\mathbb{B}$ and that we are working over a field ${\sf k}$ whose characteristic is not too small. [^1] We actually sometimes worked over ${\mathbb{C}}$, sometimes over ${\mathbb{Z}}$.; thus, $B$ and $G$ denote the base change of group schemes $\mathbb{B}$ and $\mathbb{G}$ along the morphism ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$, [18, Section 2.6]. Given a finite dimensional $B$-module $M$, let $\mathcal{L}(M)$ denote the associated (locally free) coherent sheaf on $G/B$. The examples in rank two have full exceptional collections consisting of locally free sheaves that are in fact obtained via the forgetful functor ${\mathrm{Rep}}(B)\mapsto{\mathop{\sf{Coh}}\nolimits}({G/B})$ and therefore have an underlying $B$-module.
With those rank two examples in hand, it has become apparent that the $B$-module theory developed in [47] would be vital in generalizing the approach to arbitrary rank. One way to summarize the results of loc.cit. is saying that the abelian category ${\mathrm{Rep}}(B)$ has two highest weight category structures; the costandard objects with respect to one structure are denoted $P(\lambda)$, $\lambda\in X(T)$ (where $X(T)$ is the weight lattice of a maximal torus) and are related to excellent filtrations on $B$-modules, see Section 2.7. The costandard objects with respect to another highest weight category structure are denoted $Q(\lambda)$, $\lambda\in X(T)$ and are related to relative Schubert filtrations on $B$-modules, see Section 2.3. It is these modules that will play a pivotal role in the construction of full exceptional collections on flag schemes. We refer the reader to [46] for the highest weight category approach to ${\mathrm{Rep}}(B)$; in [47] those structures are somewhat hidden in the background.
Coming back to the examples in rank two, for notational convenience let $\langle X_{v}\rangle$ denote the corresponding set of $B$-modules, $v$ running over the Weyl group $W$. Let ${\mathrm{rep}}(B)$ be the category of finitely generated $B$-modules. The first key observation is that in each type in rank two the $B$-modules $X_{v}$ have a relative Schubert filtration by modules $Q(\lambda)$ and that the dual $B$-modules $X_{v}^{*}$ have an excellent filtration by modules $P(\lambda)$; the second key observation is that the $X_{v}$’s involve in several ways the Steinberg weights $e_{v}$ known from the Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ of the representation ring $R(B)={\sf K}_{0}({\mathrm{rep}}(B))$ as a module over the representation ring $R(G)$, see Section 3.2.
Now it is the Steinberg basis and combinatorics related to it that take the central stage. In the set of modules $Q(\lambda)$ that occur in the filtration of $X_{v}$, the weight $\lambda$ of the $B$-socle of $Q(\lambda)$ is always a Steinberg weight $e_{w}$ with $w\geq v$ in the Bruhat order on $W$. Moreover, the modules $P(-\lambda)^{*}$ that occur in the other filtration of $X_{v}$ are of the form $\nabla_{\mu}\otimes P(-e_{w})^{*}$ with $\mu$ dominant, $w\leq v$ in the Bruhat order. We show in Theorem 8.16 that one still gets a basis of $R(B)$ when replacing a few ${\sf k}_{e_{v}}$’s with $Q(e_{v})$ or $P(-e_{v})^{*}$, see Section 8.
The characters of the modules $Q(e_{v})$ and $P(-e_{w})$ may be computed with Demazure operators, and with computer assisted computations in rank three we have found a close relation in the $T$-equivariant ${\sf K}$-group ${\sf K}_{T}(G/B)$ between the classes of the $\mathcal{L}(P(-e_{w}))$ and the Schubert basis of the free $R(T)$-module ${\sf K}_{T}(G/B)$. We show in Theorem 3.1 that the transition matrix between the basis $\{\mathcal{L}(P(-e_{w}))\}_{w\in W}$ and the Schubert basis of ${\sf K}_{T}(G/B)$ is always invertible and triangular, up to a permutation of rows and columns. Similarly, we show there is a triangular transition matrix comparing the basis $\{\mathcal{L}(Q(e_{v}))\}_{v\in W}$ with the opposite Bruhat cell basis, see Section 5.
This triangularity turns out to be the key to the desired ${\mathrm{Ext}}$-vanishing required for exceptional collections. Precisely, much of the required coherent cohomology vanishing is assured by the results from [47], e.g. Theorems 6.2 and 6.4. They allow to reduce the problem to Euler characteristic computations. Then the crucial vanishing results are provided by the results of Section 5. The precise statement is in Theorem 6.5.
Our way to construct full exceptional collections on flag varieties $G/B$ goes through the study of the derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a category which is linear over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ (a $G$-linear category for short), see Section 8.2. Note that the action of the latter category on the former is given by the restriction functor ${\mathrm{res}}_{B}^{G}$, cf. the aforementioned quote from Donkin. We first construct semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $G$-linear category into pieces, each equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ (a $G$-linear semiorthogonal decomposition for short, see Definition 20). Sections 8 and 9 are devoted to constructing objects $X_{p}$, $p\in W$ labelled by the elements of the Weyl group; these objects $X_{p}$, $p\in W$ will serve as the basis for constructing $G$-semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. This requires choosing a total order $\prec$ on $W$ that refines the Bruhat order. This phase itself breaks up into two steps: given an element $p\in W$, the first step constructs a $G$-semiorthogonal decomposition (called “cut at $p$”, see Section 9.1) of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ into two subcategories
$\langle\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}),\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})\;\rangle.$ (2)
In turn, to establish (2) one needs the crucial vanishing result of Corollary 6.7 that ensures semiorthogonality of those subcategories, and one also needs Theorem 8.15 which says that the triangulated hull of the union of the two categories in (2) is the whole of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. The latter property is a corollary of the generation property saying that the $B$-modules in (2) generate the abelian category ${\mathrm{rep}}(B)$. This relies on Weyl group’s combinatorics and is explained in Section 8.6. As a byproduct of the results of Section 8, we obtain an alternative proof of a theorem of Steinberg [44].
Using $G$-linear semiorthogonal decompositions as in (2), we can finally construct the objects $X_{p}$, $p\in W$ of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. In a nutshell, the procedure is as follows. Given an element $p\in W$, associated to it are the $B$-modules $P(-e_{p})^{\ast}$ and $Q(e_{p})$ where $e_{p}$ is the Steinberg weight associated with $p$. We define $X_{p}$ to be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. Similarly, we define $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. Section 10 proves that $X_{p}$ and $Y_{p}$ are isomorphic for any $p\in W$, and that the objects $X_{p}$, $p\in W$ are $B$-exceptional, i.e. $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))}(X_{p},X_{p})={\sf k}$.
Now each object $X_{p}$, $p\in W$ gives rise to a $G$-linear functor $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Cohomological vanishing statements from Section 6 combined with Theorem 11.1 prove that each $G$-linear functor $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ is full and faithful. Results of Section 8.2 then allow to establish that the collection of full triangulated subcategories ${\rm Im}(\Phi_{p})\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$, $p\in W$, forms a $G$-linear semiorthogonal decomposition of the category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ with respect to the chosen total order $\prec$ on $W$. This is the statement of Theorem 11.5.
The full exceptional collection in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ is now obtained as a consequence of Theorem 11.5; this is taken care of in Theorem 12.1 of Section 12. Specifically, we define objects ${\mathcal{X}}_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ as the images of $X_{p}$, $p\in W$ under the composition of the sheafification functor $\mathcal{L}$ with the forgetful functor ${\sf For}:{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}}(\mathop{\sf{Coh}}\nolimits^{G}(G/B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. The required cohomological vanishing for the objects ${\mathcal{X}}_{p}$, $p\in W$ is in fact provided by what has been previously done in Theorem 11.5, while the generating property is assured by the results of Section 8. For the details, see Theorem 12.1.
The appearance of $G$-equivariant vector bundles in the full exceptional collection is predicted by [39, Lemma 2.2]. We also knew from the rank two examples that we had to look for ${\mathcal{X}}_{v}$ inside the intersection of the hull of $\{\mathcal{L}(P(-e_{v})^{*})\}_{v\preceq p}$ with the hull of $\{\mathcal{L}(Q(e_{v}))\}_{v\succeq p}$.
In section 13 we extend our results to the case of a parabolic subgroup $P\subset G$ with $B\subset P$. We do not need to refer explicitly to highest weight category structures on ${\mathrm{rep}}(P)$. Such structures on ${\mathrm{rep}}(P)$ are available by [21] and appear implicitly in [40] the results of which also bear on the quivers from Section 14.
1.3 Discussion
The forgetful functor ${\mathrm{rep}}(B)\rightarrow{\mathop{\sf{Coh}}\nolimits}(G/B))$ is at the core of Kuznetsov-Polishchuk’s construction of exceptional collections on maximal Grassmannians, [34]. Starting off from a full exceptional collection in the equivariant category and suitably mutating the collection, they arrive at an exceptional collection of objects ${\mathcal{E}}^{\lambda}$ in the non-equivariant derived category where $\lambda$ runs over a certain collection of weights in $X(T)$. The relation of their construction to the one of the present paper is a subject of further research. We note that the objects ${\mathcal{X}}_{p}$, $p\in W$ are amenable to calculations; thus, given a group $G$ and a maximal parabolic subgroup $P\subset G$, the first step in relating the objects ${\mathcal{X}}_{p}$ to the objects ${\mathcal{E}}^{\lambda}$ of [34] is to compute the direct images $\pi_{\ast}{\mathcal{X}}_{p}$, $p\in W$ along the projection $\pi:G/B\rightarrow G/P$.
Furthermore, the objects ${\mathcal{E}}^{\lambda}$ of [34] are a priori complexes of coherent sheaves. Quoting from loc.cit.: “Note that a priori the exceptional objects ${\mathcal{E}}^{\lambda}$ constructed above are complexes. However, we have the following Conjecture 4.1: For any exceptional block $B\subset P_{L}^{+}$ and $\lambda\in B$ the object ${\mathcal{E}}^{\lambda}$ is a vector bundle.” At the moment, it is unclear whether the objects ${\mathcal{X}}_{p}$, $p\in W$ are pure with respect to the standard $t$-structure on $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Evidence for their purity comes from the rank two cases, see Section 14. It is natural to conjecture that ${\mathcal{X}}_{p}$, $p\in W$ are always pure.
A very interesting question is to understand a relation of our results to baric structures/staggered $t$-structures on the derived categories of equivariant coherent sheaves from [1] and [2]. Further, full exceptional collections on flag varieties appear through the computations related to the dual Steinberg basis of ${\sf K}^{0}_{G}(G/B)$, [14]. These, in turn, are related to Lusztig’s asymptotic affine Hecke algebra and to Lusztig’s canonical basis in the $T$-equivariant $\sf K$-theory of Springer fibres. Ultimately, the problem is to relate the main construction of the objects $X_{p}$, $p\in W$ of the present paper to Lusztig’s canonical basis. Finally, the full exceptional collections in rank two from Section 14 possess a property that their duals are obtained from the initial ones by applying the Grothendieck duality followed by tensoring with the square root of the canonical sheaf. In particular, the dual collections in rank two also consist of pure objects. This property allows to put a highest weight category structure on the exact category of vector bundles ${\sf Vect}(G/B)$ in rank two cases and seems to be a manifestation of Ringel duality. Thus, a natural question is to see whether that property holds in general case; in particular, this means computing the dual collection to ${\mathcal{X}}_{p}$, $p\in W$.
It could be more economical to state some of the principal results in the language of quotient stacks; e.g., the passage from the semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear category to that of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ as a ${\sf k}$-linear category could be deduced from the base change theorem for semiorthogonal decompositions, [33]. In the interests of full transparency of the arguments and for convenience of the reader, we have resorted to the more classical language in formulating and proving our results.
Notation
1.4 Semisimple groups
Let ${\mathbb{G}}\rightarrow\mathbb{Z}$ be a simply connected semisimple Chevalley group scheme (a smooth split affine group scheme over ${\rm Spec}(\mathbb{Z})$ whose geometric fibres are connected simply connected semisimple algebraic groups), and ${\mathbb{G}}/{\mathbb{B}}\rightarrow\mathbb{Z}$ be the corresponding Chevalley flag scheme (resp., the corresponding parabolic subgroup scheme ${\mathbb{G}}/{\mathbb{P}}\rightarrow{\mathbb{Z}}$ for a standard parabolic subgroup scheme ${\mathbb{P}}\subset\mathbb{G}$ over ${\mathbb{Z}}$). Then ${\mathbb{G}/\mathbb{P}}\rightarrow{\rm Spec}({\mathbb{Z}})$ is flat and any line bundle $\mathbb{L}$ on ${\mathbb{G}/\mathbb{P}}$ also comes from a line bundle on ${\mathbb{G}/\mathbb{B}}$.
Given an affine algebraic group $H$, flat over a noetherian base ring ${\sf k}$, an $H$-module is a ${\sf k}[H]$-comodule (a rational representation). The abelian category of $H$-modules is denoted ${\mathrm{Rep}}(H)$ and ${\mathrm{rep}}(H)$ denotes the full abelian subcategory of representations which are finitely generated over the base ring ${\sf k}$. (cf. [3, Section 1]).
Our base ring ${\sf k}$ is usually a field or ${\mathbb{Z}}$. We will discuss the case ${\sf k}={\mathbb{Z}}$ separately, or leave the necessary modifications to the reader. For instance, when $M$ is a finitely generated free module over ${\sf k}={\mathbb{Z}}$, then $\dim_{\sf k}(M)$ means its rank. In our references the field is often assumed algebraically closed, but this assumption is irrelevant in our context because of the straightforward behaviour of the various cohomology groups under field extensions of ${\sf k}$.
Let $G$ be obtained from ${\mathbb{G}}$ by base change along ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$. When ${\sf k}={\mathbb{Z}}$ this just means that $G=\mathbb{G}$.
Now let ${\sf k}$ be a field of arbitrary characteristic. So $G$ is a split semisimple simply connected affine algebraic group. Let $T$ be a split maximal torus of $G$, obtained by base change from a split maximal torus $\mathbb{T}$ in $\mathbb{B}$. Let $W=N_{G}({T})/T$ be the Weyl group and let $B\supseteq T$ correspond with $\mathbb{B}\supseteq\mathbb{T}$. Let $X=X(T)$ be the weight lattice and $\Phi$ be the root system. We choose the system $\Phi^{+}$ of positive roots which makes $B$ the negative Borel and let $\Pi$ be the set of simple roots. Let (,)(~{},~{})( , ) be a non-singular, symmetric, positive definite $W$-invariant form on ${\mathbb{R}}\otimes_{\mathbb{Z}}X$ and let $X(T)_{+}=\{\lambda\in X:(\lambda,\alpha^{\vee})\geq 0$ for all $\alpha\in S$}, be the set of dominant weights, where $\alpha^{\vee}=\frac{2\alpha}{(\alpha,\alpha)}$ for $\alpha\in\Phi$.
We want to emphasize once again the fact that our convention (which is that of [29]) is that ${B}$ corresponds to the negative roots (as opposed to the conventions of [25], [41], and [46]). For that reason, the translation of various results from loc.cit. that are used extensively in this paper requires some care.
Let $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ be the set of simple roots relative to our choice of $\Phi^{+}$. The fundamental weights $\omega_{i}$ satisfy $(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij}$. A parabolic subgroup of $G$ is usually denoted by $P$; in particular, for a simple root $\alpha\in\Pi$, denote ${P}_{\alpha}$ the minimal parabolic subgroup of $G$ associated to $\alpha$.
Given a $B$-module $M$ we associate to it as in [29, I 5.8] a sheaf $\mathcal{L}(M)$ on $G/B$ with fibre $M$ at the point $B/B$. Given a weight $\lambda\in{\rm X}({T})$, denote $\mathcal{L}_{\lambda}$ the corresponding line bundle on ${G}/{B}$. Here we still follow [29], so that $\mathcal{L}_{\lambda}$ or $\mathcal{L}(\lambda)$ is associated to the one dimensional $B$-module ${\sf k}_{\lambda}$ of weight $\lambda$, not associated to the dual of that representation.
The weight lattice $X$ has a natural partial order $\geq_{d}$, known as the dominance order: for $\lambda,\mu\in X$ we write $\lambda\geq_{d}\mu$ if $\lambda-\mu$ is a sum of positive roots, with repetitions allowed [^2] That is, $\lambda-\mu$ is of the form $\sum_{\alpha\in\Phi^{+}}n_{\alpha}\alpha$, for non-negative integers $n_{\alpha}$.. We denote $w_{0}$ the longest element of the Weyl group $W$ and let $\lambda^{\ast}=-w_{0}\lambda$ for $\lambda\in X(T)_{+}$ (dual or contragredient, cf. [28, Section 2.2]).
For $\lambda\in X$ let $\nabla_{\lambda}$ or $\nabla(\lambda)$ be the induced module ${\mathrm{ind}}_{B}^{G}({{\sf k}}_{\lambda})$ [29, I, Section 3.3]. It is finite-dimensional and non-zero if and only if $\lambda$ is dominant. For $\lambda\in X(T)_{+}$ we denote by $\Delta(\lambda)$ the Weyl module $(\nabla_{\lambda^{\ast}})^{\ast}$. Then $\nabla_{\lambda}$ has simple socle $L(\lambda)$ and $\Delta(\lambda)$ has simple head $L(\lambda)$. If ${\sf k}={\mathbb{Z}}$ then $\nabla_{\lambda}$, $\Delta(\lambda)$ are finitely generated and free over ${\sf k}$, but heads and socles make less sense.
Proposition 1.4.
([29, II Proposition 4.13, B.4]) Let $\lambda\in X(T)_{+}$.
${\mathrm{Ext}}^{i}_{G}({\sf k},\nabla(\lambda))={\mathrm{Ext}}^{i}_{B}({\sf k},{\sf k}_{\lambda})=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=0,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$
1.5 Triangulated categories
Let ${\sf k}$ be a field. Given a ${\sf k}$-linear triangulated category $\mathcal{D}$, equipped with a shift functor $[1]\colon\mathcal{D}\rightarrow\mathcal{D}$ and two objects $A,B\in\mathcal{D}$, we let denote $\mathop{\mathrm{Hom}}\nolimits^{\bullet}_{\mathcal{D}}(A,B)$ the graded ${\sf k}$-vector space $\bigoplus_{i\in\mathbb{Z}}{\mathop{\mathrm{Hom}}\nolimits}_{\mathcal{D}}(A,B[i])$. A full triangulated subcategory ${\sf A}\subset\mathcal{D}$ is a full subcategory which is closed under shifts and taking cones.
A strictly full subcategory is a full subcategory that contains with every object also the objects isomorphic to it.
[…]
2 Rappels: $B$-modules
2.1 Highest weight category structure on ${\mathrm{Rep}}(G)$
Let ${\sf k}$ be a field.
Recall that relative to the choice of the system of positive roots $\Phi^{+}$ there is a partial order $\leq_{d}$ on $X(T)$: given two weights $\lambda,\mu\in X(T)$ we write $\lambda\leq_{d}\mu$ if $\mu-\lambda$ is a nonnegative integer combination of elements of $\Phi^{+}$. Simple modules in ${\mathrm{Rep}}(G)$ are parametrized by the dominant weights $X(T)_{+}$, [29, II, Section 2]. By [13] and [17], there is a highest weight category structure on ${\mathrm{Rep}}(G)$ with the weight poset $(X(T)_{+},\leq_{d})$: the standard modules in this structure are the Weyl modules $\Delta_{\lambda},\lambda\in X(T)_{+}$ and the costandard modules are the dual Weyl modules $\nabla_{\lambda},\lambda\in X(T)_{+}$.
2.2 Joseph-Demazure modules
Let $H\subset G$ be a flat subgroup scheme of $G$. Associated to it are the restriction and induction functors ${{\mathrm{res}}}_{H}^{G}$ and ${{\mathrm{ind}}}_{H}^{G}$, [29, I, Section 3.1]. Recall the definition of Joseph’s functors ${\sf H}_{w},w\in W$, following [46].
Definition 1.
Let $M$ be a $B$-module and $w\in W$. Consider the Schubert variety $X_{w}$ associated to $w$. Then the functor ${\sf H}_{w}:{{\mathrm{Rep}}}(B)\rightarrow{{\mathrm{Rep}}}(B)$ is given by $M\rightarrow{\rm H}^{0}(X_{w},\mathcal{L}(M))$.
Equivalently, the functor ${\sf H}_{w}$ can be defined as in [47, Definition 7.2.1]. Consider the minimal parabolic subgroup ${P}_{\alpha}$ generated by $B$ and ${U}_{\alpha}$. Then
${\sf H}_{w}:={{\mathrm{res}}}_{B}^{{P}_{\alpha_{1}}}{{\mathrm{ind}}}_{B}^{{P}_{\alpha_{1}}}\dots{\mathrm{res}}_{B}^{{P}_{\alpha_{n}}}{{\mathrm{ind}}}_{B}^{{P}_{\alpha_{n}}},$
where $s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{n}}$ is a reduced expression for $w$.
The functor ${\sf H}_{w}$ is left exact; let ${\rm R}{\sf H}_{w}$ denote the right derived functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))$. It restricts to a functor between bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))$.
Definition 2 (Dual Joseph modules $P(\lambda)$).
Let $\lambda\in X(T)$. Let $\lambda^{+}$ be the dominant weight in the Weyl group orbit of $\lambda$ and let $w$ be minimal so that $\lambda=w\lambda^{+}$. The dual Joseph module $P(\lambda)$ is set to be ${\sf H}_{w}(\lambda^{+})$.
2.3 Relative Schubert modules
Definition 3 (Relative Schubert modules $Q(\lambda)$).
Let $\lambda=w\lambda^{+}$ as above. The relative Schubert module $Q(\lambda)$ is set to be the kernel of the (surjective) restriction homomorphism $P(\lambda)\rightarrow{\rm H}^{0}(\partial X_{w},\mathcal{L}(\lambda^{+}))$.
2.4 $B$-cohomological duality
Let $P(\lambda)$ be the dual Joseph module with socle of weight $\lambda$. The modules $Q(\mu)$ (minimal relative Schubert module with socle ${\sf k}_{\mu}$) satisfy:
${\mathrm{Ext}}^{i}_{B}(P(\lambda)^{\ast},Q(\mu))=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=-\mu,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$ (3)
See also 2.4.
2.5 Excellent order
Recall that we have fixed a Weyl group invariant inner product (,)(~{},~{})( , ) on $X({T})\otimes_{\mathbb{Z}}\mathbb{R}$.
Definition 4.
([46]) Let $\lambda,\mu\in X({T})$. Define that $\lambda$ is less than $\mu$ in the excellent order, notation $\lambda\leq_{e}\mu$, if either $(\lambda,\lambda)<(\mu,\mu)$ or $\lambda=w\nu$, $\mu=z\nu$ for some $\nu\in X({T})_{+}$, $w,z\in W$ with $w\leq z$ (in the Bruhat order on $W$).
2.6 Antipodal excellent order
Definition 5.
([46]) Define that $\lambda\leq_{a}\mu$ in the antipodal excellent order if $-\lambda\leq_{e}-\mu$ in the excellent order.
2.7 Highest weight category structures on ${\mathrm{Rep}}(B)$
Theorem 2.1.
([46, Theorem 1.6, (i)]) The category ${\mathrm{Rep}}(B)$ of rational representations of $B$ is a highest weight category with respect to the excellent order. The $P(\lambda)$ are the costandard modules for this order.
Theorem 2.2.
([46, Theorem 1.6, (ii)]) The category ${\mathrm{Rep}}(B)$ of rational representations of $B$ is a highest weight category with respect to the antipodal excellent order. The $Q(\lambda)$ are the costandard modules for this order.
Remark 2.3.
Let ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$ denote the subcategory of $B$-modules whose weights $\mu$ all satisfy $\mu\leq_{a}\lambda$. Then, in the usual terminology, cf. [19, §2], the costandard module $Q(\lambda)$ for $\leq_{a}$ is the injective hull of ${\sf k}_{\lambda}$ in ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$ and the standard module $P(-\lambda)^{*}$ for $\leq_{a}$ is the projective cover of ${\sf k}_{\lambda}$ in ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$.
2.8 Cohomology vanishing for $B$-modules
Theorem 2.4.
[47, Theorem 3.2.6] Let $\lambda,\mu\in X({T})$. Then $H^{p}(B,P(\lambda)\otimes Q(\mu))=0$ for $p>0$.
Corollary 2.5.
Let $\lambda,\mu\in X(T)$, $\nu\in X(T)_{+}$. Then $H^{p}(B,P(\lambda)\otimes Q(\mu)\otimes\nabla_{\nu})=0$ for $p>0$.
Proof.
Combine Theorem 2.4 with Theorem 2.13. ∎
Remark 2.6.
The vanishing statement of Corollary 2.5 exhibits a remarkable interplay among the three highest weight category structures: one on the category ${\mathrm{Rep}}(G)$ from Section 2.1, and the other two on ${\mathrm{Rep}}(B)$ from Section 2.7. Note also that ${\sf k}=P(0)=Q(0)=\nabla(0)$ is costandard in all three highest weight category structures.
2.9 Filtrations on $B$-modules
We recall here the fundamental statements concerning filtrations on $B$-modules. As we emphasized in the Notation section, one has to take some care when citing the literature, because our $B$ is not $B^{+}$. We have to convert to our conventions. That means that dominant often becomes antidominant and vice versa.
We begin with recalling the definition of good filtration on a $G$-module, [17].
Definition 6.
A rational $G$-module $M$ is said to admit a good filtration provided that there exists a finite or infinite increasing filtration
$0=M_{0}\subset M_{1}\subset M_{2}\subset\dots\subset M_{n}\subset\dots;\bigcup_{i>0}M_{i}=M$
such that each $M_{i}/M_{i-1}=\nabla_{\lambda_{i}}$ for some $\lambda_{i}\in X(T)_{+}$.
Remark 2.7.
An unfortunate side effect of this definition is that if $M$ admits a good filtration, its dimension is at most countable, even if $G$ acts trivially on $M$. This makes results like Theorem 2.16 a bit complicated. If one wants to keep representation theory clean, it is better to use [47, Definition 4.1.1], which is similar to the next definition.
Definition 7.
[47, Definition 2.3.6] A $B$-module $M$ is said to have an excellent filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $B$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus P(\lambda_{i})$ for some $\lambda_{i}\in X(T)$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.8.
Modules possessing excellent filtrations are called strong modules in [35].
Definition 8.
[47, Definition 2.3.8] A $B$-module $M$ is said to have a relative Schubert filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $B$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus Q(\lambda_{i})$ for some $\lambda_{i}\in X(T)$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.9.
Modules possessing filtrations by relative Schubert modules are called weak modules in [35].
Definition 9.
A weight $\lambda$ of a $B$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$.
Proposition 2.10.
[47, Proposition 2.3.11] For any $B$-invariant closed subset $S\subset G/B$ and $\lambda\in X(T)_{+}$, the module $H^{0}(S,\mathcal{L}(\lambda))$ has a relative Schubert filtration with layers $Q(w\lambda)$. Moreover, $Q(w\lambda)$ occurs if and only if $w\lambda$ is an extremal weight of $H^{0}(S,\mathcal{L}(\lambda))$ and then has multiplicity one.
Corollary 2.11.
[47, Corollary 2.3.12] The modules ${\sf H}_{w}(\lambda)$ have relative Schubert filtration for all $w\in W$ and $\lambda\in X(T)$.
Theorem 2.12.
[47, Corollary 5.2.3], [36, §5, Corollary 2] Let $\lambda\in X(T)_{+}$ be a regular dominant weight and $\mu\in X(T)$ be an arbitrary weight. Then $\lambda\otimes Q(\mu)$ has excellent filtration.
Theorem 2.13.
[47, Corollary 5.2.7] Let $\lambda\in X({T})$ and $\mu\in X({T})_{+}$, so that $P(w_{0}\mu)=\nabla_{\mu}$. Then $P(\lambda)\otimes\nabla_{\mu}$ has excellent filtration.
Theorem 2.14.
[35, Theorem 2] Let ${\sf k}$ be a field of characteristic zero. Let $n$ be an integer and $M_{1},\dots M_{n}$ be $n$ weak modules (i.e. each $M_{i}$ has a filtration by $Q(\lambda_{i})$’s, $i=1,\dots,n)$. Suppose $\lambda-n\rho$ is dominant. Then the $B$-module $M_{1}\otimes\dots\otimes M_{n}\otimes{\sf k}_{\lambda}$ is strong (i.e. has a filtration by dual Joseph modules $P(\lambda)$).
2.10 Cohomology criteria for $G$-modules
Theorem 2.15.
[17, Theorem 1.1’] If $\lambda\in X(T)_{+}$ and $M$ is a $G$-module with a good filtration then ${\rm H}^{i}(G,\nabla_{\lambda}\otimes M)=0$ for $i>0$.
Theorem 2.16.
[24, Theorem 4, Corollary 7] Let $M$ be a finitely/countably generated rational $G$-module satisfying ${\rm H}^{1}(G,\nabla_{\mu}\otimes M)=0$ for all $\mu\in X(T)_{+}$. Then $M$ admits a good filtration.
Theorem 2.17.
[29, II Proposition 4.21, B.9] Let $\lambda,\mu\in X(T)_{+}$. Then $\nabla_{\lambda}\otimes\nabla_{\mu}$ admits a good filtration.
2.11 Cohomology criteria for $B$-modules
Theorem 2.18.
[46, Theorem 3.2.7] Let $M$ be a $B$-module such that for every $\lambda\in X(T)$, $H^{1}(B,M\otimes Q(\lambda))=0$. Then, $M$ has an excellent filtration.
Remark 2.19.
For the results over ${\mathbb{Z}}$ we will use ${\mathbb{Z}}$-forms $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$ of $P(\lambda)$, $Q(\mu)$ respectively. They are constructed and studied in [47, Chapter 7]. Both $P(\lambda)_{\mathbb{Z}}$ and $Q(\mu)_{\mathbb{Z}}$ are finitely generated and free over ${\mathbb{Z}}$. They share many properties with their counterparts over fields. We often drop the ${\mathbb{Z}}$ from the notations $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$.
3 $T$-equivariant ${\sf K}$-theory of $G/B$
In this section we introduce the theorem on triangular transition matrices.
3.1 Two fixed points
Let ${\sf k}$ still be a field. We will need results from Graham–Kumar [25] on the $T$-equivariant ${\sf K}$-theory ${\sf K}_{T}(G/B)$ of $G/B$. Their $B$ is our $B^{+}$. As both $B$ and $B^{+}$ will be needed, let us consider the $G$-variety $\mathcal{B}$ of Borel subgroups. If $x$ is a $T$ fixed point in $\mathcal{B}$, then its stabilizer $B(x)$ is a Borel subgroup and one identifies $\mathcal{B}$ with $G/B(x)$. Say $x_{+}$ is the $T$ fixed point so that $B(x_{+})=B^{+}$. And $x_{-}$ is the fixed point so that $B(x_{-})=B$. Choose a representative $\dot{w_{0}}$ of $w_{0}$. One has $x_{+}=\dot{w_{0}}x_{-}$. We simply write $x_{+}={w_{0}}x_{-}$. Let $\phi_{0}$ be the isomorphism $B^{+}\to B$ sending $b$ to $\dot{w_{0}}b\dot{w_{0}}^{-1}$. If ${\mathcal{F}}$ is a $G$ equivariant vector bundle on $\mathcal{B}$, and $x$ is fixed by $T$, then the fibre ${\mathcal{F}}_{x}$ is a $B(x)$-module. We have ${\mathcal{F}}_{x_{+}}\cong\phi_{0}^{*}({\mathcal{F}}_{x_{-}})$. Recall that when $M$ is a finite dimensional $B$-module, we denote as in [29] by $\mathcal{L}(M)$ the $G$ equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{-}}=M$. If $M$ is a finite dimensional $B^{+}$-module, we denote by $\mathcal{L}^{+}(M)$ the $G$ equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{+}}=M$.
If ${\mathcal{F}}$ is a $T$ equivariant coherent sheaf on $G/B$ and $N$ is a finite dimensional $T$-module, then
${\mathcal{F}}\otimes_{\sf k}N$
is a $T$ equivariant coherent sheaf.
When working with $B^{+}$ one should define excellent filtrations and relative Schubert filtrations in terms of $P^{+}(\lambda):=\phi_{0}^{*}(P(w_{0}\lambda))$ and $Q^{+}(\lambda):=\phi_{0}^{*}(Q(w_{0}\lambda))$. They have $B^{+}$ socles of weight $\lambda$ and are the costandard modules of [46].
One has
$H^{i}(B^{+},P^{+}(\lambda)\otimes_{\sf k}Q^{+}(\mu))=\begin{cases}{\sf k}&\text{if $i=0$ and $\lambda+\mu=0$}\\ 0&\text{else.}\end{cases}$
3.2 The Steinberg basis
For $v\in W$ the Steinberg weight $e_{v}$ is given by
$e_{v}=v^{-1}\sum_{v^{-1}\alpha_{i}<0}\omega_{i}.$
The Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ consists of the corresponding one dimensional $B$-modules.
It provides by [44] a basis of the representation ring $R(T)$ as a module over the representation ring $R(G)={\sf K}_{0}({\mathrm{rep}}(G))$, where ${\mathrm{rep}}(G)$ is the category of finitely generated $G$-modules. We will find later that one still gets a basis if one replaces a few ${\sf k}_{e_{v}}$ by $Q(e_{v})$ or $P(-e_{v})^{*}$.
3.3 Schubert varieties and opposite Schubert varieties
Let the Schubert variety $X^{+}_{w}$ be the closure of $B^{+}wx_{+}$ and let the opposite Schubert variety $X^{w}$ be the closure of $B^{-}wx_{+}$. Its ‘boundary’ $\partial X^{w}$ is the union of the $X^{v}$ that are strictly contained in $X^{w}$.
The $T$-equivariant ${\sf K}$-theory ${\sf K}_{T}(G/B)$ of $G/B$ is a module for the representation ring $R(T)$. Generators of $R(T)$ are often written $e^{\lambda}$ instead of $[{\sf k}_{\lambda}]$. If $M$ is a finitely generated $T$-module, then its class $[M]$ in $R(T)$ is also written ${\mathrm{Char}}(M)$, the formal character of $M$.
If ${\mathcal{F}}$ is a $T$ equivariant coherent sheaf and $N$ is a finite dimensional $T$-module, then ${\mathcal{F}}\otimes_{\sf k}N$ represents $[N]\cdot[{\mathcal{F}}]$.
If ${\sf k}={\mathbb{C}}$, then we learn from [25] that the $R(T)$-module ${\sf K}_{T}(G/B)$ has a Schubert basis $\{[\mathcal{O}_{X^{+}_{w}}]\}_{w\in W}$. And also an ‘opposite Bruhat cell’ basis $\{[\mathcal{O}_{X^{w}}(-\partial X^{w})]\}_{w\in W}$. They are orthogonal under the $R(T)$-bilinear symmetric pairing $\langle$-,$\rangle$ - , - ⟩ on ${\sf K}_{T}(G/B)$ given by
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(G/B,{\mathcal{F}}\otimes{\mathcal{G}}))\in R(T).$
This pairing makes sense over any field ${\sf k}$.
If ${\mathcal{F}}\otimes{\mathcal{G}}$ is supported on a $T$ stable closed subscheme $Y$, then
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(Y,{\mathcal{F}}\otimes{\mathcal{G}}))\in R(T).$
We put
${\mathcal{P}}_{v}=\mathcal{L}(P(-e_{v}))$
and
${\mathcal{Q}}_{v}=\mathcal{L}(Q(e_{v})).$
They are $G$-equivariant vector bundles, but we also view them as $T$-equivariant vector bundles.
Write $\alpha_{vw}=\langle[\mathcal{O}_{X^{+}_{w}}],[{\mathcal{Q}}_{v}]\rangle$. So
$[{\mathcal{Q}}_{v}]=\sum\alpha_{vw}[\mathcal{O}_{X^{w}}(-\partial X^{w})]$
when ${\sf k}={\mathbb{C}}$.
Write $\beta_{vw}=\langle[\mathcal{O}_{X^{w}}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$. So
$[{\mathcal{P}}_{v}]=\sum\beta_{vw}[\mathcal{O}_{X^{+}_{w}}]$
when ${\sf k}={\mathbb{C}}$.
Our main result concerning these matrices is that, with a suitable reordering of rows and columns, the matrices $(\alpha_{wv})$ and $(\beta_{vw})$ are upper triangular and invertible.
Theorem 3.1 (Triangular transition matrices).
  1. 1.
    $\beta_{vw}=\langle[\mathcal{O}_{X^{w}}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$ vanishes unless $v\;w_{0}\leq w$ in the Bruhat order.
  2. 2.
    $\alpha_{vw}=\langle[\mathcal{O}_{X^{+}_{w}}],[{\mathcal{Q}}_{v}]\rangle$ vanishes unless $w\leq v\;w_{0}$ in the Bruhat order.
  3. 3.
    If $v\,w_{0}=w$, then $\alpha_{vw}=[{\sf k}_{ve_{v}}]$, $\beta_{vw}=[{\sf k}_{-ve_{v}}]$.
Before proving this, we need to recall things about Weyl groups.
4 Rappels: Weyl groups
4.1 The Demazure product
Let $(W,S=\{s_{\alpha_{1}},\dots,s_{\alpha_{r}}\})$ be the Coxeter presentation of the Weyl group of a simple simply connected group $G$ of rank $r$. The Demazure product on $W$ is defined by replacing the relation $s_{\alpha}\cdot s_{\alpha}=e$ for a simple root $\alpha$ with $s_{\alpha}\star s_{\alpha}=s_{\alpha}$. Equivalently, for any $w\in W$ and any simple root $\alpha$, define $w\star s_{\alpha}$ to be the longer of $ws_{\alpha}$ or $w$. A sequence of simple reflections is reduced if and only if the usual product $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ is equal to the Demazure product $s_{\alpha_{i}}\star\dots\star s_{\alpha_{i_{k}}}$. If $v,w\in W$, then $\overline{BvB}~{}\overline{BwB}=\overline{Bv\star wB}$.
4.2 Total order $\prec$ on $W$, refining the Bruhat order
Choose a total order $\prec$ on $W$ that refines the Bruhat order $<$. Thus $w\leq v$ implies $w\preceq v$. And $w\succ v$ implies $w\nleq v$. To get a nice fit with Theorem 3.1 one may also arrange that $w\preceq v$ implies $ww_{0}\succeq vw_{0}$. This is optional.
4.3 Combinatorics
First we recall some facts from [27] about $W$ and its action on $X(T)$. As usual $w_{0}$ denotes the longest element of $W$.
Lemma 4.1.
Let $v,w\in W$. The following are equivalent
  • $v<w$
  • $v^{-1}<w^{-1}$
  • $w_{0}w<w_{0}v$
  • $ww_{0}<vw_{0}$.
Proof.
For the first three, see [27, page 119]. Use that $(ww_{0})^{-1}=w_{0}w^{-1}$, $(vw_{0})^{-1}=w_{0}v^{-1}$. ∎
Lemma 4.2.
Let $w\in W$ and let $\alpha$ be a simple root. Then $\ell(ws_{\alpha})>\ell(w)$ if and only if $w(\alpha)>0$.
Proof.
See [27, page 116]. ∎
Lemma 4.3.
Let $w\in W$. One can successively multiply $w$ on the right by simple reflections (increasing the length by 1) until this is no longer possible and $w_{0}$ is obtained.
Proof.
See [27, page 16]. We usually prefer to multiply from the left. ∎
Lemma 4.4.
Let $w,z\in W$ with $\ell(wz)=\ell(w)+\ell(z)$. Then $w\leq w_{0}z^{-1}$.
Proof.
If $wz=w_{0}$, then it is clear. We argue by induction on $\ell(w_{0})-\ell(wz)$. If $\ell(w_{0})>\ell(wz)$, then there is a simple $s$ with $\ell(swz)>\ell(wz)$. We have $\ell(sw)=\ell(w)+1$ and $\ell(swz)=\ell(sw)+\ell(z)$. So by induction hypothesis $sw\leq w_{0}z^{-1}$. And thus $w\leq sw\leq w_{0}z^{-1}$
Lemma 4.5.
If $\lambda$ is dominant, then its stabilizer $W_{I}$ in $W$ is generated by the simple reflections that fix $\lambda$.
One puts $W^{I}=\{w\in W\mid\ell(ws)>\ell(w)\mbox{ for every simple reflection $s$ that fixes }\lambda\}$.
Every $w\in W$ may be written uniquely as $vz$, where $z\in W_{I}$ and $v\in W^{I}$ is the minimal coset representative of the coset $wW_{I}$.
Proof.
Compare [27, 1.15, page 22]. ∎
Recall that we denote the Demazure product on $W$ by $\star$.
Lemma 4.6.
Let $v,w\in W$. There are $v^{\prime}\leq v$, $w^{\prime}\leq w$ so that $v^{\prime}w=v\star w=vw^{\prime}$, $\ell(v^{\prime})+\ell(w)=\ell(v\star w)=\ell(v)+\ell(w^{\prime})$.∎
Lemma 4.7.
Let $\lambda$, $\mu$ be weights in the same $W$-orbit. Then $\lambda<_{e}\mu$ if and only if $\lambda>_{a}\mu$.
Proof.
If $\lambda\leq_{e}\mu$, then there is a dominant weight $\omega$ and there are $w,z\in W$ with $w\leq z$, $\lambda=w\omega$, $\mu=z\omega$. Then $ww_{0}\geq zw_{0}$ and $-\lambda=wz_{0}(-w_{0}\omega)$, $-\mu=zz_{0}(-w_{0}\omega)$, so $-\lambda\geq_{e}-\mu$, so $\lambda>_{a}\mu$. The converse is proved similarly. ∎
5 Triangularity of transition matrices
In this section we will show that, with a suitable reordering of rows and columns, the matrices $(\alpha_{wv})$ and $(\beta_{vw})$ from subsection 3.3 are upper triangular and invertible.
Recall that a weight $\lambda$ of a $B$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$
Lemma 5.1.
Let $S,S^{\prime}$ be unions of Schubert varieties $\overline{BuB}/B$ in $G/B$ and let $\lambda$ be dominant.
  • The extremal weights of $\Gamma(S,\mathcal{L}(\lambda))$ are the $w\lambda$ with $wx_{-}\in S$.
  • $\Gamma(G/B,\mathcal{L}(\lambda))\to\Gamma(S,\mathcal{L}(\lambda))$ is surjective.
  • If $S^{\prime}\subset S$, then $\Gamma(S,\mathcal{L}(\lambda))\to\Gamma(S^{\prime},\mathcal{L}(\lambda))$ is surjective.
Proof.
See [47, Proposition 2.2.15, Lemma 2.2.11]. ∎
Proposition 5.2.
Let $\mu$ be a dominant weight with $\alpha^{\vee}(\mu)\leq 1$ for all simple roots $\alpha$. Then $Q(-\mu)={\sf k}_{-\mu}$.
Proof.
We may assume ${\sf k}={\mathbb{C}}$, because of the base change properties [47, Chapter 7] of the $Q(\mu)$. Suppose $Q(-\mu)$ is larger than its socle ${\sf k}_{-\mu}$. Let $\nu$ be a weight of the socle of $Q(-\mu)/{\sf k}_{-\mu}$ and let $f$ be a nonzero weight vector of $Q(-\mu)$, of weight $\nu$, mapping to a vector of the socle of $Q(-\mu)/{\sf k}_{-\mu}$. As the socle of $Q(-\mu)$ is its weight space ${\sf k}_{-\mu}$, the vector $f$ can not be fixed by the unipotent radical $U$ of $B$. But it is well known that $U({\mathbb{C}})$ is generated by the $x_{-\alpha}(t)$ with $\alpha$ simple. So there must be such an $x_{-\alpha}(t)$ with $x_{-\alpha}(t)(f)-f$ a nonzero vector in ${\sf k}_{-\mu}$. In particular, $-\mu=\nu+n(-\alpha)$ for some integer $n$. But $\nu$ must be strictly shorter than the unique extremal weight $-\mu$ of $Q(-\mu)$. However, there is no strictly shorter weight in $-\mu+{\mathbb{Z}}\alpha$. ∎
Lemma 5.3.
Let $w\in W$. If $M$ is a finite dimensional $B^{+}$-module with a relative Schubert filtration, then $H^{i}(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(M))=0$ for $i>0$.
Proof.
Compare [46, Theorem 1.9.(a)(ii)]. ∎
Fix $v\in W$. Let $I$ consist of the simple reflections that fix the dominant weight $ve_{v}$. Let $s$ be simple. From the definition of $e_{v}$ and Lemma 4.2 it follows that $\ell(v^{-1}s)>\ell(v^{-1})$ if and only if $s\in I$. In particular, $v^{-1}$ is a minimal coset representative of $v^{-1}W_{I}$.
Proposition 5.4.
Let $u,v\in W$.
  1. 1.
    $-ve_{v}$ is an extremal weight of $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$, but not of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
  2. 2.
    If $-uve_{v}\neq-ve_{v}$, then $-uve_{v}$ is an extremal weight of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
Proof.
Part (1). The extremal weights of $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ are the elements in the $W$ orbit of $-w_{0}ve_{v}$, hence of $-ve_{v}$. Now suppose $-ve_{v}$ is an extremal weight of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Then $-ve_{v}$ can be written as $y(-w_{0}ve_{v})$ with $y\leq z\star v^{-1}w_{0}$ for some $z<v$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq zv^{-1}w_{0}$ for some $z<v$. Put $u=yw_{0}$. Then $-ve_{v}=-uve_{v}$ with $u\geq zv^{-1}$. But $-ve_{v}=-uve_{v}$ implies $u\in W_{I}$, so $u\geq zv^{-1}$ implies $zv^{-1}\in W_{I}$. Thus $z^{-1}\in v^{-1}W_{I}$. This contradicts the minimality of $v^{-1}$ in its coset $v^{-1}W_{I}$.
Part (2). Now consider a weight of the form $-uve_{v}$ with $-uve_{v}\neq-ve_{v}$. We may replace $u$ by its minimal coset representative. As $\ell(u)\geq 1$ there is a simple reflection $s$ with $s\notin I$ and $\ell(u)=\ell(us)+1$. In particular, $u\geq s$. Now $v^{-1}s<v^{-1}$ by the construction of $e_{v}$. Put $z=sv$, $y=uw_{0}$. Then $z<v$, $-uve_{v}=y(-w_{0}ve_{v})$, with $y\leq sw_{0}=z\star v^{-1}w_{0}$. ∎
Proposition 5.5.
Let $v,w\in W$ such that $vw_{0}\nleq w$. Every extremal weight of $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is also an extremal weight of $\Gamma((\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
Proof.
Suppose $\ell((ww_{0})\star(v^{-1}w_{0}))=\ell(ww_{0})+\ell(v^{-1}w_{0})$. By Lemma 4.4 $(ww_{0})\leq v$, so $vw_{0}\leq w$, contrary to our assumption. So $\ell((ww_{0})\star(v^{-1}w_{0}))<\ell(ww_{0})+\ell(v^{-1}w_{0})$ and we have $y<ww_{0}$ with $(ww_{0})\star(v^{-1}w_{0}))=yv^{-1}w_{0}$, $\ell(yv^{-1}w_{0})=\ell(y)+\ell(v^{-1}w_{0})$. Every extremal weight of $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is of the form $z(-w_{0}ve_{v})$ with $z\leq yv^{-1}w_{0}$. And $\Gamma((\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ maps onto $\Gamma(\overline{Byv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ and then onto $\Gamma(\overline{Bzx_{-}},\mathcal{L}(-w_{0}ve_{v}))$ by Lemma 5.1. ∎
Let $S$ be a union of Schubert varieties $\overline{BuB}/B$ in $G/B$ and let $w\in W$.
Proposition 5.6.
Let ${\sf k}$ be a field of finite characteristic. Let $\hat{S}$ be the inverse image of $S$ in $G$.
  1. 1.
    The multiplication map $\overline{BwB}\times^{B}S\to G/B$ has connected fibres.
  2. 2.
    The multiplication map $\hat{S}\times^{B}\overline{BwB}/B\to G/B$ has connected fibres.
Proof.
The first part is [47, Proposition 6.1.9]. (See also [35, Proposition 18].) It implies that the multiplication map $\overline{BwB}\times^{B}\hat{S}\to G$ has connected fibres. Taking inverses in $G$, ($x\mapsto x^{-1}$), one concludes that a map like the multiplication map $\hat{S}\times^{B}\overline{BwB}\to G$ also has connected fibres. ∎
Remark 5.7.
The result also holds if ${\sf k}$ is a field of characteristic zero, but we will not need that.
Proposition 5.8.
Let $\lambda$ be a dominant weight.
  1. 1.
    $H^{0}(\overline{BwB}S,\mathcal{L}(\lambda))\to H^{0}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))$ is an isomorphism.
  2. 2.
    $H^{0}(\hat{S}\overline{BwB},\mathcal{L}(\lambda))\to H^{0}(S,\mathcal{L}(H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda)))$ is an isomorphism.
  3. 3.
    $H^{i}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))=0$ for $i>0$.
  4. 4.
    $H^{i}(S,\mathcal{L}(H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda)))=0$ for $i>0$.
Proof.
First assume ${\sf k}$ is a field of finite characteristic.
Proof of part 1. Consider the multiplication map $m:\overline{BwB}\times^{B}S\to\overline{BwB}S$. It has connected fibres, is separable, proper, has a target that is Frobenius split, so by [47, Proposition 6.1.6] we have $m_{*}\mathcal{O}_{\overline{BwB}\times^{B}S}=\mathcal{O}_{\overline{BwB}S}$. By the projection formula we have $m_{*}m^{*}\mathcal{L}(\lambda)=\mathcal{L}(\lambda)$. So $H^{0}(\overline{BwB}S,\mathcal{L}(\lambda))\cong H^{0}(\overline{BwB}\times^{B}S,m^{*}\mathcal{L}(\lambda))$. And $\overline{BwB}\times^{B}S\to\overline{BwB}$ is a Serre fibration with fibre $S$, so $H^{0}(\overline{BwB}\times^{B}S,m^{*}\mathcal{L}(\lambda))\cong H^{0}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))$.
Part 2 follows in a similar way.
Part 3 is basically the fact that $H^{0}(S,\mathcal{L}(\lambda))$ has relative Schubert filtration and is therefore ${\sf H}_{w}$ acyclic, cf. Lemma 5.3.
Part 4. By the previous part the result holds when $S$ is a Schubert variety. There are only finitely many possibilities for $S$, so we may assume the result for any union of Schubert varieties that is strictly contained in $S$. If $S$ is not a Schubert variety, write $S=S^{\prime}\cup S^{\prime\prime}$ where $S^{\prime}$, $S^{\prime\prime}$ are strictly smaller. Put $M=H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda))$. Consider the Mayer–Vietoris sequence
$0\to H^{0}(S,\mathcal{L}(M))\to H^{0}(S^{\prime},\mathcal{L}(M))\oplus H^{0}(S^{\prime\prime},\mathcal{L}(M))\to H^{0}(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M))\to\cdots.$
By part 1 the map $H^{0}(S^{\prime\prime},\mathcal{L}(M))\to H^{0}(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M))$ may be identified with the surjective map $H^{0}(\overline{BwB}S^{\prime\prime},\mathcal{L}(\lambda))\to H^{0}(\overline{BwB}(S^{\prime}\cap S^{\prime\prime}),\mathcal{L}(\lambda))$. The result follows.
By the universal coefficient theorem all four parts go through when ${\sf k}={\mathbb{Z}}$ or when ${\sf k}$ is a field of characteristic zero [29, I Proposition 4.18, 5.8(2)].
5.1 The matrix $\beta_{vw}$
Recall that $\beta_{vw}=\langle[\mathcal{O}_{X^{w}}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$. We want to pair $[\mathcal{O}_{X^{w}}(-\partial X^{w})]$ with $[{\mathcal{P}}_{v}]=[\mathcal{L}(P(-e_{v}))]$. As $0\to\mathcal{O}_{X^{w}(-\partial X^{w})}\to\mathcal{O}_{X^{w}}\to\mathcal{O}_{\partial X^{w}}\to 0$ is exact, it suffices to compute $\langle[\mathcal{O}_{X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ minus $\langle[\mathcal{O}_{\partial X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$.
Note that $X^{w}=\overline{Bww_{0}x_{-}}$.
And $P(-e_{v})=\Gamma(\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
We hope that $\langle[X^{w}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$ vanishes unless $vw_{0}\leq w$. But the propositions above show that $\langle[\mathcal{O}_{X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Similarly $\langle[\mathcal{O}_{\partial X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$, where $\partial^{w,v}$ equals $(\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}B}x_{-}$. Both modules have a relative Schubert filtration and we can get a grip on them by inspecting the extremal weights. (The weights that have the same length as the dominant weight $ve_{v}$.)
There are three cases.
  • $vw_{0}=w$. So that is about $Q(\overline{Bw_{0}x_{-}},\partial^{vw_{0},v},\mathcal{L}(-w_{0}ve_{v}))$, the kernel of the surjective map from $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ to $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$. One checks it to be $Q(-ve_{v})={\sf k}_{-ve_{v}}$.
  • $vw_{0}\nleq w$. Then one sees that $Q(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$ vanishes.
  • $vw_{0}<w$. No claims here.
We conclude that the matrix $\beta_{vw}=\langle[\mathcal{O}_{X^{w}}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$ behaves as claimed in Theorem 3.1.
5.2 The matrix $\alpha_{vw}$
So let us turn to $\alpha_{vw}=\langle[\mathcal{O}_{X_{w}}],[{\mathcal{Q}}_{v}]\rangle$.
Now we are dealing with the situation where the Borel subgroup has positive roots, and anti-dominant weights are to be used as in [47]. For instance, Proposition 5.8 gives
Proposition 5.9.
Let $\lambda$ be an anti-dominant weight. And let $S$ be a $B^{+}$ invariant closed subset of $G/B^{+}$.
  1. 1.
    $H^{0}(\overline{B^{+}wB^{+}}S,\mathcal{L}^{+}(\lambda))\to H^{0}({\overline{B^{+}wB^{+}}/B^{+}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))$ is an isomorphism.
  2. 2.
    $H^{i}({\overline{B^{+}wx_{+}}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))=0$ for $i>0$. ∎
We now need to understand $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\phi_{0}^{*}(Q(e_{v}))))$, the kernel of the map $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))\to\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$, which is a surjective map by Lemma 5.1, Lemma 5.9.
Lemma 5.10.
Let $J=w_{0}Iw_{0}$ consist of the simple reflections that fix $w_{0}ve_{v}$.
  1. 1.
    $w_{0}v^{-1}w_{0}$ is a minimal coset representative in $W/W_{J}$.
  2. 2.
    $\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}})$ contains the union of the $\overline{B^{+}w_{0}v^{-1}w_{0}sx_{+}}$ with $s$ simple, $s\notin J$.
Proof.
Let $\alpha$ be a simple root. As $e_{v}$ is a Steinberg weight, $w_{0}v^{-1}(\alpha)>0$ if and only if $s_{\alpha}\notin I$. So $w_{0}v^{-1}w_{0}(-w_{0}(\alpha))<0$ if and only if $s_{-w_{0}\alpha}\notin J$. Thus if $\beta$ is a simple root, then $w_{0}v^{-1}w_{0}(\beta)<0$ if and only if $s_{\beta}\notin J$. ∎
Proposition 5.11.
Let $u,v\in W$.
  1. 1.
    $ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$, but not of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
  2. 2.
    If $uve_{v}\neq ve_{v}$, then $uve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v})))$.
Proof.
Part (1). The extremal weights of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ are elements in the $W$ orbit of $w_{0}e_{v}$. The socle of $P^{+}(w_{0}e_{v})$ has weight $w_{0}e_{v}=w_{0}v^{-1}w_{0}w_{0}ve_{v}$ and $\ell(vw_{0})+\ell(w_{0}v^{-1}w_{0})=\ell(w_{0})$, so $(vw_{0})\star(w_{0}v^{-1}w_{0})=w_{0}$, and $w_{0}w_{0}ve_{v}$ is a weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}ve_{v})))=\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(H^{0}(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}},\mathcal{L}^{+}(w_{0}ve_{v})))$. Now suppose $ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. Then $ve_{v}$ can be written as $y(w_{0}ve_{v})$ with $y\leq vw_{0}\star z$ for some $z<w_{0}v^{-1}w_{0}$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq vw_{0}z$ for some $z<w_{0}v^{-1}w_{0}$. Put $u=w_{0}y$. Then $ve_{v}=yw_{0}ve_{v}=w_{0}uw_{0}ve_{v}$ with $u\geq w_{0}vw_{0}z$. But $uw_{0}ve_{v}=w_{0}ve_{v}$ implies $u\in W_{J}$, so $u\geq w_{0}vw_{0}z$ implies $w_{0}vw_{0}z\in W_{J}$. This contradicts the minimality of $w_{0}v^{-1}w_{0}$ in its coset $w_{0}v^{-1}w_{0}W_{J}$.
Part (2). Now consider a weight of the form $uve_{v}$ with $uve_{v}\neq ve_{v}$. Then $w_{0}uw_{0}w_{0}ve_{v}\neq w_{0}ve_{v}$ so $w_{0}uw_{0}\notin W_{J}$. There is a simple $s$, $s\notin J$, with $w_{0}uw_{0}\geq s$. Then $uw_{0}\leq w_{0}s=(vw_{0})\star(w_{0}v^{-1}w_{0}s)$ and $\overline{w_{0}v^{-1}w_{0}sx_{+}}$ is contained in $\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}})$ by Lemma 5.10. So $uw_{0}w_{0}ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. ∎
Proposition 5.12.
Let $v,w\in W$ such that $w\nleq vw_{0}$. Every extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ is also an extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v})))$.
Proof.
The extremal weights of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ are of the form $uw_{0}ve_{v}$ with $u\leq w\star(w_{0}v^{-1}w_{0})$. Suppose $\ell(w)+\ell(w_{0}v^{-1}w_{0})=\ell(ww_{0}v^{-1}w_{0})$. Then $w\leq w_{0}w_{0}vw_{0}$, contrary to the assumption. So $\ell(w)+\ell(w_{0}v^{-1}w_{0})<\ell(ww_{0}v^{-1}w_{0})$ and we may chose $z<w_{0}v^{-1}w_{0}$ with $w\star(w_{0}v^{-1}w_{0})=w\star z$. As $u\leq w\star z$, $uw_{0}ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. ∎
One can now deal with the matrix $\{\alpha_{vw}\}$ in the same manner as for $\{\beta_{vw}\}$.
This ends the proof of Theorem 3.1. We next develop its consequences.
6 ${\mathrm{ind}}$-vanishing for $B$-modules
In this section we draw conclusions about $B$-modules by combining the vanishing in Corollary 2.5 with Theorem 3.1 on triangular transition matrices.
6.1 Cohomological descent from $G/B$ to $B$
Let ${\sf k}$ be a field or ${\mathbb{Z}}$.
Let $\mathcal{F}=\mathcal{L}({\sf F})={G}\times^{B}\sf F$ be a $G$-equivariant sheaf on $G/B$ given by a finitely generated $B$-module $\sf F$, which is projective over ${\sf k}$. By [29, I 3.3, Proposition 5.12] there are isomorphisms
${\rm H}^{p}({B},{\sf k}[{G}]\otimes{\sf F})=\mathrm{R}^{p}{\mathrm{ind}}_{B}^{G}(\sf F)={\rm H}^{p}({G/B},\mathcal{F}),$ (4)
[…]
Lemma 6.1.
Assume $H^{p}(B,{\sf F}\otimes\nabla_{\mu})=0$ for $p>0$ and all $\mu$.
  1. 1.
    Then ${\rm H}^{p}({G/B},\mathcal{F})=0$ for all $p>0$ and ${\rm H}^{0}({G/B},\mathcal{F})$ has a good filtration with with associated graded
    $\bigoplus_{\lambda\in X({T})_{+}}\nabla_{\lambda}\otimes{\rm H}^{0}({B},\nabla_{-w_{0}\lambda}\otimes{\sf F}).$
  2. 2.
    If moreover $H^{0}(B,{\sf F}\otimes\nabla_{\mu})=0$ for all $\mu$, then ${\rm H}^{0}({G/B},\mathcal{F})=0$.
Proof.
[…]
(2) This follows from the preceding item. ∎
Theorem 6.2.
Let $\lambda,\mu\in X(T)$. Then
  1. 1.
    $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))=0$ for $i>0$.
  2. 2.
    ${\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))$ has a good filtration.
Proof.
By Corollary 2.5 this follows from (4) and Lemma 6.1 above. ∎
Corollary 6.3.
$\langle[\mathcal{L}(P(\lambda))],[\mathcal{L}(Q(\mu))]\rangle={\mathrm{ind}}_{B}^{G}(P(\lambda)\otimes Q(\mu))$.∎
The proof of Theorem 6.2 also works over ${\mathbb{Z}}$. Or one may use the universal coefficient theorem [29, I 4.18]. Thus
Theorem 6.4.
Let $\lambda,\mu\in X(T)$. Then
  1. 1.
    $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbb{B}}^{\mathbb{G}}(P(\lambda)_{\mathbb{Z}}\otimes Q(\mu)_{\mathbb{Z}})=0$ for $i>0$.
  2. 2.
    ${\mathrm{ind}}_{\mathbb{B}}^{\mathbb{G}}(P(\lambda)_{\mathbb{Z}}\otimes Q(\mu)_{\mathbb{Z}})$ has a good filtration.∎
Theorem 6.5.
If ${\sf k}={\mathbb{C}}$, then
  1. 1.
    ${\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))$ vanishes unless $w\leq v$.
  2. 2.
    ${\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v}))={\sf k}$.
Proof.
$\langle[{\mathcal{P}}_{v}],[{\mathcal{Q}}_{w}]\rangle=\sum_{y,z\in W}\alpha_{wz}\beta_{vy}\langle[\mathcal{O}_{X^{+}_{y}}],[\mathcal{O}_{X^{z}}(-\partial X^{z})]\rangle=\sum_{y\in W}\alpha_{wy}\beta_{vy}$. Now apply Theorem 3.1. ∎
We wish to get rid of the restriction ${\sf k}={\mathbb{C}}$. As Euler characteristics are robust, this provides no difficulty:
The ${\mathrm{ind}}_{\mathbb{B}}^{\mathbb{G}}(P(\lambda)_{\mathbb{Z}}\otimes Q(\mu)_{\mathbb{Z}})$ are free ${\mathbb{Z}}$-modules of finite rank, so the universal coefficient theorem [29, I 4.18] shows that the theorem generalizes from ${\sf k}={\mathbb{C}}$ to ${\mathbb{Z}}$, which in turn implies the case where ${\sf k}$ is a field, again by the universal coefficient theorem.
Thus
Theorem 6.6.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $v,w\in W$.
  1. 1.
    $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$}\\ 0&\text{else.}\end{cases}$
  2. 2.
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=0$ for all $i$.
Corollary 6.7.
Let $M$, $N$ be finite dimensional $G$-modules.
  1. 1.
    ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{G}(M,N)$ for all $i$.
  2. 2.
    If $w\not\leq v$ then ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$.
Proof.
$(1).$ We have a spectral sequence [29, I Proposition 4.5]
${\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},\mathrm{R}^{j}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w})))\Rightarrow{\mathrm{Ext}}^{i+j}_{B}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{w})).$
By Theorem 6.2 (1), $\mathrm{R}^{j}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=0$ for $j>0$. Thus, the above spectral sequence degenerates, and with Theorem 6.6 this gives ${\mathrm{Ext}}^{i}_{B}(M\otimes N^{*},P(-e_{v})\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},\mathrm{R}^{0}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{v})))={\mathrm{Ext}}^{i}_{G}(M\otimes N^{*},{\sf k})={\mathrm{Ext}}^{i}_{G}(M,N)$.
$(2).$ If $w\not\leq v$, the above spectral sequence still degenerates by Theorem 6.2, and Theorem 6.6 gives ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$. ∎
Remark 6.8.
Let ${\sf k}={\mathbb{Z}}$. Then one should take $M$, $N$ finitely generated and free over ${\mathbb{Z}}$ in this proof. But see Corollary 8.3 below.
7 Rappels: triangulated categories
7.1 Triangulated hull
Definition 10 (Triangulated hull).
Given a set $S$ of objects of a triangulated category $\mathcal{D}$, its triangulated hull ${\mathrm{hull}}(S)$ is the smallest strictly full triangulated subcategory that contains $S$. The triangulated hull of a sequence of subcategories is defined similarly.
7.2 Thick hull
Definition 11 (Thick).
A subcategory of a triangulated category is called thick if it is closed under direct summands of objects.
Definition 12 (Thick hull).
Given a set $S$ of objects of a triangulated category $\mathcal{D}$, its thick hull ${\mathrm{hull}}(S)_{\oplus}$ is the smallest strictly full triangulated subcategory that contains $S$ which is closed under direct summands of objects.
7.3 Admissible subcategories
We follow the exposition of [20, Section 2.1]. Let $\mathcal{N}\subset\sf D$ be a full triangulated subcategory. The right orthogonal to $\mathcal{N}$ is the full subcategory $\mathcal{N}^{\perp}\subset\sf D$ consisting of all objects $X$ such that $\mathop{\mathrm{Hom}}\nolimits_{\sf D}(Y,X)=0$ for any $Y\subset\mathcal{N}$. The left orthogonal ${}^{\perp}\mathcal{N}$ is defined analogously. The orthogonals are also triangulated subcategories.
Definition 13.
A full triangulated subcategory ${\sf A}$ of $\mathcal{D}$ is called right admissible if the inclusion functor ${\sf A}\hookrightarrow\mathcal{D}$ has a right adjoint. Similarly, ${\sf A}$ is called left admissible if the inclusion functor has a left adjoint. Finally, ${\sf A}$ is admissible if it is both right and left admissible.
Lemma 7.1.
[6, Lemma 3.1] Let $\mathcal{N}\subset\mathcal{D}$ be a full triangulated subcategory, and consider its right orthogonal $\mathcal{N}^{\perp}\subset\mathcal{D}$. Then the following are equivalent:
  • (1)
    The category $\mathcal{D}$ is generated by $\mathcal{N}$ and $\mathcal{N}^{\perp}$, i.e. ${\mathrm{hull}}(\mathcal{N}^{\perp},\mathcal{N})=\mathcal{D}$.
  • (2)
    For each object $x\in\mathcal{D}$ there exists an exact triangle $b\rightarrow x\rightarrow c$ with $b\in\mathcal{N},c\in\mathcal{N}^{\perp}$.
  • (3)
    The embedding functor $i_{\ast}:\mathcal{N}\hookrightarrow\mathcal{D}$ has a right adjoint $i^{!}:\mathcal{D}\rightarrow\mathcal{N}$.
  • (4)
    The embedding functor $j_{\ast}:\mathcal{N}^{\perp}\hookrightarrow\mathcal{D}$ has a left adjoint $j^{\ast}:\mathcal{D}\rightarrow\mathcal{N}^{\perp}$.
If the subcategory $\mathcal{N}$ is right (resp., left) admissible, then $\mathcal{N}$ is thick and one can consider the Verdier localization functor with respect to $\mathcal{N}$. We also record the following version of Lemma 7.1:
Proposition 7.2.
[30, Proposition 3.2.8] Let $\mathcal{N}\subset\mathcal{D}$ be a thick subcategory. Then the following are equivalent:
  • (1)
    The inclusion functor $\mathcal{N}\subset\mathcal{D}$ has a right adjoint (i.e., $\mathcal{N}$ is right admissible in $\mathcal{D}$).
  • (2)
    For each object $b\in\mathcal{D}$ there exists an exact triangle $a\rightarrow b\rightarrow c$ with $a\in\mathcal{N},c\in\mathcal{N}^{\perp}$.
  • (3)
    The canonical functor $\mathcal{D}\rightarrow\mathcal{D}/\mathcal{N}$ has a right adjoint.
  • (4)
    The composite $\mathcal{N}^{\perp}\hookrightarrow\mathcal{D}\rightarrow\mathcal{D}/\mathcal{N}$ is a triangle equivalence.
7.4 Semiorthogonal decompositions
If a full triangulated category $\mathcal{N}\subset\mathcal{D}$ is right admissible then by Lemma 7.1, every object $x\in\mathcal{D}$ fits into a unique (up to unique isomorphism) distinguished triangle
$\dots\longrightarrow Y\longrightarrow X\longrightarrow Z\longrightarrow Y[1]\rightarrow\dots$ (5)
with $Y\in\mathcal{N}$ and $Z\in\mathcal{N}^{\perp}$. One then says that there is a semiorthogonal decomposition of $\mathcal{D}$ into the subcategories $(\mathcal{N}^{\perp},\ \mathcal{N})$. More generally, assume given a sequence of full triangulated subcategories $\mathcal{N}_{1},\dots,\mathcal{N}_{n}\subset\mathcal{D}$. Denote $\langle\mathcal{N}_{1},\dots,\mathcal{N}_{n}\rangle:={\mathrm{hull}}(\mathcal{N}_{1},\dots,\mathcal{N}_{n})$, the triangulated subcategory of $\mathcal{D}$ generated by $\mathcal{N}_{1},\dots,\mathcal{N}_{n}$
Definition 14.
A semiorthogonal decomposition of a triangulated category $\mathcal{D}$ is a sequence of full triangulated subcategories $(\mathcal{N}_{1},\dots,\mathcal{N}_{n})$ in $\mathcal{D}$ such that $\mathcal{N}_{i}\subseteq\mathcal{N}_{j}^{\perp}$ for $1\leq i<j\leq n$ and for every object $X\in\mathcal{D}$ there exists a chain of morphisms in $\mathcal{D}$,
[…]
such that a cone $A_{k}$ of the morphism $X_{k}\rightarrow X_{k-1}$ belongs to $\mathcal{N}_{k}$ for $k=1,\dots,n$.
7.5 Mutations
Let $\mathcal{D}$ be a triangulated category and assume $\mathcal{D}$ admits a semiorthogonal decomposition $\mathcal{D}=\langle{\sf A},{\sf B}\rangle$ with ${\sf A}$ and ${\sf B}$ admissible.
Definition 15.
The left mutation of ${\sf B}$ through ${\sf A}$ is defined to be ${\bf L}_{{\sf A}}({\sf B}):={\sf A}^{\perp}$. The right mutation of ${\sf A}$ through ${\sf B}$ is defined to be ${\bf R}_{{\sf B}}({\sf A}):={{}^{\perp}{\sf B}}$.
One obtains semiorthogonal decompositions $\mathcal{D}=\langle{\bf L}_{{\sf A}}({\sf B}),{\sf A}\rangle$ and $\mathcal{D}=\langle{\sf A},{\bf R}_{{\sf B}}({\sf A})\rangle$.
Lemma 7.3.
[32, Lemma 2.7] There are equivalences ${\bf L}_{{\sf A}}:{\sf B}\simeq\mathcal{D}/{\sf A}\simeq{\bf L}_{{\sf A}}({\sf B})$ and ${\bf R}_{{\sf A}}:{\sf A}\simeq\mathcal{D}/{\sf B}\simeq{\bf R}_{{\sf A}}({\sf B})$.
Definition 16.
Let $\mathcal{D}=\langle{\sf A}_{1},\dots,{\sf A}_{n}\rangle$ be a semiorthogonal decomposition with the ${\sf A}_{i}$ admissible. The left dual semiorthogonal decomposition $\mathcal{D}=\langle{\sf B}_{n},\dots,{\sf B}_{1}\rangle$ is defined by
${\sf B}_{i}:={\bf L}_{{\sf A}_{1}}{\bf L}_{{\sf A}_{2}}\dots\bf L_{{\sf A}_{i-1}}{\sf A}_{i}={\bf L}_{\langle{\sf A}_{1},\dots,{\sf A}_{i-1}\rangle}{\sf A}_{i},\quad 1\leq i\leq n.$ (6)
The right dual semiorthogonal decomposition $\mathcal{D}=\langle{\mathcal{C}}_{n},\dots,{\mathcal{C}}_{1}\rangle$ is defined by
${\mathcal{C}}_{i}:={\bf R}_{{\sf A}_{n}}{\bf R}_{{\sf A}_{n-1}}\dots\bf R_{{\sf A}_{i+1}}{\sf A}_{i}={\bf R}_{\langle{\sf A}_{i+1},\dots,{\sf A}_{n}\rangle}{\sf A}_{i},\quad 1\leq i\leq n.$ (7)
Lemma 7.4.
[32, Lemma 2.10] Let $\mathcal{D}=\langle{\sf A}_{1},\dots,{\sf A}_{n}\rangle$ be a semiorthogonal decomposition such that the components ${\sf A}_{k}$ and ${\sf A}_{k+1}$ are completely orthogonal, i.e., $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k},{\sf A}_{k+1})=0$ and $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\sf A}_{k+1},{\sf A}_{k})=0$. Then
${\bf L}_{{\sf A}_{k}}{\sf A}_{k+1}={\sf A}_{k+1}\qquad\qquad and\qquad\qquad{\bf R}_{{\sf A}_{k+1}}{\sf A}_{k}={\sf A}_{k},$ (8)
and both the left mutation of ${\sf A}_{k+1}$ through ${{\sf A}_{k}}$ and the right mutation of ${\sf A}_{k}$ through ${\sf A}_{k+1}$ boil down to a permutation and
$\mathcal{D}=\langle{\sf A}_{1},\dots,{\sf A}_{k-1},{\sf A}_{k+1},{\sf A}_{k},{\sf A}_{k+2},\dots\sf A_{n}\rangle$ (9)
is the resulting semiorthogonal decomposition of $\mathcal{D}$.
Given an object $F\in\mathcal{D}$, we also record explicit descriptions of left and right mutations of $F$ through an admissible admissible subcategory $i:{\sf A}\hookrightarrow\mathcal{D}$. Let $i^{\ast}$ and $i^{!}$ denote left and right adjoint to functors to $i$, respectively. Then the left mutation ${\bf L}_{{\sf A}}(F)$ the (reps. the right mutations ${\bf R}_{{\sf A}}(F)$ of $F$ through ${\sf A}$) are given by:
${\bf L}_{{\sf A}}(F):={\sf Cone}(ii^{!}(F)\rightarrow F),\qquad\qquad{\bf R}_{{\sf A}}(F):={\sf Cone}(F\rightarrow ii^{\ast}(F))[-1].$ (10)
7.6 Exceptional collections
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let ${\sf k}\textrm{-}\mathrm{Mod}$ denote the category of ${\sf k}$-modules, and ${\sf k}\textrm{-}\mathrm{mod}$ the subcategory of finitely generated ${\sf k}$-modules. Exceptional collections in $\sf k$-linear triangulated categories are a special case of semiorthogonal decompositions with each component of the decomposition being equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\sf k}\textrm{-}\mathrm{mod}})$. The above properties of mutations thus specialize to this special case.
Definition 17.
Let $\mathcal{D}$ be a $\sf k$-linear triangulated category.
An object $E\in\mathcal{D}$ of $\mathcal{D}$ is said to be exceptional if there is an isomorphism of graded $\sf k$-algebras
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(E,E)={\sf k}.$ (11)
A collection of exceptional objects $(E_{0},\dots,E_{n})$ in $\mathcal{D}$ is called exceptional if for $1\leq i<j\leq n$ one has
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(E_{j},E_{i})=0.$ (12)
The collection $(E_{0},\dots,E_{n})$ in $\mathcal{D}$ is said to be full if ${\mathrm{hull}}(E_{0},\dots,E_{n})^{\perp}=0.$
7.7 Admissible subcategories from exceptional collections
Assume given an exceptional collection $E_{0},\dots,E_{n}$ in $\mathcal{D}$, and assume
$\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(-,-):\mathcal{D}\times\mathcal{D}\to\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf k}\textrm{-}\mathrm{mod}).$
One proves [6, Theorem 3.2] that the subcategory ${\mathrm{hull}}(E_{0},\dots,E_{n})\subseteq\mathcal{D}$ is admissible in $\mathcal{D}$. If the collection is full, then ${\mathcal{D}}={\mathrm{hull}}(E_{0},\dots,E_{n})$.
If ${\sf A}\subset\mathcal{D}$ is generated by an exceptional object $E$, then by (10) the left and right mutations of an object $F\in\mathcal{D}$ through ${\sf A}$ are given by the following distinguished triangles:
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(E,F)\otimes E\rightarrow F\rightarrow{\bf L}_{\langle E\rangle}(F),\qquad{\bf R}_{\langle E\rangle}(F)\rightarrow F\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(F,E)^{\ast}\otimes E.$ (13)
7.8 Dual exceptional collections
Let ${\sf k}$ be a field.
Definition 18.
Let $X$ be a smooth projective variety, and assume given a full exceptional collection $(E_{0},\dots,E_{n})$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$. The right dual exceptional collection $(F_{n},\dots,F_{0})$ to $(E_{0},\dots,E_{n})$ is defined as
$F_{i}:={\bf R}_{\langle E_{i+1},\dots,E_{n}\rangle}E_{i},\quad{\rm for}\quad 1\leq i\leq n.$ (14)
The left dual exceptional collection $(G_{n},\dots,G_{0})$ to $(E_{0},\dots,E_{n})$ is defined as
$G_{i}:={\bf L}_{\langle E_{1},\dots,E_{i-1}\rangle}E_{i},\quad{\rm for}\quad 1\leq i\leq n.$ (15)
Proposition 7.5.
[20, Proposition 2.15] Let $(E_{0},\dots,E_{n})$ be a full exceptional collection in a triangulated category $\mathcal{D}$. The left dual exceptional collection $\langle F_{n},\dots,F_{0}\rangle$ is uniquely determined by the following property:
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(E_{i},F_{j}[l])=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad l=0,\ i=j,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$ (16)
Similarly, the right dual exceptional collection $\langle G_{n},\dots,G_{0}\rangle$ is uniquely determined by the following property:
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(G_{i},E_{j}[l])=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad l=0,\ i=j,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$ (17)
8 Generating the categories ${\mathrm{rep}}(B)$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
8.1 The categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$
Let ${\sf k}$ be a field or ${\mathbb{Z}}$.
We recall that ${\mathrm{Rep}}(B)$ (resp., ${\mathrm{rep}}(B)$) denotes the abelian category of rational $B$-modules (resp., the abelian subcategory of rational $B$-modules which are finitely generated over ${\sf k}$). We let denote $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(B))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(G))$) the unbounded derived category of ${\mathrm{Rep}}(B)$ (resp., of ${\mathrm{Rep}}(G)$), and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ (resp., $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$) the bounded derived category of the small category ${\mathrm{rep}}(B)$ (resp., of ${\mathrm{rep}}(G)$).
Let ${\mathrm{rep}}_{\mathrm{fr}}(B)$ denote the full subcategory of ${\mathrm{rep}}(B)$ consisting of the representations that are free over ${\sf k}$. Define ${\mathrm{rep}}_{\mathrm{fr}}(G)$ similarly.
We have the exact bifunctors:
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
and (the internal $\mathop{\mathrm{Hom}}\nolimits_{\sf k}$ on ${\mathrm{rep}}_{\mathrm{fr}}(B)$):
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B).$
Put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$, for $M\in{\mathrm{rep}}_{\mathrm{fr}}(B)$. By [29, I, Section 2.7, (4)], there is a natural $B$-module structure on $M^{*}$.
Proposition 8.1 (Resolution property).
Let $H$ be a flat affine group scheme over a PID ${\sf k}$. Then for every finitely generated $H$-module $N$ there is an exact sequence
$0\to L\to M\to N\to 0$
with the $H$-modules $L$, $M$, finitely generated and free over ${\sf k}$.
Proof.
This is a special case of [43, Proposition 3]. ∎
Let ${\sf k}$ be a field or ${\mathbb{Z}}$ again.
The Proposition implies in standard fashion that every bounded complex in ${\mathrm{rep}}(B)$ is quasi-isomorphic to a bounded complex with objects in ${\mathrm{rep}}_{\mathrm{fr}}(B)$. The above functors can thus be extended as exact bifunctors to the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$:
$(-)\otimes_{\sf k}(-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\otimes\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
and
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\times\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\to\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B)).$
In other words, we are using that by the dual of [9, Theorem 10.22, Remark 10.23], the bounded derived category of the abelian category ${\mathrm{rep}}(B)$ is equivalent to the bounded derived category of its exact subcategory ${\mathrm{rep}}_{\mathrm{fr}}(B)$.
We often write $\otimes_{\sf k}$ as $\otimes$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. By [29, I, Section 2.7, (5)], there is an isomorphism $\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,N)=M^{\ast}\otimes N$ for $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. There is $(\mathop{\mathrm{Hom}}\nolimits_{\sf k},\otimes_{\sf k})$-adjunction:
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-\otimes_{\sf k}-,-)=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-)).$
Let $\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathrm{rep}}(B)}({\mathrm{Rep}}(B))$ denote the derived category of complexes of ${\mathrm{Rep}}(B)$ whose cohomology lie in ${\mathrm{rep}}(B)$.
Lemma 8.2.
There is an equivalence of triangulated categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{{\mathrm{rep}}(B)}({\mathrm{Rep}}(B))$.
Proof.
This is a particular case of [37, Appendix, Lemma A.3]. ∎
Let ${\sf k}={\mathbb{Z}}$.
Corollary 8.3.
Let $M,N\in{\mathrm{rep}}(\mathbb{B})$.
  1. 1.
    ${\mathrm{Ext}}^{i}_{\mathbb{B}}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{v}))={\mathrm{Ext}}^{i}_{\mathbb{G}}(M,N)$ for all $i$.
  2. 2.
    If $w\not\leq v$ then ${\mathrm{Ext}}^{i}_{\mathbb{B}}(M\otimes P(-e_{v})^{*},N\otimes Q(e_{w}))=0$ for all $i$.
Proof.
We know this already for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(\mathbb{B})$, cf. Remark 6.8. By Proposition 8.1 the results follow. ∎
8.2 The category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear category
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The restriction functor ${\mathrm{res}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(B))$ is $t$-exact. Its right adjoint is the induction functor ${\mathrm{Rind}}_{B}^{G}$ and the fact that ${\mathrm{res}}_{B}^{G}$ gives a full faithful embedding is a consequence of the Kempf vanishing theorem, [10, Theorem 1.2], [29, B.3] and Section 14.2: it implies that ${\mathrm{Rind}}_{B}^{G}{\mathrm{res}}_{B}^{G}={\sf id}_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))}$.
Proposition 8.4.
The functor ${\mathrm{Rind}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))$ restricts to a functor ${\mathrm{Rind}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Proof.
Set $X=G/B$, $Y={\sf pt}$, $f:X\rightarrow Y$, $H=G$ in [37, Proposition A.12]. ∎
[…]
Proposition 8.5.
[…]
Proof.
Put $X=G/B,H=G$ in [37, Appendix, Corollary A.5]
[…]
Let ${\sf Inv}^{G}$ denote the derived functor of invariants $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\sf k}\textrm{-}\mathrm{Mod})$.
Proposition 8.6.
Let ${\sf F},{\sf G}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{rep}}(B))$ and $\mathcal{F}=\mathcal{L}({\sf F})$, ${\mathcal{G}}=\mathcal{L}({\sf G})\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{G}(G/B)$ be the associated equivariant complexes of coherent sheaves. There there is a canonical functorial isomorphism
[…]
Proof.
Put $X=G/B$ and $H=G$ in [37, Appendix, Proposition A.6]
[…]
8.3 $G$-linear semiorthogonal decompositions
Definition 19.
A triangulated category $\mathcal{D}$ is called $G$-linear if $\mathcal{D}$ is equipped with a monoidal action of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$, i.e. there is a bifunctor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\times\mathcal{D}\rightarrow\mathcal{D}$. Compare [31, Section 2.7].
Our main concern is the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ which becomes a $G$-linear triangulated category under the restriction functor ${\mathrm{res}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Later we will also need $P$, so let us use $P$ instead of $B$, having in mind $P=B$ as an important case.
Definition 20.
A $G$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ is a collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ of full triangulated $G$-linear subcategories such that the collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ is a semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ in the sense of Definition 14.
The following proposition is just a variation of [33, Lemma 2.7]:
Proposition 8.7.
A pair of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear subcategories ${\sf A},{\sf B}\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ is semiorthogonal (i.e.  ${\sf A}\subset{\sf B}^{\perp}$) if and only if the equality ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M)=0$ holds for all $M\in{\sf A}$, $N\in{\sf B}$.
Proof.
($\Leftarrow$). Assume given $M\in{\sf A}$, $N\in{\sf B}$ and assume ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M)=0$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N,M)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\sf k},N^{*}\otimes M)=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}({\sf k},{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M))=0.$
($\Rightarrow$). Let $M\in{\sf A}$, $N\in{\sf B}$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N,M)=0$ for all such $M$, $N$. Let be $L$ an arbitrary object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M))=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\mathrm{res}}_{P}^{G}L,N^{\ast}\otimes_{{\sf k}}M)$
and
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\mathrm{res}}_{P}^{G}L,N^{\ast}\otimes_{{\sf k}}M)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N\otimes_{{\sf k}}{\mathrm{res}}_{P}^{G}L,M)=0;$
the latter equality holds since the subcategory $\sf B$ is $G$-linear and hence stable under tensoring with objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. It follows that $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M))=0$ for an arbitrary object $L\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$; hence, ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M)=0$. ∎
Let us simplify notation and write $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}$ as $\mathop{\mathrm{RHom}}\nolimits_{P}$, and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}$ as $\mathop{\mathrm{RHom}}\nolimits_{G}$.
Definition 21.
A functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))$ is called $G$-linear if for all $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))$, $N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ there are given bifunctorial isomorphisms
$\Phi({\mathrm{res}}_{P}^{G}(N)\otimes M)={\mathrm{res}}_{P}^{G}(N)\otimes\Phi(M).$
Proposition 8.8.
  1. 1.
    Assume given a fully faithful $G$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Then $\Phi$ is isomorphic to the functor $\Phi_{E}(-)=(-)\otimes E$ where $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that $\mathop{\mathrm{RHom}}\nolimits_{P}(E,E)={\sf k}$.
  2. 2.
    Each object $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that ${\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)={\sf k}$ gives a fully faithful $G$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$.
  3. 3.
    Under the assumptions of (1), the object $E$ satisfies ${\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)={\sf k}$.
Proof.
(1) Take $E=\Phi({\sf k})$.
(2) For $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ we have
$\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes E,N\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes N^{\ast},E^{\ast}\otimes E)=$
$\mathop{\mathrm{RHom}}\nolimits_{G}(M\otimes N^{\ast},{\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E))=\mathop{\mathrm{RHom}}\nolimits_{G}(M\otimes N^{\ast},{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{G}(M,N).$
(3) Under the assumptions of (1)
$\mathop{\mathrm{RHom}}\nolimits_{G}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes E,E)=$
$\mathop{\mathrm{RHom}}\nolimits_{P}(M,E^{\ast}\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{G}(M,{\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)).$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{P}^{G}(X_{v}^{\ast}\otimes E)={\sf k}$. ∎
Remark 8.9.
See Section 6.1 for the precise relation between the two vanishing conditions ${\rm R}^{i}{\mathrm{ind}}_{B}^{G}(E^{\ast}\otimes E)=0$ for $i>0$ and $\mathop{\mathrm{RHom}}\nolimits^{i}_{B}(E,E)=0$ for $i>0$.
8.4 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(G/P))$
Definition 22.
Let $\mathcal{D}$ be a compactly generated triangulated category. A set $S$ of compact objects of $\mathcal{D}$ is called a generating set if $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(S,X)=0$ implies $X=0$ and $S$ is closed under the shift functor, i.e. $S=S[1]$.
Proposition 8.10.
Let $X$ be a quasi–compact, separated scheme, and $\mathcal{L}$ be an ample line bundle on $X$. Then the set $\langle\mathcal{L}^{\otimes m}[n]\rangle,m,n\in\mathbb{Z}$ is a generating set for $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(X))$.
Proof.
See [38, Examples 1.10 and 1.11]. ∎
Corollary 8.11.
The set of line bundles $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(T)_{+},n\in\mathbb{Z}\rangle$ is a generating set for $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(G/P))$.
Corollary 8.12.
The smallest thick full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P)$ containing $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(T)_{+},n\in\mathbb{Z}\rangle$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P)$.
Proof.
Follows from [38, Corollary 2.3]. ∎
The goal of the rest of this section—and the two sections after it—is to construct a collection of objects $X_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$, $p\in W$, each satisfying the condition in (2) of Proposition 8.8; by part (1) of the Proposition each of those objects will also be $B$-exceptional. Furthermore, the collection of objects $X_{p}$ will form a $G$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$; the ultimate statement is Theorem 11.5.
8.5 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$
Proposition 8.13.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The smallest strictly full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ that contains the set of modules $\nabla_{\lambda}$, $\lambda\in X(T)_{+}$, is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Proof.
First we show that ${\mathrm{rep}}(G)$ lies in the subcategory. We choose a real valued additive injective height function ${\mathrm{ht}}$ on the weight lattice which is positive on positive roots. Consider a nonzero $M\in{\mathrm{rep}}(G)$. Say $\mu$ is its highest weight with respect to ${\mathrm{ht}}$. Thus ${\mathrm{ht}}(\mu)\geq 0$. Assume that all representations with a smaller highest weight are in the subcategory. We have a map $\Delta_{\mu}\to\nabla_{\mu}$ whose kernel and cokernel have lower weights, so $\Delta_{\mu}$ lies in the subcategory. Let $M_{\mu,\mathrm{triv}}$ denote the weight space $M_{\mu}$ provided with a trivial $G$ action. By [23, Proposition 21] there is a natural map $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}\to M$ whose kernel and cokernel have lower weights. Now notice that $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}$ lies in the subcategory. So $M$ does too. So ${\mathrm{rep}}(G)$ lies in the subcategory. Now use that every object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ is quasi-isomorphic to a bounded complex in ${\mathrm{rep}}(G)$ and that bounded complexes are repeated cones of pure objects. (cf. “Stupid truncations” [33, 2.5].) ∎
Proposition 8.14.
Let $S$ be a set of objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
Then ${\mathrm{hull}}(\{\nabla_{\nu}\otimes M\mid\nu\in X(T)_{+},M\in S\})$ is $G$-linear.
Proof.
[…]
8.6 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
Let ${\sf k}$ be a field or ${\mathbb{Z}}$.
Theorem 8.15.
Given a $p\in W$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(B))$ of the two categories
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}),$
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})$
is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
Proof.
This will take the rest of the section. ∎
8.7 Generating ${\mathrm{rep}}(B)$
We say that a full subcategory of an abelian category has the 2 out of 3 property if, whenever $0\to N_{1}\to N_{2}\to N_{3}\to 0$ is exact and two of the $N_{i}$ are in the subcategory, then so is the third.
8.8 The set $\{M_{v}\}_{v\in W}$
Let us be given a set $\{M_{v}\}_{v\in W}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}(B)$ with the following properties. The multiplicity of the weight $e_{v}$ in $M_{v}$ is one. Every weight $\lambda$ of $M_{v}$ satisfies $(\lambda,\lambda)\leq(e_{v},e_{v})$. If $\lambda$ is a weight of $M_{v}$ with $(\lambda,\lambda)=(e_{v},e_{v})$, then $\lambda$ is a weight of $P(-e_{v})^{*}$. So all weights $\lambda$ of $M_{v}$ precede $e_{v}$ in the antipodal excellent order, notation $\lambda\leq_{a}e_{v}$. Examples of possible choices of $M_{v}$ are ${\sf k}_{e_{v}}$, $Q(e_{v})$, $P(-e_{v})^{*}$. We use $e^{\lambda}$ as another notation for ${\sf k}_{\lambda}$.
Theorem 8.16 (Generation).
The smallest strictly full additive subcategory that
  • contains the $M_{v}$,
  • has the 2 out of 3 property and
  • contains with every $e^{\lambda}$ and every fundamental representation $\nabla(\omega_{i})$ also $e^{\lambda}\otimes\nabla(\omega_{i})$,
is the category ${\mathrm{rep}}(B)$ of finite dimensional $B$-modules.
We first prove two propositions.
Proposition 8.17.
Let $\lambda$ be a weight in the $W$ orbit of the dominant weight $\varpi$. Assume $\alpha_{i}^{\vee}(\varpi)\leq 1$ for all simple $\alpha_{i}$. Take $w$ minimal so that $\lambda=w\varpi$. Let there be an $\alpha_{j}$ such that $\alpha_{j}^{\vee}(\varpi)=1$ and $w\leq ws_{j}$. Let $\mu\in w(\varpi-\omega_{j})+W\omega_{j}$. Then $(\mu,\mu)\leq(\varpi,\varpi)$ and if $(\mu,\mu)=(\varpi,\varpi)$, then there is $v\geq w$ with $\mu=v\varpi$. So $\mu\leq_{a}\lambda$.
Proof.
Put $\tau=\varpi-\omega_{j}$. Let $W_{I}$, $W_{J}$, $W_{K}$ be the stabilizers in $W$ of $\tau$, $\omega_{j}$, $\varpi$ respectively. Choose $z$ minimal so that $w^{-1}\mu=\tau+z\omega_{j}$. Assume $(\mu,\mu)\geq(\varpi,\varpi)$. Choose a reduced expression $s_{k}s_{k-1}\cdots s_{1}$ for $z$ and put $z_{0}={\mathrm{id}}$, $z_{i}=s_{i}z_{i-1}$. Observe for $0<i\leq k$ that $z_{i-1}\omega_{j}$ is a weight of $P(z_{i}\omega_{j})$, which has socle of weight $z_{i}\omega_{j}$. The socle must lie in the $B$-submodule generated by the weight space of weight $z_{i-1}\omega_{j}$. This implies $z_{i}\omega_{j}\leq_{d}z_{i-1}\omega_{j}$, where $\leq_{d}$ is the dominance partial order, defined in 2.1. The path $z_{0}\omega_{j},\cdots,z_{k}\omega_{j}$ from $\omega_{j}$ to $z\omega_{j}$ is thus strictly descending for $\leq_{d}$. Along the path $(\tau,z_{i}\omega_{j})$ can only go down, but $(\tau,u\omega_{j})<(\tau,v\omega_{j})$ implies $(\tau+u\omega_{j},\tau+u\omega_{j})<(\tau+v\omega_{j},\tau+v\omega_{j})$. So the path must consist of steps in directions perpendicular to $\tau$. So $z\in W_{I}$. Now $w$ is minimal in its coset $wW_{K}$, and moreover we have $w\leq ws_{j}$. That makes it minimal in $wW_{I}$ also. This shows $wz\geq w$. Now note that $\mu=w(\tau+z\omega_{j})=w(z\tau+z\omega_{j})=wz\varpi$. ∎
Proposition 8.18.
Let $\lambda$ be a weight in the $W$ orbit of the dominant weight $\varpi$. Assume $\alpha_{j}^{\vee}(\varpi)>1$ for some simple $\alpha_{j}$. Take $w$ minimal so that $\lambda=w\varpi$. Let $\mu\in w(\varpi-\omega_{j})+W\omega_{j}$. Then $(\mu,\mu)\leq(\varpi,\varpi)$ and if $(\mu,\mu)=(\varpi,\varpi)$, then there is $v\geq w$ with $\mu=v\varpi$. So $\mu\leq_{a}\lambda$.
Proof.
Put $\tau=\varpi-\omega_{j}$. Let $W_{I}$, $W_{J}$ be the stabilizers in $W$ of $\tau$, $\omega_{j}$ respectively. Observe that $W_{I}$ is also the stabilizer $W_{K}$ of $\varpi$. Choose $z$ minimal so that $w^{-1}\mu=\tau+z\omega_{j}$. Assume $(\mu,\mu)\geq(\varpi,\varpi)$. Along the path $z_{0}\omega_{j},\cdots,z_{k}\omega_{j}$ from $\omega_{j}$ to $z\varpi_{j}$, given by an irreducible expression for $z$, the inner product $(\tau,z_{i}\omega_{j})$ can only go down, so the path must consist of steps in directions perpendicular to $\tau$. So $z\in W_{I}=W_{K}$. Now $w$ is minimal in its coset $wW_{K}=wW_{I}$. This shows $wz\geq w$. Now note that $\mu=w(\tau+z\omega_{j})=w(z\tau+z\omega_{j})=wz\varpi$. ∎
Proof of Theorem 8.16.
Let $\varpi$ be dominant and let $\lambda$ be a weight in its $W$ orbit. Choose $w$ minimal so that $\lambda=w\varpi$. Assume that for all $v\geq w$ with $v\varpi\neq\lambda$ the representation $e^{v\varpi}$ is in the subcategory. Assume also that for all weights $\mu$ with $(\mu,\mu)<(\lambda,\lambda)$ the one dimensional representation $e^{\mu}$ of weight $\mu$ is in the subcategory. In other words, assume that $e^{\mu}$ is in the subcategory for all $\mu$ with $\mu<_{a}\lambda$.
We claim that then $e^{\lambda}$ is in the subcategory. The theorem easily follows from the claim by induction along a suitable order on weights.
So let us prove the claim.
There are several cases.
If $\lambda$ is a Steinberg weight $e_{v}$, then one uses the given properties of $M_{v}$.
So we may assume it is not a Steinberg weight.
There are two cases.
The first case is that $\alpha_{i}^{\vee}(\varpi)\leq 1$ for all $i$. As $\lambda$ is not a Steinberg weight there must be a $j$ with $\alpha_{j}^{\vee}(\varpi)=1$ but $w(\alpha_{j})$ positive. Then $\ell(ws_{j})>\ell(w)$ and we are in the situation of Proposition 8.17. Let $\tau=\varpi-\omega_{j}$. Then $\tau$ is strictly shorter than $\lambda$, so $e^{w\tau}$ is in our subcategory, and therefore $N=e^{w\tau}\otimes\nabla(\omega_{j})$ is in the subcategory. The weights $\mu$ of $N$ lie in the convex hull of $w\tau+W\omega_{j}$ and are either shorter than $\lambda$ or they are of the form $v\varpi$ with $v\geq w$. That is $\mu\leq_{a}\lambda$. So $e^{\mu}$ is in the subcategory for all weights of $N$ different from $\lambda$ and $\lambda$ has multiplicity one in $N$. It follows that $e^{\lambda}$ is in the subcategory.
The second case is that some $\alpha_{j}^{\vee}(\varpi)>1$. Put $\tau=\varpi-\omega_{j}$. Again $\tau$ is strictly shorter than $\varpi$. By Proposition 8.18 the weights $\mu$ of $N=e^{w\tau}\otimes\nabla(\omega_{j})$ are either shorter than $\lambda$ or they are of the form $v\varpi$ with $v\geq w$. So $e^{\mu}$ is in the subcategory for all weights of $N$ different from $\lambda$ and $\lambda$ has multiplicity one in $N$. It follows that $e^{\lambda}$ is in the subcategory.
Corollary 8.19.
Let ${\sf k}$ be an algebraically closed field. In the Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ of $R(B)$ over $R(G)$, one may replace some ${\sf k}_{e_{v}}$ by $Q(e_{v})$ and some other ${\sf k}_{e_{w}}$ by $P(-e_{w})^{*}$. The result will still be a basis.
Remark 8.20.
The results of this section give an alternative proof of Steinberg’s theorem [44] that the ${\sf k}_{e_{w}}$ generate $R(B)$ over $R(G)$. See also Corollary 10.2. Our proof assumes only that $G$ is split and ${\sf k}$ is a field or ${\mathbb{Z}}$, while Steinberg takes the field algebraically closed. But by [43, §3] this makes no difference.
Proof of Theorem 8.15.
By Theorem 8.16, for any $p\in W$ the set $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ generates ${\mathrm{rep}}(B)$ as an abelian category. Considering the canonical truncation of an object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ and resolving each cohomology via the above generators of ${\mathrm{rep}}(B)$, we get the statement. One could also use stupid truncations, cf. [33, 2.5]. ∎
9 Construction of the objects $X_{p}$ and $Y_{p}$
By Theorem 8.15, we have
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B)).$ (18)
for all $p\in W$.
9.1 Cut at $p\in W$
Introduce the following notation:
  • ${\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}})$
  • ${\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}})$
  • ${\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})$
  • ${\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$
Denote $\mathcal{D}:={\mathrm{hull}}({\sf Q}_{\succ p}\cup{\sf P}_{\preceq p})$. Then (18) gives that $\mathcal{D}$ is just $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
$\mathcal{D}={\mathrm{hull}}({\sf Q}_{\succ p}\cup{\sf P}_{\preceq p}).$ (19)
By Corollary 6.7, for all $p\in W$:
${\sf Q}_{\succ p}\subseteq{\sf P}_{\preceq p}^{\perp}.$ (20)
Note that we are taking the right orthogonal in $\mathcal{D}$.
By Lemma 7.1 we have a left adjoint of the inclusion of ${\sf P}_{\preceq p}^{\perp}$ into $\mathcal{D}$. Let $j^{*}$ denote this left adjoint of the embedding functor ${\sf P}_{\preceq p}^{\perp}\hookrightarrow\mathcal{D}$. Then we have
$j^{*}(\mathcal{D})={\sf P}_{\preceq p}^{\perp},$ (21)
$j^{*}({\sf P}_{\preceq p})=0,$ (22)
$j^{*}({\sf Q}_{\succ p})={\sf Q}_{\succ p}.$ (23)
Now consider the full subcategory of $\mathcal{D}$ whose objects are the $X$ with $j^{*}(X)\in{\sf Q}_{\succ p}$. It is triangulated and contains both ${\sf P}_{\preceq p}$ and ${\sf Q}_{\succ p}$. So it contains ${\mathrm{hull}}({\sf Q}_{\succ p}\cup{\sf P}_{\preceq p})$ and must be all of $\mathcal{D}$. In other words, $j^{*}(\mathcal{D})={\sf Q}_{\succ p}$. Taken together with equation (21) this shows
${\sf P}_{\preceq p}^{\perp}={\sf Q}_{\succ p}.$ (24)
Thus, by Definition 14, $\mathcal{D}$ has a semiorthogonal decomposition:
$\mathcal{D}=\langle\;{\sf Q}_{\succ p},{\sf P}_{\preceq p}\;\rangle.$ (25)
Then the inclusion ${\sf Q}_{\succ p}\hookrightarrow\mathcal{D}$ has a left adjoint and the inclusion ${\sf P}_{\preceq p}\hookrightarrow\mathcal{D}$ has a right adjoint. So, ${\sf Q}_{\succ p}$ is left admissible in $\mathcal{D}$ and ${\sf P}_{\preceq p}$ is right admissible in $\mathcal{D}$. Dually to (24) we also have
${\sf P}_{\preceq p}={}^{\perp}{\sf Q}_{\succ p}.$ (26)
Moreover:
Proposition 9.1.
$\mathcal{D}=\langle{\sf Q}_{\succeq p},{\sf P}_{\prec p}\rangle$ (27)
Proof.
If $p\neq{\mathrm{id}}$, then $p$ has a predecessor $p^{\prime}$ for the total order $\prec$ on $W$ and ${\sf Q}_{\succeq p}={\sf Q}_{\succ p^{\prime}}$, ${\sf P}_{\prec p}={\sf P}_{\preceq p^{\prime}}$. And if $p={\mathrm{id}}\in W$, then actually $Q(e_{p})=P(-e_{p})^{*}={\sf k}$. ∎
9.2 Defining $X_{p}$ and $Y_{p}$
Let $X_{p}$ be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of ${\sf Q}_{\succeq p}$ into $\mathcal{D}$. More precisely, denote $i^{q}_{\succeq p}:{\sf Q}_{\succeq p}\hookrightarrow\mathcal{D}$ the embedding functor and let ${i^{q}}_{\succeq p}^{\ \ast}$ be its left adjoint. Then
$X_{p}:={i^{q}}_{\succeq p}^{\ \ast}(P(-e_{p})^{*}).$ (28)
By definition, $X_{p}\in{\sf Q}_{\succeq p}$. We have an exact triangle
$\dots\rightarrow P(-e_{p})^{*}\rightarrow X_{p}\rightarrow{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\rightarrow\dots$ (29)
Note that ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$. Therefore,
$X_{p}\in{\mathrm{hull}}(P(-e_{p})^{*}\cup{\sf P}_{\prec p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})={\sf P}_{\preceq p}.$
So,
$X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}.$ (30)
Now let $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of ${\sf P}_{\preceq p}=\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. More precisely, denote $i^{p}_{\preceq p}:{\sf P}_{\preceq p}\hookrightarrow\mathcal{D}$ the embedding functor and let ${i^{p}}_{\preceq p}^{\ !}$ be its right adjoint. Then
$Y_{p}:={i^{p}}_{\preceq p}^{\ !}(Q(e_{p})).$ (31)
By definition, $Y_{p}\in{\sf P}_{\preceq p}$. We have an exact triangle
$\dots\rightarrow Y_{p}\rightarrow Q(e_{p})\rightarrow{\sf cone}(Y_{p}\rightarrow Q(e_{p}))\rightarrow\dots$ (32)
Note that ${\sf cone}(Y_{p}\rightarrow Q(e_{p}))\in{\sf P}_{\preceq p}^{\perp}={\sf Q}_{\succ p}$. Therefore,
$Y_{p}\in{\mathrm{hull}}(Q(e_{p})\cup{\sf Q}_{\succ p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}})={\sf Q}_{\succeq p}.$
So,
$Y_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}.$ (33)
Lemma 9.2.
Let $p\in W$.
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\preceq p})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}).$
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Y_{v}\}_{v\succeq p,\lambda\in X(T)_{+}})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}}).$
In particular, ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})=\mathcal{D}$.
Proof.
As ${\mathsf{cone}}(P(-e_{p})^{*}\to X_{p})$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$, we may argue by induction on the size of $\{\;v\in W\mid v\preceq p\;\}$.
Similarly, ${\mathsf{cone}}(Y_{v}\to Q(e_{v}))$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}})$ and we may use induction on the size of $\{\;v\in W\mid v\succeq p\;\}$. ∎
9.3 Computing morphisms between $X_{p}$ and $Y_{p}$
The objects $X_{v},Y_{v}$ are the key ingredient for constructing semiorthogonal sequences in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Combined with the results of next Section 10, the objects $X_{v}$’s will give the sought-for exceptional collection in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. The goal of this section is to compute $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(X_{v},Y_{v})$. This computation is achieved by the results of Section 6.
Lemma 9.3.
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
  1. 1.
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Y_{v}[i])=\begin{cases}{\sf k}&\text{if $i=0$},\\ 0&\text{else.}\end{cases}$
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(Y_{v},X_{w}[i])=0$ for all $i$.
Proof.
$(1).$ Setting $M=N={\sf k}$ in Corollary 6.7 or Corollary 8.3, we see with Proposition 1.4 that $\mathop{\mathrm{RHom}}\nolimits_{B}(P(-e_{v})^{*},Q(e_{v}))={\rm H}^{\ast}(G,{\sf k})={\sf k}$. Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,Q(e_{v}))$ to the triangle (29), we obtain
$\dots\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathsf{cone}}(P(-e_{v})^{*}\to X_{v}),Q(e_{v}))\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{p},Q(e_{v}))\rightarrow$ (34)
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(P(-e_{v})^{*},Q(e_{v}))\rightarrow\dots$
Now ${\mathsf{cone}}(P(-e_{v})^{*}\to X_{v})\in{\sf P}_{\prec v}$ and $Q(e_{v})\in{\sf P}_{\prec v}^{\perp}$. Thus,
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathsf{cone}}(P(-e_{v})^{*}\to X_{v}),Q(e_{v})[i])=0$
for all $i\in\mathbb{Z}$. We conclude that $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Q(e_{v}))[i]=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(P(-e_{v})^{*},Q(e_{v}))[i]$ for all $i\in\mathbb{Z}$, and by Theorem 6.6 the latter group is trivial if $i\neq 0$ and is isomorphic to $\sf k$ for $i=0$. Thus,
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Q(e_{v}))[i]=\begin{cases}{\sf k}&\text{if $i=0$}\\ 0&\text{else.}\end{cases}$ (35)
Now applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},-)$ to the triangle (32), we obtain
$\dots\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Y_{v})\rightarrow\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Q(e_{v}))\rightarrow$ (36)
$\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},{\sf cone}(Y_{v}\rightarrow Q(e_{v}))\rightarrow\dots$
and ${\sf cone}(Y_{v}\rightarrow Q(e_{v})\in{\sf P}_{\preceq v}^{\perp}$, while $X_{v}\in{\sf Q}_{\succeq v}\cap{\sf P}_{\preceq v}\subset{\sf P}_{\preceq v}$. We conclude that $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},{\sf cone}(Y_{v}\rightarrow Q(e_{v})[i])=0$ for all $i\in\mathbb{Z}$. From the previous exact triangle we obtain isomorphisms $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Y_{v}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(X_{v},Q(e_{v})[i])$ for all $i\in\mathbb{Z}$ and from (35) the latter group is isomorphic to ${\sf k}$ for $i=0$ and is trivial otherwise.
$(2).$ Is an easier consequence of Corollary 6.7 or Corollary 8.3. ∎
As is clear from Lemma 9.3, it is desirable to know that $X_{v}$ is isomorphic to $Y_{v}$ for all $v\in W$.
10 Isomorphism of $X_{p}$ with $Y_{p}$
Let $\lambda\in X(T)$. Choose a dominant weight $\lambda^{+}\in X({T})_{+}\cap W\lambda$ where $W\lambda$ is the $W$-orbit through $\lambda$ and $w\in W$ such that $\lambda^{+}=w\lambda$ with $w$ being minimal with that property.
Then put
${\rm Phys.~{}Rev.~{}E}_{\lambda}=\{v\mid(e_{v},e_{v})<(\lambda,\lambda)\}\cup\{v\mid e_{v}=v^{-1}\lambda^{+}\neq\lambda,~{}v>w\}.$
That is:
${\rm Phys.~{}Rev.~{}E}_{\lambda}=\{v\in W\mid e_{v}<_{a}\lambda\},$
where $<_{a}$ is the antipodal excellent order 5.
We have the following technical variation on Theorem 8.16.
Theorem 10.1.
Let $\lambda,\mu\in X(T)$, $\mu\neq\lambda$. Assume that $\mu$ is a weight of $P(-\lambda)^{*}$ or that $(\mu,\mu)<(\lambda,\lambda)$.
Then $e^{\mu}$ is an object of the triangulated hull of $\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(T)_{+},~{}v\in{\rm Phys.~{}Rev.~{}E}_{\lambda}\}.$
Proof.
We adapt the proof of Theorem 8.16. First let us rephrase what we need to show. We need to show that if $\mu<_{a}\lambda$, then $e^{\mu}$ is in the triangulated hull of
$\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(T)_{+},~{}v\in W,~{}e_{v}<_{a}\lambda\}.$ By Proposition 8.14 this hull is $G$-linear. We want to show by induction along $<_{a}$ that $e^{\mu}$ is in the hull. So assume $e^{\sigma}$ is in the hull for $\sigma<_{a}\mu$.
There are two cases: $\mu$ is a Steinberg weight or it is not. If it is a Steinberg weight, then it is an $e_{v}$ with $e_{v}<_{a}\lambda$. And all weights $\sigma$ of $M_{v}$ that are distinct from $\mu=e_{v}$ satisfy $\sigma<_{a}\mu$. If $\mu$ is not a Steinberg weight, then one reasons as in the proof Theorem 8.16 to see that there is a module $N$ in $\{\nabla_{\nu}\otimes e^{\tau}\mid\nu\in X(T)_{+},~{}(\tau,\tau)<(\mu,\mu)\}$ with $\mu$ a weight of $N$ of multiplicity one, and all weights $\sigma$ of $N$ satisfying $\sigma\leq_{a}\mu$. But $(\tau,\tau)<(\mu,\mu)$ implies $\tau<_{a}\mu$. ∎
For completeness we mention
Corollary 10.2.
Let $\lambda\in X(T)$. Then $e^{\lambda}$ lies in the $R(G)$-submodule of $R(B)$ generated by the $[M_{v}]$ with $e_{v}\leq_{a}\lambda$.
Proof.
‘Same proof’ by induction along $\leq_{a}$. ∎
Let $f_{p}$ be the natural map from $P(-e_{p})^{*}$ to $Q(e_{p})$. Notice that $f_{p}$ factors as a surjection $P(-e_{p})^{*}\to{\sf k}_{e_{p}}$, followed by an injection ${\sf k}_{e_{p}}\to Q(e_{p})$. We have an exact triangle
$\dots\rightarrow{\mathsf{ker}}(f_{p})[1]\to{\mathsf{cone}}(f_{p})\to{\mathsf{coker}}(f_{p})\rightarrow\dots$
Take $M_{v}=Q(e_{v})$ for $v\succ p$ and $M_{v}=P(-e_{v})^{*}$ for $v\prec p$. We do not need to specify $M_{p}$, because $p\notin{\rm Phys.~{}Rev.~{}E}_{e_{p}}$.
Corollary 10.3.
Let $p\in W$. Then ${\mathsf{cone}}(f_{p})$ belongs to the hull of the union of following three sets:
$\{\;\nabla_{\lambda}\otimes Q(e_{v})$$\mid$ $\lambda\mbox{ is dominant, }v\succ p\mbox{ and }(e_{v},e_{v})<(e_{p},e_{p})\;\}\cup$
$\{\;\nabla_{\lambda}\otimes Q(e_{v})$$\mid$ $\lambda\mbox{ is dominant, }v\succ p\mbox{ and }v^{-1}pe_{p}=e_{v}\neq e_{p}\;\}\cup$
$\{\;\nabla_{\lambda}\otimes P(-e_{v})^{*}$$\mid$ $\lambda\mbox{ is dominant, }v\prec p\mbox{ and }(e_{v},e_{v})<(e_{p},e_{p})\;\}.$
In particular, ${\mathsf{cone}}(f_{p})$ belongs to the hull of ${\sf P}_{\prec p}\cup{\sf Q}_{\succ p}$.
Proof.
Apply Theorem 10.1 with $\lambda=e_{p}$ to ${\mathsf{ker}}(f_{p})$ and ${\mathsf{coker}}(f_{p})$. ∎
Remark 10.4.
Observe that the proof involves the partial order $<_{a}$, but the conclusion, that ${\mathsf{cone}}(f_{p})$ belongs to the hull of ${\sf P}_{\prec p}\cup{\sf Q}_{\succ p}$, refers to $\prec$. This conclusion just uses that only $M_{v}$ with $v\neq p$ are used.
We are now in a position to prove the isomorphism $X_{p}=Y_{p}$. Note that ${i^{p}}_{\preceq p}^{\ !}Q(e_{p})=Y_{p}$ and ${i^{p}}_{\preceq p}^{\ !}(X_{p})=X_{p}$ as ${i^{p}}_{\preceq p}^{\ !}={\sf id}_{{\sf P}_{\preceq p}}$ on ${\sf P}_{\preceq p}$. Moreover, the functor ${i^{p}}_{\preceq p}^{\ !}$ annihilates ${\sf Q}_{\succ p}$ as ${\sf Q}_{\succ p}={\sf P}_{\preceq p}^{\perp}$. That is:
${i^{p}}_{\preceq p}^{\ !}({\sf Q}_{\succ p})=0,$
and
${i^{p}}_{\preceq p}^{\ !}({\sf P}_{\prec p})\subset{\sf P}_{\prec p}.$
As ${\mathsf{cone}}(f_{p})$ belongs to the hull of ${\sf P}_{\prec p}\cup{\sf Q}_{\succ p}$, we conclude that
${i^{p}}_{\preceq p}^{\ !}({\mathsf{cone}}(f_{p}))\in{\sf P}_{\prec p}.$ (37)
Consider the exact triangle
$\dots\rightarrow P(-e_{v})^{*}\rightarrow Q(e_{v})\rightarrow{\mathsf{cone}}(f_{p})\rightarrow\dots$ (38)
Applying to the triangle the functor ${i^{p}}_{\preceq p}^{\ !}$ and remembering that ${i^{p}}_{\preceq p}^{\ !}(Q(e_{v}))=Y_{p}$, we obtain
$\dots\rightarrow P(-e_{v})^{*}\rightarrow Y_{p}\rightarrow{i^{p}}_{\preceq p}^{\ !}({\mathsf{cone}}(f_{p}))\rightarrow\dots$ (39)
with ${i^{p}}_{\preceq p}^{\ !}({\mathsf{cone}}(f_{p}))\in{\sf P}_{\prec p}$. Thus
${\mathsf{cone}}(P(-e_{v})^{*}\rightarrow Y_{p})\in{\sf P}_{\prec p}.$ (40)
Now apply to the triangle
$\dots\rightarrow P(-e_{v})^{*}\rightarrow Y_{p}\rightarrow{\mathsf{cone}}(P(-e_{v})^{*}\rightarrow Y_{p})\rightarrow\dots$ (41)
the projection functor ${i^{q}}_{\succeq p}^{\ \ast}$ onto ${\sf Q}_{\succeq p}$. We have ${i^{q}}_{\succeq p}^{\ \ast}(P(-e_{v})^{*})=X_{p}$ and ${i^{q}}_{\succeq p}^{\ \ast}(Y_{p})=Y_{p}$, as $Y_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ by (33). Further ${i^{q}}_{\succeq p}^{\ \ast}({\mathsf{cone}}(P(-e_{v})^{*}\rightarrow Y_{p}))=0$ because of (40), as ${i^{q}}_{\succeq p}^{\ \ast}$ annihilates ${\sf P}_{\prec p}$. Thus, $X_{p}=Y_{p}$ and by Lemma 9.3, (1), the object $X_{p}$ is exceptional in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
11 Semi-orthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $G$-linear category
Let ${\sf k}$ be a field.
Theorem 11.1.
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Let $M$, $N$ be finite dimensional $G$-modules. Let $v,w\in W$.
  1. 1.
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v}[i])={\mathrm{Ext}}^{i}_{G}(M,N)$ for all $i$.
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{w}[i])=0$ for all $i$.
Proof.
$(1).$ Recall that for $p\in W$, the object $X_{p}$ is both the projection of $P(-e_{v})^{*}$ onto the category ${\sf Q}_{\succeq p}$ and the projection of $Q(e_{p})$ onto ${\sf P}_{\preceq p}$. Tensoring the triangles (29) and (32) with modules $M$ and $N$, respectively, we get:
$\dots\rightarrow M\otimes P(-e_{p})^{*}\rightarrow M\otimes X_{p}\rightarrow M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\rightarrow\dots$ (42)
and
$\dots\rightarrow N\otimes X_{p}\rightarrow N\otimes Q(e_{p})\rightarrow N\otimes{\sf cone}(X_{p}\rightarrow Q(e_{p}))\rightarrow\dots$ (43)
We have inclusions ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and ${\sf cone}(Y_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$, and both subcategories ${\sf P}_{\prec p}$ and ${\sf Q}_{\succ p}$ are $G$-linear, thus $M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and $N\otimes{\sf cone}(Y_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$.
We have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p}),N\otimes X_{p})[i])=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ and ${\sf Q}_{\succeq p}={\sf P}_{\prec p}^{\perp}$ by (27). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,N\otimes X_{p})$ to the triangle (42), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])$ for all $i$.
Next, we have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes{\sf cone}(Y_{p}\rightarrow Q(e_{p})[i]))=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf P}_{\preceq p}$ and ${\sf Q}_{\succ p}={\sf P}_{\preceq p}^{\perp}$ by (25). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},-)$ to the triangle (43), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p})[i])$ for all $i$. The latter group is isomorphic to ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p}))$. Now the statement in (1) follows by Corollary 6.7, (1).
$(2).$ The second part follows similarly using Corollary 6.7, (2). ∎
Remark 11.2.
If ${\sf k}={\mathbb{Z}}$, one may use the resolution property Proposition 8.1 to choose a representative of the isomorphism class of $X_{v}$ so that the representative is a bounded complex of modules which are flat over ${\mathbb{Z}}$. With such choice the proof of Theorem 11.1 goes through for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(G)$ and by Proposition 8.1 the result then follows for $M,N\in{\mathrm{rep}}(G)$.
In fact we have
Theorem 11.3.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ and $v,w\in W$.
  1. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,N)$,
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{w})=0$.
Proof.
Part (1). It suffices to treat the case $M={\sf k}$ because $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(L\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(X_{v},L^{*}\otimes N\otimes X_{v})$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,N)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}({\sf k},L^{*}\otimes N)$. By Theorem 11.1 or Remark 11.2 we know the result when $N\in{\mathrm{rep}}(G)$ and $\mathop{\mathrm{RHom}}\nolimits_{G}({\sf k},N)$ is concentrated in at most one degree. For instance, this is the case when $N=\nabla_{\nu}$ for some dominant $\nu$. So it holds for $N$ in the hull $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ of the $\nabla_{\nu},\nu\in X(T)_{+}$.
Part (2) is easier. ∎
Corollary 11.4.
  1. 1.
    ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$
  2. 2.
    If $w\succ v$ then ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w})=0$
Proof.
Part (1)
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},X_{v})=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M,X_{v}^{\ast}\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,{\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})),$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.
Part (2) is easier. ∎
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. In the latter case $G$ means $\mathbb{G}$.
Theorem 11.5.
For $v\in W$, denote $\sf X_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ generated by $\{M\otimes X_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\}$. Then the category $\mathcal{D}$ has a $G$-linear semiorthogonal decomposition
$\mathcal{D}=\langle\sf X_{v}\rangle_{v\in W}$ (44)
with respect to the order $\prec$ on the Weyl group $W$. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Proof.
Denote $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathcal{D}$ the $G$-linear functor $M\rightarrow M\otimes_{{\sf k}}X_{p}$. Now Corollary 11.4 and Proposition 8.8 give that $\Phi_{p}$ is fully faithful.
Let now $w\succ v$. By Corollary 11.4 and [29, I Proposition 3.6] we have ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w}\otimes M)={\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w})\otimes M=0$ for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By Proposition 11.6 below, each subcategory ${\sf X}_{v},v\in W$ is admissible in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Proposition 8.7 then gives that the sequence $\langle\sf X_{v}\rangle_{v\in W}$ of admissible subcategories of $\mathcal{D}$ is $G$-semiorthogonal with respect to the order $\prec$ on $W$. Lemma 9.2 states that the triangulated hull of $\langle\sf X_{v}\rangle_{v\in W}$ coincides with $\mathcal{D}$. Thus, it is a $G$-linear semi-orthogonal decomposition of $\mathcal{D}$. ∎
Let ${\sf k}$ be a field or ${\mathbb{Z}}$.
Proposition 11.6.
For each $p\in W$ the subcategory $\sf X_{v}$ is an admissible subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
Proof.
Recall that by Proposition 8.1 we may extend the exact bifunctors
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
and
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
to the bounded derived category of ${\mathrm{rep}}(B)$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. Then the right adjoint of the inclusion of $\sf X_{v}$ is
${\mathrm{Rind}}_{B}^{G}(-\otimes_{\sf k}X_{p}^{*})\otimes_{\sf k}X_{p}$
and the left adjoint is
$({\mathrm{Rind}}_{B}^{G}((-)^{*}\otimes_{\sf k}X_{p}))^{*}\otimes_{\sf k}X_{p}.$
12 Full exceptional collections in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$
We recall first the sheafification functor $\mathcal{L}$. Let $G$ be a flat $\sf k$-group scheme acting freely (from the right) on a flat $\sf k$-scheme $X$ such that $X/G$ is a scheme. Associated to each $G$-module $M$ is a sheaf $\mathcal{L}(M)=\mathcal{L}_{X/G}(M)$ on $X/G$, [29, I, Section 5.8]. The functor $\mathcal{L}:{{\mathrm{Rep}}}({G})\rightarrow{\rm Sh}(X/G)$ is exact and lands in the subcategory ${\mathsf{QCoh}}(X/G)$ of quasi–coherent $\mathcal{O}_{X/G}$-modules of ${\rm Sh}(X/G)$, [29, I, Proposition 5.9]. Since $\mathcal{L}$ is exact, denote also $\mathcal{L}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({{\mathrm{Rep}}}({G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathsf{QCoh}}(X/G)}{\rm Sh}(X/G)$ [^3] The derived category of complexes on ${\rm Sh}(X/G)$ with quasi-coherent cohomology. its derived functor.
Now set $G:=B,X=G$ (so $X=G/B$) in the definition of the functor $\mathcal{L}$. It restricts to the functor between the bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$. We put ${\mathcal{X}}_{v}:=\mathcal{L}(X_{v})$ for $v\in W$.
[…]
Theorem 12.1.
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$. Let $v,w\in W$.
  1. 1.
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{v}[i])=\begin{cases}{\sf k}&\text{if $i=0$},\\ 0&\text{else.}\end{cases}$
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{w}[i])=0$ for all $i$.
  3. 3.
    The triangulated hull of $\{{\mathcal{X}}_{v}\mid v\in W\}$ is $\mathcal{D}$.
In other words, the collection of objects $({\mathcal{X}}_{v})_{v\in W}$ is a full exceptional collection in $\mathcal{D}$.
Proof.
$(1)$. By Proposition 8.6, the functor $\mathop{\mathrm{RHom}}\nolimits_{G/B}(\mathcal{L}(-),\mathcal{L}(-))$ canonically factors through ${\mathrm{Rind}}_{B}^{G}((-)^{\ast}\otimes(-))$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{G/B}(\mathcal{L}(X_{v}),\mathcal{L}(X_{v}))={\sf For}\circ{\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})$ where $\sf For$ is the forgetful functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\sf k}\textrm{-}\mathrm{mod}})$. This also holds when ${\sf k}={\mathbb{Z}}$. Now by Corollary 11.4 (1), ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.
$(2)$. This is the same argument as in (1), using Corollary 11.4 (2).
[…]
By Lemma 9.2, ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ coincides with the hull of ${\mathrm{rep}}(B)$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. In particular, the abelian category ${\mathrm{rep}}(B)\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ is contained in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. By definition, ${\mathcal{X}}_{v}=\mathcal{L}(X_{v})$; applying the functor $\mathcal{L}$ to the above inclusion, we see that the category of $G$-equivariant coherent sheaves ${\sf Coh}^{G}(G/B)=\mathcal{L}({\mathrm{rep}}(B))$ is contained in the image under $\mathcal{L}$ of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. Recall that the functor $\mathcal{L}$ is monoidal; thus, $\mathcal{L}(\nabla_{\lambda}\otimes X_{v})=\mathcal{L}(\nabla_{\lambda})\otimes{\mathcal{X}}_{v}$. Now $\mathcal{L}(\nabla_{\lambda})$ is a ($G$-equivariant) trivial vector bundle on $G/B$, so the image of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ under $\mathcal{L}$ coincides with the hull $\langle{\mathcal{X}}_{v},v\in W\rangle$ of $\{{\mathcal{X}}_{v},v\in W\}$. We conclude that $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{G}(G/B)$.
Now since $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{G}(G/B)\subset\langle{\mathcal{X}}_{v},v\in W\rangle$, the category $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains all line bundles $\mathcal{L}_{\lambda},\lambda\in X(T)_{+}$, the group $G$ being simply connected. Now $\langle{\mathcal{X}}_{v},v\in W\rangle^{\perp}=0$ by Corollary 8.11. By Lemma 7.1 we get $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))=\langle{\mathcal{X}}_{v},v\in W\rangle$. ∎
12.1 Variations
The ${\mathcal{X}}_{p}$ may depend on the choice of the total order $\prec$ on $W$. Therefore it is not clear that ${\mathcal{X}}_{p}$ is perpendicular to ${\mathcal{X}}_{q}$ when $\ell(p)=\ell(q)$. In fact computer assisted computations in ${\sf K}_{T}(G/B)$ indicate that this fails already for type ${\bf B}_{3}$ with $\ell(p)=\ell(q)=3$ and for type ${\bf C}_{3}$ with $\ell(p)=\ell(q)=2$.
13 Generalised flag varieties
Let $P$ be a parabolic subgroup containing $B$. (If ${\sf k}={\mathbb{Z}}$, then $P$ means $\mathbb{P}$.) We seek a full exceptional collection on $G/P$. Let $W_{P}$ be the Weyl group of $P$, generated inside $W$ by the $s_{\alpha}$ with $\alpha$ simple and $P_{\alpha}\subset P$. Let $W^{P}$ be the set of minimal coset representatives of $W/W_{P}$, cf.  Lemma 4.5. According to Steinberg [44] we may use as generators of the $R(G)$-module $R(P)$, over ${\mathbb{C}}$, the classes of the irreducible $P$-modules with highest weight $e_{v}$ where $v$ runs over $W^{P}$, and $e_{v}$ is still defined as in 3.2. Inspired by that, we restrict our total order $\prec$ from $W$ to $W^{P}$.
Remark 13.1.
The map $W\to W^{P}$ which sends $w\in W$ to the minimal representative of the coset $wW_{P}$ is a poset map. If one has chosen a total order on $W^{P}$ refining the (restriction to $W^{P}$) of the Bruhat order, then this chosen order can be extended to a total order on $W$ that refines the Bruhat order on $W$. We already had to choose $\prec$ on $W$ in section 4.2, so it makes sense to keep that order $\prec$ and restrict it to $W^{P}$.
For $v\in W^{P}$ we will find ${\hat{X}}_{v}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that
Theorem 13.2.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ and $v,w\in W^{P}$.
  1. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes{\hat{X}}_{v},N\otimes{\hat{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,N)$,
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes{\hat{X}}_{v},N\otimes{\hat{X}}_{w})=0$.
Theorem 13.3.
For $v\in W^{P}$, denote ${\hat{\sf X}}_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ generated by $\{M\otimes{\hat{X}}_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\}$. Then the category $\mathcal{D}$ has a $G$-linear semiorthogonal decomposition
$\mathcal{D}=\langle{\hat{\sf X}}_{v}\rangle_{v\in W^{P}}$ (45)
with respect to the order $\prec$ on $W^{P}$. Each subcategory ${\hat{\sf X}}_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Definition 23.
We put $\hat{\mathcal{X}}_{v}:=\mathcal{L}_{G/P}({\hat{X}}_{v})$ for $v\in W^{P}$.
Theorem 13.4.
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Let $v,w\in W^{P}$.
  1. 1.
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(\hat{\mathcal{X}}_{v},\hat{\mathcal{X}}_{v}[i])=\begin{cases}{\sf k}&\text{if $i=0$},\\ 0&\text{else.}\end{cases}$
  2. 2.
    If $w\succ v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(\hat{\mathcal{X}}_{v},\hat{\mathcal{X}}_{w}[i])=0$ for all $i$.
  3. 3.
    The triangulated hull of $\{\hat{\mathcal{X}}_{v}\mid v\in W^{P}\}$ is $\mathcal{D}$.
In other words, the collection of objects $(\hat{\mathcal{X}}_{v})_{v\in W}$ is a full exceptional collection in $\mathcal{D}$.
To prove these Theorems we will need to find replacements–often with hatted notation–of the key ingredients used in the case $P=B$. This will take the rest of the section.
Recall that by [29, II Corollary 4.7] we may view ${\mathrm{rep}}(G)$ as a subcategory of ${\mathrm{rep}}(P)$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ as a subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$, so that we may suppress ${\mathrm{res}}^{G}_{P}$ in notations. Similarly we may suppress ${\mathrm{res}}^{P}_{B}$.
Remark 13.5.
The strategy for constructing $\hat{\mathcal{X}}_{v}$ is the same as for ${\mathcal{X}}_{v}$, but we do not know the precise relation between ${\mathcal{X}}_{v}$, $\hat{\mathcal{X}}_{v}$ and $G/B\to G/P$. More specifically, we do not know if ${\mathcal{X}}_{v}$ is always the pull back of $\hat{\mathcal{X}}_{v}$ for $v\in W^{P}$.
Recall that $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ is the set of simple roots and $(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij}$. If $\alpha=\alpha_{i}$, we also write $\omega_{i}$ as $\omega_{\alpha}$.
Lemma 13.6.
Let $\lambda\in X(T)$, $\alpha\in\Pi$. Then $s_{\alpha}\lambda\leq_{e}\lambda$ if and only if $s_{\alpha}\lambda\geq_{d}\lambda$.
Proof.
Write $\lambda=w\nu$ with $\nu$ dominant and $w$ minimal. We have $s_{\alpha}w>w$ or $s_{\alpha}w<w$. If $s_{\alpha}w>w$, then $\lambda$ is a weight of $P(s_{\alpha}\lambda)$ and the socle of $P(s_{\alpha}\lambda)$, which has weight $s_{\alpha}\lambda$, must lie in the $B$-submodule generated by the weight space of weight $\lambda$. This implies $s_{\alpha}\lambda\leq_{d}\lambda$. The other case is similar. ∎
Notations 13.7.
Let $\Pi^{P}$ be the set of simple roots $\alpha$ for which $P_{\alpha}\subseteq P$. Let $\Pi^{\notin P}$ be the set of simple roots outside $\Pi^{P}$. We say that $\lambda$ is $P$-dominant if $(\alpha^{\vee},\lambda)\geq 0$ for all $\alpha\in\Pi^{P}$. The set of $P$-dominant weights is denoted $X^{P}_{+}$. Every $W_{P}$-orbit of weights intersects $X^{P}_{+}$ in a unique element. Let $G_{P}$ be the semisimple subgroup of $P$ with $\Pi^{P}$ as simple roots. Note that $G_{P}$ is simply connected.
Lemma 13.8.
Let $v\in W^{P}$. Then $P(-e_{v})$ is a $P$-module and therefore $P(-e_{v})^{*}$ is a $P$-module. Conversely, if $e_{w}$ is $P$-dominant, then $w\in W^{P}$.
Proof.
By [47, Proposition 2.2.5], there is a sequence $\alpha_{1},\cdots,\alpha_{n}$ with $\alpha_{i}\in\Pi^{P}$ and ${\mathrm{ind}}_{B}^{P}={\mathrm{ind}}_{B}^{P_{\alpha_{1}}}\circ\cdots\circ{\mathrm{ind}}_{B}^{P_{\alpha_{n}}}$. It therefore suffices to see that ${\mathrm{ind}}_{B}^{P_{\alpha}}(P(-e_{v}))=P(-e_{v})$ for $\alpha\in\Pi^{P}$. Now $v$ is a minimal coset representative in $W/W_{P}$, so $vs_{\alpha}>v$ for $\alpha\in\Pi^{P}$. Then $s_{\alpha}v^{-1}>v^{-1}$, so $s_{\alpha}v^{-1}w_{0}<v^{-1}w_{0}$, and $s_{\alpha}\star v^{-1}w_{0}=v^{-1}w_{0}$. Thus ${\mathrm{ind}}_{B}^{P_{\alpha}}(P(-e_{v}))={\sf H}_{s_{\alpha}}{\sf H}_{v^{-1}w_{0}}(-w_{0}ve_{v})={\sf H}_{v^{-1}w_{0}}(-w_{0}ve_{v})=P(-e_{v})$.
Conversely, let $e_{w}$ be $P$-dominant and $\alpha\in\Pi^{P}$. Now $we_{w}$ is dominant and $w^{-1}$ is minimal amongst the $z$ with $zwe_{w}=e_{w}$. Thus if $s_{\alpha}e_{w}=e_{w}$ then $s_{\alpha}w^{-1}>w^{-1}$. If $s_{\alpha}e_{w}\neq e_{w}$ then $P(e_{w})$ is no $P_{\alpha}$-module, because its $B$-socle has weight $e_{w}$. Thus ${\mathrm{ind}}_{B}^{P_{\alpha}}P(e_{w})={\sf H}_{s_{\alpha}}{\sf H}_{w^{-1}}P(we_{w})$ with $s_{\alpha}\star w^{-1}\neq w^{-1}$ and $s_{\alpha}w^{-1}>w^{-1}$. We have $ws_{\alpha}>w$ in either case, so $w\in W^{P}$. ∎
Lemma 13.9.
Let $\lambda\in X^{P}_{+}$. Let $N\in{\mathrm{rep}}(B)$ be such that all weights $\mu$ of $N$ satisfy $\mu\leq_{a}\lambda$. Then all weights $\mu$ of ${\mathrm{ind}}_{B}^{P}(N)$ satisfy $\mu\leq_{a}\lambda$ and the natural map ${\mathrm{ind}}_{B}^{P}(N)\to N$ induces an isomorphism ${\mathrm{ind}}_{B}^{P}(N)_{\lambda}\to N_{\lambda}$.
Proof.
Recall there is a sequence $\alpha_{1},\cdots,\alpha_{n}$ with $\alpha_{i}\in\Pi^{P}$ and ${\mathrm{ind}}_{B}^{P}={\mathrm{ind}}_{B}^{P_{\alpha_{1}}}\circ\cdots\circ{\mathrm{ind}}_{B}^{P_{\alpha_{n}}}$. Therefore we may assume $P$ is a minimal parabolic, say $P=P_{\alpha}$. Thus $P/B$ is a projective line and we may use [29, II Proposition 5.2]. If $\mu$ is a weight and $(\alpha^{\vee},\mu)\geq 0$, then $s_{\alpha}\mu\leq_{a}\mu$, $\mathrm{R}^{1}{\mathrm{ind}}_{B}^{P}(\mu)=0$, $({\mathrm{ind}}_{B}^{P}(\mu))_{\mu}={\sf k}_{\mu}$ and all weights $\nu$ of ${\mathrm{ind}}_{B}^{P}(\mu)$ lie on the line segment joining $\mu$ with $s_{\alpha}\mu$. In particular, they satisfy $\nu\leq_{a}\mu$. And if $(\alpha^{\vee},\mu)<0$, then ${\mathrm{ind}}_{B}^{P}(\mu)=0$ and all weights $\nu$ of $\mathrm{R}^{1}{\mathrm{ind}}_{B}^{P}(\mu)=0$ lie strictly between the endpoints of the line segment. Thus $\nu<_{a}\mu$. If $N$ has just weight $\lambda$, then it is well known that ${\mathrm{ind}}_{B}^{P}(N)_{\lambda}\to N_{\lambda}$ is an isomorphism [23, Propositions 17, 20]. Now use induction on the number of weights of $N$. ∎
13.1 Generating ${\mathrm{rep}}(P)$
We say that a full subcategory of an abelian category has the 2 out of 3 property if, whenever $0\to N_{1}\to N_{2}\to N_{3}\to 0$ is exact and two of the $N_{i}$ are in the subcategory, then so is the third.
13.2 The set $\{\hat{M}_{v}\}_{v\in W^{P}}$
For $P$-dominant $\lambda$ we use ^$\hat{\nabla}_{\lambda}$ as another notation for ${\mathrm{ind}}_{B}^{P}{\sf k}_{\lambda}$. Let us be given a set $\{\hat{M}_{v}\}_{v\in W^{P}}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}(P)$ with the following properties. For each ${v\in W^{P}}$ there is a diagram of $P$-modules
$\hat{M}_{v}\stackrel{{\scriptstyle g}}{{\longrightarrow}}\hat{N}_{v}\stackrel{{\scriptstyle f}}{{\longleftarrow}}\hat{\nabla}_{e_{v}}$
such that every weight $\mu$ of ${\mathsf{ker}}(f)$, ${\mathsf{ker}}(g)$, ${\mathsf{coker}}(f)$, ${\mathsf{coker}}(g)$ satisfies $\mu<_{a}e_{v}$.
Examples of possible choices of $\hat{M}_{v}$ are ^$\hat{\nabla}_{{e_{v}}}$ with $f=g={\mathrm{id}}$, ${\mathrm{ind}}_{B}^{P}Q(e_{v})$ with $g={\mathrm{id}}$, $P(-e_{v})^{*}$ with $f={\mathrm{id}}$. The multiplicity of the weight $e_{v}$ is one in $\hat{M}_{v}$, $\hat{N}_{v}$, ^$\hat{\nabla}_{e_{v}}$. Every weight $\lambda$ of $\hat{M}_{v}$ satisfies $\lambda\leq_{a}e_{v}$.
Theorem 13.10 (Generation).
The smallest strictly full additive subcategory of ${\mathrm{rep}}(P)$ that
  • contains the $\hat{M}_{v}$,
  • has the 2 out of 3 property and
  • contains with every ^$\hat{\nabla}_{\lambda}$ also ^$\hat{\nabla}_{\lambda}\otimes\nabla_{\omega_{i}}$ for every fundamental representation $\nabla_{\omega_{i}}$,
is the category ${\mathrm{rep}}(P)$ of finite dimensional $P$-modules.
More specifically
Theorem 13.11 (Generation of initial interval).
Let $\tau\in X^{P}_{+}$. Let $V\in{\mathrm{rep}}(P)$ such that every weight $\mu$ of $V$ satisfies $\mu<_{a}\tau$. Then $V$ is an object of the smallest strictly full additive subcategory ${\mathcal{S}}$ of ${\mathrm{rep}}(P)$ satisfying
  • ${\mathcal{S}}$ contains $\hat{M}_{v}$ for $v\in W^{P}$ with $e_{v}<_{a}\tau$,
  • ${\mathcal{S}}$ has the 2 out of 3 property,
  • If ${\mathcal{S}}$ contains ^$\hat{\nabla}_{\lambda}$, then ${\mathcal{S}}$ also contains ^$\hat{\nabla}_{\lambda}\otimes\nabla_{\omega_{i}}$, for every fundamental representation $\nabla_{\omega_{i}}$.
Proof.
We argue by induction along the well-ordered partial order $\leq_{a}$. So we assume the Theorem when $\tau$ is replaced by a $\sigma\in X_{+}^{P}$ with $\sigma<_{a}\tau$. Let $\sigma\in X_{+}^{P}$ with $\sigma<_{a}\tau$. Thus $V\in{\mathcal{S}}$ when all weights $\xi$ of $V$ satisfy $\xi<_{a}\sigma$.
Step 1. First we wish to show that ${\mathcal{S}}$ contains at least one $P$-module $N$ with $N_{\sigma}={\sf k}_{\sigma}$ and with weights $\xi$ that satisfy $\xi\leq_{a}\sigma$. If $\sigma$ is a Steinberg weight $e_{v}$, then $v\in W^{P}$ by Lemma 13.8 and we simply take $N=\hat{M}_{v}$. If $\sigma$ is not a Steinberg weight, then we know from the proof Theorem 8.16 that there is a fundamental weight $\omega_{j}$ and a weight $\mu$ with $(\mu,\mu)<(\sigma,\sigma)$, so that the $B$-module $L=\nabla_{\omega_{j}}\otimes e^{\mu}$ has $\sigma$ as a weight of multiplicity one, and so that all weights $\nu$ of $L$ satisfy $\nu\leq_{a}\sigma$. Now take $N={\mathrm{ind}}_{B}^{P}L=\nabla_{\omega_{j}}\otimes{\mathrm{ind}}_{B}^{P}e^{\mu}$. By Lemma 13.9 every weight $\nu$ of $N$ satisfies $\nu\leq_{a}\sigma$ and $N_{\sigma}={\sf k}_{\sigma}$. In particular, ${\mathrm{ind}}_{B}^{P}e^{\mu}$ is nonzero, so $\mu\in X^{P}_{+}$ and ${\mathrm{ind}}_{B}^{P}e^{\mu}=\hat{\nabla}_{\mu}$. And every weight $\nu$ of ^$\hat{\nabla}_{\mu}$ satisfies $(\nu,\nu)<(\sigma,\sigma)$, hence $\nu<_{a}\sigma$, so ^$\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$ and $N=\nabla_{\omega_{j}}\otimes\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$.
Step 2. Next we wish to show that ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$. To this end we look for more $P$-modules $N$ in ${\mathcal{S}}$ with $N_{\sigma}={\sf k}_{\sigma}$ and such that the weights $\xi$ of $N$ satisfy $\xi\leq_{a}\sigma$. Start with the $N$ from Step 1. Let $N^{1}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin P}$. Then $N^{1}$ is a $P$-submodule that lies in ${\mathcal{S}}$. So we may replace $N$ with $N/N^{1}$ and further assume $N^{1}=0$. Let $N^{2}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin P}$. Then $N^{2}$ is a $P$-submodule and $N/N^{2}$ lies in ${\mathcal{S}}$. So we may replace $N$ with $N^{2}$ and further assume $N^{1}=0$, $N=N^{2}$. Let $N^{3}$ be the $P$-submodule of $N$ generated by $N_{\sigma}$. Then $N/N^{3}$ is in ${\mathcal{S}}$, so we may replace $N$ with $N^{3}$. Now the unipotent radical of $P$ acts trivially on $N$ and the projection $p$ of $N$ onto its weight space $N_{\sigma}$ is $T$-equivariant and $G_{P}$-equivariant, hence $P$-equivariant. By Lemma 13.9 the map $N\to\hat{\nabla}_{\sigma}$ induced by $p$ has kernel and cokernel in ${\mathcal{S}}$. So $\nabla_{\sigma}$ lies in ${\mathcal{S}}$.
Step 3. Finally we want to show that ${\mathcal{S}}$ contains every $P$-module $V$ all whose weights $\mu$ satisfy $\mu<_{a}\tau$. Consider such a $V$. We may and shall assume that ${\mathcal{S}}$ contains every $P$-module whose set of weights is a proper subset of the set of weights of $V$. Say $V$ is nonzero. Choose an extremal weight $\sigma$ of $V$ that is $P$-dominant. Let $V^{1}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin P}$. This is a $P$-submodule, and if $V^{1}$ is nonzero, then $V$ is in ${\mathcal{S}}$ because both $V^{1}$ and $V/V^{1}$ are. Let $V^{2}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin P}$. This is a $P$-submodule, and if $V^{2}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{2}$ and $V/V^{2}$ are. So we further assume $V^{1}=0$ and $V=V^{2}$. Then the unipotent radical of $P$ acts trivially on $V$. Let $V^{3}$ be the $P$-submodule generated by $V_{\sigma}$. If $V^{3}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{3}$ and $V/V^{3}$ are. So we may assume $V=V^{3}$. The weights $\xi$ of $V$ now satisfy $\xi\leq_{a}\sigma$. The projection $p$ of $V$ onto its weight space $V_{\sigma}$ is $T$-equivariant and $G_{P}$-equivariant, hence $P$-equivariant. We have ${\mathrm{ind}}_{B}^{P}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, where $(V_{\sigma})_{\mathrm{triv}}$ is $V_{\sigma}$ with trivial $P$-action, cf. [23, Proposition 17]. Consider the map $f:V\to{\mathrm{ind}}_{B}^{P}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, corresponding with $p$. Both ${\mathsf{ker}}(f)$ and ${\mathsf{coker}}(f)$ lie in ${\mathcal{S}}$. As ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$, so does ^$\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$. So $V$ lies in ${\mathcal{S}}$. ∎
For completeness we mention
Corollary 13.12.
Let $\lambda\in X_{+}^{P}$. Then [^$[\hat{\nabla}_{\lambda}]$ lies in the $R(G)$-submodule of $R(P)$ generated by the $[\hat{M}_{v}]$ with $v\in W^{P},\;e_{v}\leq_{a}\lambda$.
Proof.
By induction along $\leq_{a}$. Either $\lambda$ is a Steinberg weight or it is not …∎
Definition 24.
For $v\in W^{P}$ put $\hat{Q}(e_{v})={\mathrm{ind}}_{B}^{P}Q(e_{v})$. And then for $p\in W^{P}$
  • $\hat{\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succeq p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succ p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,v\in W^{P},\lambda\in X(T)_{+}})$
We get from Theorem 13.10 the following replacement of Theorem 8.15
Theorem 13.13.
Given a $p\in W^{P}$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(P))$ of the two categories $\hat{\sf P}_{\preceq p}$ and $\hat{\sf Q}_{\succ p}$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$.
We also have the following replacement of Theorem 6.6
Theorem 13.14.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $v,w\in W^{P}$.
  1. 1.
    $\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$,}\\ 0&\text{else.}\end{cases}$
  2. 2.
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{w}))=0$ for all $i$.
Proof.
By [46, Theorem 1.9.(a)(ii)] the module $Q(e_{w})$ is ${\mathrm{ind}}_{B}^{P}$ acyclic. We get
$\mathrm{R}^{j}{\mathrm{ind}}_{B}^{P}(P(-e_{v})\otimes Q(e_{w}))=P(-e_{v})\otimes\mathrm{R}^{j}{\mathrm{ind}}_{B}^{P}(Q(e_{w}))=\begin{cases}P(-e_{v})\otimes\hat{Q}(e_{w})&\text{if $j=0$,}\\ 0&\text{else.}\end{cases}$
So $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{w}))$. But the left hand side is known from Theorem 6.6. ∎
With these replacements we now proceed as in the case $P=B$: For $p\in W^{P}$, let $\hat{X}_{p}$ be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of $\hat{\sf Q}_{\succeq p}$ into $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. And let $\hat{Y}_{p}$ be the image of $\hat{Q}(e_{p})$ under the right adjoint of the inclusion of $\hat{\sf P}_{\preceq p}$ into $\mathcal{D}$.
Let $f_{p}$ be the natural map from $P(-e_{p})^{*}$ to $Q(e_{p})$. Recall that $f_{p}$ factors as a surjection $P(-e_{p})^{*}\to{\sf k}_{e_{p}}$, followed by an injection ${\sf k}_{e_{p}}\to Q(e_{p})$. Let $\hat{f}_{p}:P(-e_{p})^{*}\to\hat{Q}(e_{p})$ be induced by $f_{p}$. Then $\hat{f}_{p}$ induces an isomorphism of $T$-modules $(P(-e_{p})^{*})_{e_{p}}\to(\hat{Q}(e_{p}))_{e_{p}}$ and the weights $\mu$ of ${\mathsf{ker}}(\hat{f}_{p})$, ${\mathsf{coker}}(\hat{f}_{p})$ satisfy $\mu<_{a}e_{p}$. We have an exact triangle
$\dots\rightarrow\ker(\hat{f}_{p})[1]\to{\mathsf{cone}}(\hat{f}_{p})\to{\mathsf{coker}}(\hat{f}_{p})\rightarrow\dots$
in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Take $\hat{M}_{v}=Q(e_{v})$ for $p\prec v\in W^{P}$ and $\hat{M}_{v}=P(-e_{v})^{*}$ for $p\succ v\in W^{P}$. It does not matter what we choose for $\hat{M}_{p}$ itself.
Using Theorem 13.11 instead of Theorem 10.1, we find that ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\{\hat{M}_{v}\otimes\nabla_{\lambda}\mid v\in W^{P},\ e_{v}<_{a}e_{p},\ \lambda\in X(T)_{+}\}$. So ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\hat{\sf P}_{\prec p}\cup\hat{\sf Q}_{\succ p}$. One uses this to show that $\hat{X}_{p}$ and $\hat{Y}_{p}$ are isomorphic. Mutatis mutandis the old constructions and proofs go through and Theorems 13.2, 13.3, 13.4 follow.
13.3 A pairing
Let $X=G/B$, $Y=G/P$ and let $\pi:X\to Y$ be the projection. It induces a map
$\pi^{*}:K_{T}(Y)\to K_{T}(X)$ and therefore, by subsection 3.3, a pairing
$K_{T}(Y)\times K_{T}(X)\to R(T),\quad([{\mathcal{F}}],[{\mathcal{G}}])\mapsto\langle\pi^{*}[{\mathcal{F}}],[{\mathcal{G}}]\rangle.$
For $v,w\in W^{P}$ with $vw_{0}\not\leq ww_{0}$ the pairing sends $([\mathcal{L}_{Y}(P(-e_{v}))],[\mathcal{O}_{X^{ww_{0}}}(-\partial X^{ww_{0}})])$ to $\beta_{v,ww_{0}}=0$. And it sends $([\mathcal{O}_{\overline{B^{+}vP}/P}],[\mathcal{O}_{X^{ww_{0}}}(-\partial X^{ww_{0}})])$ to $\delta_{vw}$ for $v,w\in W^{P}$. If ${\sf k}={\mathbb{C}}$ it follows that $[\mathcal{L}_{Y}(P(-e_{v}))]$ is in the $R(T)$-span of the Schubert classes $[\mathcal{O}_{\overline{B^{+}wP}/P}]$ with $w\in W^{P}$, $w\leq v$.
14 Explicit calculations in rank 2
This section works out in detail the case of flag varieties of rank two groups. Unless specified otherwise, we work over $\mathbb{Z}$. The ensuing full exceptional collections consist of pure objects. These collections have the property that their duals in the sense of the definition of Section 7.8 are given by by applying the Grothendieck duality to the initial ones followed by tensoring with the square root of the canonical sheaf $\mathcal{L}(2\rho)$.
We first record several fundamental vanishing theorems that will be used throughout this section.
14.1 Borel-Weil-Bott theorem
Define
[…]
if ${\rm char}(k)=0$ and
[…]
if ${\rm char}(k)=p>0$.
Theorem 14.1.
[29, Part II, Corollary 5.5]
  • [(a)]
  • (a)
    If $\lambda\in\overline{C}_{\mathbb{Z}}$ with $\lambda\notin X({T})_{+}$, then
    ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})=0$
    for all $w\in W$. Here $\cdot$ is the dot–action of the Weyl group $W$ on $X({T})$.
  • (b)
    If $\lambda\in\overline{C}_{\mathbb{Z}}\cap X({T})_{+}$, then for all $w\in W$ and $i\in\mathbb{N}$
    ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})={\rm H}^{0}({G}/{B},\mathcal{L}_{\lambda}),$
    if $i=l(w)$ and otherwise ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})=0$. Here $l(w)$ is the length function.
14.2 Kempf vanishing theorem
Theorem 14.2.
[29, Part II, Proposition 4.5] and [18, Section 2.6] If $\lambda\in X({T})_{+}$, then
${\rm H}^{i}({G}/{B},\mathcal{L}_{\lambda})=0$
for all $i>0$.
By [3, Corollary 2.6, (ii)], ${\rm H}^{0}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})$ are free $\mathbb{Z}$-modules for $\lambda\in X({\mathbb{T}})_{+}$. Combined with the universal coefficient theorem [29, II, Section 8.7, (1)] and Theorem 14.2, this gives ${\rm H}^{i}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})=0$ for $i>0$.
14.3 Kodaira vanishing theorem
Theorem 14.3.
[7, Theorem 1.2.9] Let $k$ be a field of characteristic $p>0$ and $X$ be a Frobenius split $k$-variety. Assume given an ample line bundle $\mathcal{L}$ on $X$. Then $H^{i}(X,\mathcal{L}\otimes\omega_{X})=0$ for $i>0$.
Generalized flag varieties $G/P$ are Frobenius split, [7, Theorem 2.2.5], thus Kodaira vanishing holds for these.
Corollary 14.4.
Let $X$ be a Frobenius split Fano $k$-variety such that ${\rm Pic}(X)=\mathbb{Z}$ with $\mathcal{L}\in{\rm Pic}(X)$ being the positive generator. Let $n$ be the index of $X$, i.e. $\omega_{X}^{-1}=\mathcal{L}^{\otimes n}$. Then the collection of line bundles $\mathcal{O}_{X},\mathcal{L},\dots,\mathcal{L}^{\otimes(n-1)}$ is an exceptional collection on $X$.
Proof.
Setting $\mathcal{L}=\omega_{X}^{-1}$ in Theorem 14.3, we obtain $H^{i}(X,\mathcal{O}_{X})=0$ for $i>0$. It follows that for any line bundle $\mathcal{M}$ on $X$ one has ${\mathrm{Ext}}_{X}^{i}({\mathcal{M}},{\mathcal{M}})=H^{i}(X,\mathcal{O}_{X})=0$ for $i>0$, hence the collection $\mathcal{O}_{X},\mathcal{L},\dots,\mathcal{L}^{\otimes(n-1)}$ consists of exceptional bundles. The same theorem implies that ${\mathrm{Ext}}_{X}^{k}(\mathcal{L}^{\otimes j},\mathcal{L}^{\otimes i})=H^{k}(X,\mathcal{L}^{\otimes(i-j)}))=0$ for $i<j$ and $k>0$ since $-n+1\leq i-j<0$. Finally, $\mathop{\mathrm{Hom}}\nolimits_{X}(\mathcal{L}^{\otimes j},\mathcal{L}^{\otimes i})=H^{0}(X,\mathcal{L}^{\otimes(i-j)})=0$ since the bundle $\mathcal{L}^{\otimes(j-i)}$ is ample and hence has a non-trivial section. ∎
14.4 Demazure functors
By [29, Section 5.17-5.18], the Joseph-Demazure functors ${\sf H}_{w}$ from Section 2.2 can be related to functors which are defined geometrically. Let $R$ be the root system of a simple simply connected group $G$ of rank $r$ and $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}\subset R$ the simple roots. Let $s_{\alpha_{i}},i=1,\dots,r$ be the associated simple reflections of the Weyl group $W$. Given a simple root $\alpha$, let $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ denote the projection. Associated to $\alpha$ is an endofunctor ${\rm R}^{0}{\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\rm R}^{0}{\pi_{\alpha}}_{\ast}$ of ${\mathsf{QCoh}}^{G}({G}/{B})$ (the functor $\pi_{\alpha}^{\ast}$ is exact, the morphism $\pi_{\alpha}$ being smooth). Let $w$ be an element of the Weyl group. For a reduced expression of $w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{n}}$ consider the functor ${\rm R}^{0}{\sf D}_{w}={\rm R}^{0}{\sf D}_{\alpha_{1}}\circ\dots\circ{\rm R}^{0}{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{n}}}_{\ast}$. The functor ${\rm R}^{0}{\sf D}_{w}$ is also a left exact functor.
Similarly to the case of ${\sf H}_{w}$, let ${\sf D}_{w}={\sf D}_{\alpha_{1}}\circ\dots\circ{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\pi_{\alpha_{n}}}_{\ast}$ denote the right derived functor of ${\rm R}^{0}{\sf D}_{w}$; it is an endofunctor of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathsf{QCoh}}^{G}({G}/{B}))$ [^4] Note that the functor ${\sf D}_{w}$ is also defined on the derived category of non-equivariant coherent sheaves on $G/B$ and eventually will be considered on that category.. Since the morphisms $\pi_{\alpha_{i}}$’s are smooth and proper, the functors ${\sf D}_{w}$’s restrict to endofunctors of the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf Coh}^{G}({G}/{B}))$ [^5] The same footnote as above applies..
Now set ${G}:={B},X={G}$ (res., ${G}:={P}_{\alpha},X={G}$) in the definition of the sheafification functor $\mathcal{L}$ and denote $\mathcal{L}_{B}$ (resp., $\mathcal{L}_{{P}_{\alpha}}$) the corresponding functors ${{\mathrm{Rep}}}({B})\rightarrow{\mathsf{QCoh}}^{G}({G}/{B})$ (resp., ${{\mathrm{Rep}}}({P}_{\alpha})\rightarrow{\mathsf{QCoh}}^{G}({G}/{P}_{\alpha}$)).
Proposition 14.5.
Let $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))$ (resp., $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))$). Then there is an isomorphism $\mathcal{L}_{B}({\rm R}{\sf H}_{w}(M))={\sf D}_{w}\mathcal{L}_{B}(M)$.
Proof.
See [35, Lemma 5]. ∎
To unburden the notation, let $X$ denote ${G}/{B}$. Occasionally though when there is a need, we will stress the notation ${G}/{B}$. For a simple root $\alpha_{i}$ let ${P}_{\alpha_{i}}$ denote the minimal parabolic generated by ${B}$ and $U_{\alpha_{i}}$, and denote $Y_{i}:={G}/{P}_{\alpha_{i}}$ with $\pi_{\alpha_{i}}:X\rightarrow Y_{i}$. Given a simple root $\alpha_{i}$, we denote $\sf D_{i}$ the Demazure operator associated to $\alpha_{i}$, i.e. $\sf D_{i}:=\sf D_{\alpha_{i}}$. Throughout this section, we employ the notation of Section 1.5. The following theorem is the starting point for unraveling the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({B}))$ as a module category over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(G))$:
Theorem 14.6 (Demazure, [15]).
Let $w_{0}$ be the longest element of $W$. There is an isomorphism of functors ${\sf D}_{w_{0}}=\pi^{\ast}\pi_{\ast}=\sf D_{\mathcal{O}_{X}}$, where $\pi:X\rightarrow{\rm pt}$ is the projection to a point.
Proof.
Follows from rationality of Demazure resolutions applied to the longest word $w_{0}\in W$; more generally, ${d_{w}}_{\ast}\mathcal{O}_{Z_{w}}=\mathcal{O}_{X_{w}}$, where $d_{w}:Z_{w}\rightarrow X_{w}$ is a Demazure resolution of the Schubert variety $X_{w}$ and $w\in W$ is a reduced word, [11]. ∎
The functors ${\sf D}_{i}$ are idempotents: ${\sf D}_{i}\circ{\sf D}_{i}={\sf D}_{i}$. Indeed, by the projection formula one has ${\sf D}_{i}\circ{\sf D}_{i}={\sf D}_{i}\otimes{\sf D}_{i}\mathcal{O}_{X}={\sf D}_{i}$. This leads to the appearance of the 00–Hecke monoid. The following statement is a very simple consequence of that idempotency property that will be used throughout:
Proposition 14.7.
Let $w\in W$ and $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$ starting from $s_{\alpha_{i}}$. Let $\mathcal{E},\mathcal{F}$ be two objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ and ${\sf D}_{w}$ be the Demazure functor associated to $w$. Then
$\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{i}\mathcal{E},{\sf D}_{w}\mathcal{F})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{i}\mathcal{E},{\sf D}_{s_{\alpha_{i}}w}\mathcal{F}).$ (46)
Proof.
We have ${\sf D}_{w}={\sf D}_{i}{\sf D}_{s_{\alpha_{i}}w}$, hence $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{i}\mathcal{E},{\sf D}_{w}\mathcal{F})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{i}\mathcal{E},{\sf D}_{i}{\sf D}_{s_{\alpha_{i}}w}\mathcal{F})$. Now the functor ${\pi_{i}}^{\ast}$ is fully faithful, hence the latter group is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{Y_{i}}({\pi_{i}}_{\ast}\mathcal{E},{\pi_{i}}_{\ast}{\sf D}_{s_{\alpha_{i}}w}\mathcal{F})$. By adjunction $({\pi_{i}}^{\ast},{\pi_{i}}_{\ast})$ we see that the latter group is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{i}\mathcal{E},{\sf D}_{s_{\alpha_{i}}w}\mathcal{F})$. ∎
Let $\omega_{i}$ be a fundamental weight. The space of global sections $\Gamma(X,\mathcal{L}({\omega_{i}}))$ is isomorphic to $\nabla_{\omega_{i}}$ (the induced module). Let $\omega_{X}$ (resp., $\omega_{Y_{i}}$) denote the canonical line bundle on $X$ (resp., on $Y_{i}$). Then ${\pi_{\alpha_{i}}}_{\ast}\omega_{X}=\omega_{Y_{i}}[-1]$. The relative canonical sheaf $\omega_{X/Y_{i}}$ is isomorphic to $\mathcal{L}(-\alpha_{i})$ [16, Lemme 6], and hence ${\pi_{i}}_{\ast}\mathcal{L}({-\alpha_{i}})=\mathcal{O}_{Y_{i}}[-1]$. More generally, given an object $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, Grothendieck duality for the morphism ${\pi_{\alpha_{i}}}$ gives an isomorphism ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}^{\vee})^{\vee}$ where the inner $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{X}(\mathcal{F},\mathcal{O}_{X})$ on $X$ and the outer $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{Y_{i}}(\mathcal{F},\mathcal{O}_{Y_{i}})$ on $Y_{i}$. Since the pullback ${\pi_{\alpha_{i}}^{\ast}}$ commutes with $(-)^{\vee}$, we obtain an isomorphism ${\sf D}_{i}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\sf D}_{i}(\mathcal{F}^{\vee})^{\vee}$.
Given a line bundle $\mathcal{L}(\chi),\chi\in X({T})$ and a minimal parabolic ${P}_{\alpha}$ generated by $B$ and $U_{\alpha}$, the degree of the restriction of $\mathcal{L}(\chi)$ to fibres of the projection $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ is equal to $\langle\chi,\alpha^{\vee}\rangle$, [15, Lemme 2].
Given a fundamental weight $\omega_{i}$, there is a short exact sequence [16, Section 3]:
$0\rightarrow\mathcal{L}({\omega_{i}-\alpha_{i}})\rightarrow{\sf D}_{i}\mathcal{L}({\omega_{i}})\rightarrow\mathcal{L}({\omega_{i}})\rightarrow 0$ (47)
Proposition 14.8.
The right adjoint functor ${\sf D}_{i}^{\vee}(-)$ to ${\sf D}_{i}$ is isomorphic to ${\sf D}_{i}(-)\otimes\mathcal{L}(-{\alpha_{i}})[1]$.
Proof.
There are adjunction pairs of functors $({\pi_{i}^{\ast}},{\pi_{i}}_{\ast})$ and $({\pi_{i}}_{\ast},{\pi_{i}^{!}})$ is an adjoint pair. Now ${\sf D}_{i}={\pi_{i}^{\ast}}{\pi_{i}}_{\ast}$, hence its right adjoint is isomorphic to ${\sf D}_{i}^{\vee}={\pi_{i}^{!}}{\pi_{i}}_{\ast}$. The Grothendieck duality for the morphism $\pi_{i}:X\rightarrow Y_{i}$ gives an isomorphism ${\pi_{i}^{!}}(-)={\pi_{i}^{\ast}}(-)\otimes\omega_{X/Y_{i}}[1]$ where $\omega_{X/Y_{i}}$ is the relative canonical sheaf for the projection $\pi_{i}$ and the shift $[1]$ appears since the relative dimension $X$ over $Y_{i}$ is equal to $1$. Finally, recalling the isomorphism $\omega_{X/Y_{i}}=\mathcal{L}(-\alpha_{i})$ we obtain the result. ∎
14.5 Demazure resolutions
Let $w\in W$ and $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$. Associated to $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ is the Bott-Samelson-Demazure-Hansen resolution $Z_{w}$ of the Schubert variety $X_{w}$ associated to $w$, [7, Section 2.2] and [47, Section 1.2]. Both $X_{w}$ and $Z_{w}$ are acted upon by $B$ and one can form $G$-varieties by setting ${\mathcal{X}}_{w}=G\times^{B}X_{w}\subset X\times X$ and ${\mathcal{Z}}_{w}=G\times^{B}Z_{w}$. Then ${\mathcal{Z}}_{w}$ is a resolution of singularities of ${\mathcal{X}}_{w}$. Note that the variety ${\mathcal{Z}}_{w}$ depends on the choice of a reduced expression of $w$.
Given two smooth varieties $X$ and $Y$ over ${\sf k}$ and an object $\mathcal{P}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X\times Y)$, denote $\Phi_{\mathcal{P}}(-):={\pi_{Y}}_{\ast}(\pi_{X}^{\ast}(-)\otimes\mathcal{P})$ the integral transform between $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y)$, where $\pi_{X},\pi_{Y}$ are the projections of $X\times Y$ onto corresponding factors.
By the construction, the variety ${\mathcal{Z}}_{w}$ is a closed subvariety of $X^{l(w)}$. Denote $p_{w}$ and $q_{w}$ the two projections $Z_{w}$ onto the first and the last factor of $X^{l(w)}$, respectively.
Theorem 14.9.
[47, Corollary 2.2.7]
Let $w\in W$ and consider a reduced decomposition of $w$ as above. There is an isomorphism of endofunctors of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$:
${\sf D}_{w}=\Phi_{\mathcal{O}_{{\mathcal{X}}_{w}}}={q_{w}}_{\ast}p_{w}^{\ast}$ (48)
where $\Phi_{\mathcal{O}_{{\mathcal{X}}_{w}}}$ is the integral functor associated to $\mathcal{O}_{{\mathcal{X}}_{w}}$.
Proof.
Follows from the flat base change and rationality of singularities of Schubert varieties, [4]. ∎
14.6 Braid relations for Demazure functors
Let $(W,S=\{s_{\alpha_{1}},\dots,s_{\alpha_{r}}\})$ be the Coxeter presentation of the Weyl group of a simple simply connected group $G$ of rank $r$. The basic relations are $(s_{\alpha_{i}}s_{\alpha_{j}})^{m_{ij}}=1$ where $m_{ij}\in\{2,3,4,6\}$.
The braid relations hold for Demazure operators ${\sf D}_{i}$’s:
$\underbrace{{\sf D}_{i}{\sf D}_{j}{\sf D}_{i}\dots}_{m_{ij}}=\underbrace{{\sf D}_{j}{\sf D}_{i}{\sf D}_{j}\dots}_{m_{ij}}$ (49)
(i.e., each side is a product of $m_{ij}$ Demazure operators). Their validity follows from Theorem 14.9.
In type ${\sf A}_{n}$ all the exponents $m_{ij}\leq 3$ and the braid relations simplify to
${\sf D}_{i}{\sf D}_{j}{\sf D}_{i}={\sf D}_{j}{\sf D}_{i}{\sf D}_{j}$ (50)
if $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=-1$; otherwise ${\sf D}_{i}$ and ${\sf D}_{j}$ commute:
${\sf D}_{i}{\sf D}_{j}={\sf D}_{j}{\sf D}_{i}$ (51)
if $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=0$.
14.7 The bundle $\mathcal{L}(-\rho)$
Recall that the canonical sheaf $\omega_{X}$ is isomorphic to $\mathcal{L}(-2\rho)$ where $\rho$ is the sum of fundamental weights. The following proposition will turn to be useful in checking semiorthogonality properties of exceptional sequences.
Proposition 14.10.
Let $\mathcal{E}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ and $w\in W,l(w)\geq 1$ with the associated Demazure operator ${\sf D}_{w}$. Then $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{w}\mathcal{E},\mathcal{L}(-\rho))=0$.
Proof.
Let $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$. Then ${\sf D}_{w}={\sf D}_{i_{1}}{\sf D}_{i_{2}}\dots\sf D_{i_{k}}={\sf D}_{i_{1}}{\sf D}_{s_{\alpha_{1}}w}$. By adjunction $({\pi_{i}}^{\ast},{\pi_{i}}_{\ast})$, we obtain
$\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{w}\mathcal{E},\mathcal{L}(-\rho))=\mathop{\mathrm{RHom}}\nolimits_{Y_{1}}({\pi_{i_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\mathcal{E},{\pi_{i}}_{\ast}\mathcal{L}(-\rho)).$ (52)
Now $\mathop{\mathrm{RHom}}\nolimits_{Y_{1}}({\pi_{i_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\mathcal{E},{\pi_{i}}_{\ast}\mathcal{L}(-\rho))=0$ since ${\pi_{i}}_{\ast}\mathcal{L}(-\rho)=0$ as $\langle-\rho,\alpha_{i}^{\vee}\rangle=-1$. ∎
14.8 An exact triangle
Let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$, and consider the associated projective bundle $\pi:\mathbb{P}(\mathcal{E})\rightarrow X$. Denote $\mathcal{O}_{\pi}(-1)$ the line bundle on $\mathbb{P}(\mathcal{E})$ of relative degree $-1$, such that $\pi_{\ast}\mathcal{O}_{\pi}(1)=\mathcal{E}^{\ast}$. One has [26, Corollary 8.36]:
Theorem 14.11.
The category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbb{P}(\mathcal{E}))$ has a semiorthogonal decomposition:
$\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbb{P}(\mathcal{E}))=\langle\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\otimes\mathcal{O}_{\pi}(-r+1),\dots,\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\otimes\mathcal{O}_{\pi}(-1),\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\rangle.$ (53)
The following proposition generalizes the sequence (47) for an arbitrary object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$:
Proposition 14.12.
For any object $\mathcal{E}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ there is an exact triangle
$\dots\longrightarrow{\sf D}_{i}(\mathcal{E}\otimes\mathcal{L}(-\omega_{i}))\otimes\mathcal{L}(\omega_{i}-\alpha_{i})\longrightarrow{\sf D}_{i}\mathcal{E}\rightarrow\mathcal{E}\stackrel{{\scriptstyle[1]}}{{\longrightarrow}}\dots$ (54)
Proof.
The triangle is obtained by Theorem 14.11. Note that the cone of the canonical adjunction morphism ${\sf D}_{i}\mathcal{E}\rightarrow\mathcal{E}$ is of the form $\pi_{i}^{\ast}(?)\otimes\mathcal{L}(\omega_{i}-\alpha_{i})$. To calculate the unknown object ???? in the cone, tensor the adjunction map with $\mathcal{L}(-\omega_{i})$ and apply ${\pi_{i}}_{\ast}$. One obtains ${\pi_{i}}_{\ast}{\rm Cone}({\sf D}_{i}\mathcal{E}\rightarrow\mathcal{E})=?[-1]={\pi_{i}}_{\ast}(\mathcal{E}\otimes\mathcal{L}(-\omega_{i}))$ giving the triangle (54). ∎
Remark 14.13.
Recall the formula for the functor $Q(-)$ (Section 2.3 and [47, Definition 7.2.1]):
$Q:={{\mathrm{res}}}_{B}^{P_{\alpha_{1}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{B}^{P_{\alpha_{1}}}({\mathbb{Z}}_{s_{\alpha_{1}}(\rho)}\otimes_{\mathbb{Z}}(-))\dots{\mathrm{res}}_{B}^{P_{\alpha_{n}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{B}^{P_{\alpha_{n}}}({\mathbb{Z}}_{s_{\alpha_{n}}(\rho)}\otimes_{\mathbb{Z}}(-))$ (55)
The cone ${\sf D}_{i}(-\otimes\mathcal{L}(-\omega_{i}))\otimes\mathcal{L}(\omega_{i}-\alpha_{i}))$ in (54) of Proposition 14.12 is the derived version of the factor ${{\mathrm{res}}}_{\bf B}^{{\bf P}_{\alpha_{i}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{\bf B}^{{\bf P}_{\alpha_{1}}}({\mathbb{Z}}_{s_{\alpha_{i}}(\rho)}\otimes_{\mathbb{Z}}(-))$ of (55): note that $\langle s_{\alpha_{i}}(\rho),\alpha_{i}^{\vee}\rangle=\langle-\omega_{i},\alpha_{i}^{\vee}\rangle=-1$ and $\langle-\rho,\alpha_{i}^{\vee}\rangle=\langle\omega_{i}-\alpha_{i},\alpha_{i}^{\vee}\rangle=-1$. Iterating the cones, we get a right derived functor of $Q$.
14.9 More on Demazure functors
We collect here various auxiliary statements about the effect of Demazure functors on objects of interest.
Let $G$ be a simple simply connected algebraic group and $\omega_{1},\dots,\omega_{r}$ be the fundamental weights of ${G}$. Given a fundamental weight $\omega_{i}$, define the vector bundle $\Psi_{1}^{\omega_{i}}$ as the kernel of the canonical surjective morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow\mathcal{L}(\omega_{i})$. Thus, there is the short exact sequence:
$0\rightarrow\Psi_{1}^{\omega_{i}}\rightarrow\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow\mathcal{L}(\omega_{i})\rightarrow 0,$ (56)
and $\Psi_{1}^{\omega_{i}}$ is the left mutation of $\mathcal{L}(\omega_{i})$ through $\mathcal{L}(0)$. For referential purposes, we also record the dual sequence
$0\rightarrow\mathcal{L}(-\omega_{i})\rightarrow(\nabla_{\omega_{i}})^{\ast}\otimes\mathcal{L}(0)\rightarrow(\Psi_{1}^{\omega_{i}})^{\ast}\rightarrow 0$ (57)
We can now state a few simple propositions describing the action of elementary Demazure operators ${\sf D}_{k}$.
Proposition 14.14.
${\sf D}_{k}(\Psi_{1}^{\omega_{l}})^{\ast}=\left\{\begin{array}[]{l}(\nabla_{\omega_{l}})^{\ast}\otimes\mathcal{L}(0),\quad k=l,\\ \\ (\Psi_{1}^{\omega_{l}})^{\ast},\quad k\neq l.\\ \end{array}\right.$ (58)
Proof.
Consider the sequence (57) for $i=l$ and apply to it the functor ${\pi_{\alpha_{k}}}_{\ast}$. Now observe that $\langle-\omega_{l},\alpha_{k}^{\vee}\rangle=-\delta_{kl}$, hence ${\sf D}_{k}\mathcal{L}(-\omega_{k})=0$ and ${\sf D}_{k}\mathcal{L}(-\omega_{l})=\mathcal{L}(-\omega_{k})$ if $l\neq k$.
The middle term of (57) is the trivial bundle, hence remains unaffected by ${\sf D}_{k}$, and the statement follows. ∎
Proposition 14.15.
Let the Dynkin type of $G$ be ${\sf A}_{n}$. Then
${\sf D}_{k}((\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}))=\left\{\begin{array}[]{l}\mathcal{L}(-\omega_{k-1}-\omega_{k+1}),\quad k=l,\\ \\ (\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}),\quad k\neq l.\\ \end{array}\right.$ (59)
The weight $\omega_{k-1}$ (resp., $\omega_{k+1}$) is set to be 00 if $k=1$ (resp., $k=n$).
Proof.
Tensoring the sequence (57) for $i=l$ with $\mathcal{L}(-\omega_{l})$, we obtain
$0\rightarrow\mathcal{L}(-2\omega_{l})\rightarrow(\nabla_{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l})\rightarrow(\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l})\rightarrow 0$ (60)
Applying the functor ${\pi_{\alpha_{k}}}_{\ast}$ to it and using once again the equality $\langle-\omega_{l},\alpha_{k}^{\vee}\rangle=-\delta_{kl}$, we see that the sequence is unaffected by ${\pi_{\alpha_{k}}}_{\ast}$ if $k\neq l$ which asserts the second line of (59). If $k=l$ then ${\pi_{\alpha_{l}}}_{\ast}\mathcal{L}(-\omega_{l})=0$ and we need to compute ${\pi_{\alpha_{l}}}_{\ast}\mathcal{L}(-2\omega_{l})$. Now $-2\omega_{l}=-\alpha_{l}-\omega_{l-1}-\omega_{l+1}$ (the Cartan matrix of ${\sf A}_{n}$) where $\omega_{l-1}$ (resp., $\omega_{l+1}$) is set to be 00 if $l=1$ (resp., $l=n$). Using the isomorphism ${\pi_{\alpha_{l}}}_{\ast}\mathcal{L}(-\alpha_{l})=\mathcal{O}_{Y_{l}}[-1]$ and projection formula, we obtain an isomorphism ${\sf D}_{l}\mathcal{L}(-2\omega_{l})=\mathcal{L}(-\omega_{l-1}-\omega_{l+1})[-1]$. Now the exact triangle obtained by applying ${\pi_{\alpha_{l}}}_{\ast}$ to (60) gives the statement. ∎
Remark 14.16.
The above Proposition 14.15 can easily be generalized to all simple groups, but the formulae for ${\sf D}_{k}((\Psi_{1}^{\omega_{k}})^{\ast}\otimes\mathcal{L}(-\omega_{k}))$ will then depend on the fundamental weight $\omega_{k}$. We record the examples of ${B}_{2}$ and of ${G}_{2}$ that we will need later. Let $\alpha,\beta$ be the simple roots with $\alpha$ being the short root. Then if $G$ is of type ${B}_{2}$, we have:
${\sf D}_{k}((\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}))=\left\{\begin{array}[]{l}\mathcal{L}(-\omega_{\beta}),\quad k=l,\quad\alpha_{k}=\alpha,\\ \\ \mathcal{L}(-2\omega_{\alpha}),\quad k=l,\quad\alpha_{k}=\beta,\\ \\ (\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}),\quad k\neq l.\\ \end{array}\right.$ (61)
If $G$ is of type ${G}_{2}$, we have
${\sf D}_{k}((\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}))=\left\{\begin{array}[]{l}\mathcal{L}(-\omega_{\beta}),\quad k=l,\quad\alpha_{k}=\alpha,\\ \\ \mathcal{L}(-3\omega_{\alpha}),\quad k=l,\quad\alpha_{k}=\beta,\\ \\ (\Psi_{1}^{\omega_{l}})^{\ast}\otimes\mathcal{L}(-\omega_{l}),\quad k\neq l.\\ \end{array}\right.$ (62)
Proposition 14.17.
${\sf D}_{k}(\Psi_{1}^{\omega_{l}}\otimes\mathcal{L}(-\omega_{l}))=\left\{\begin{array}[]{l}\mathcal{L}(0)[-1],\quad k=l,\\ \\ \Psi_{1}^{\omega_{l}}\otimes\mathcal{L}(-\omega_{l}),\quad k\neq l.\\ \end{array}\right.$ (63)
Proof.
Tensoring the sequence (56) for $i=l$ with $\mathcal{L}(-\omega_{l})$ and applying ${\sf D}_{k}$ to the result, and taking into account $\langle-\omega_{l},\alpha_{k}^{\vee}\rangle=-\delta_{kl}$, the ensuing exact triangle gives the result. ∎
Proposition 14.18.
The functors ${\sf D}_{w}$ preserve the subcategory $\langle\mathcal{O}_{X}\rangle^{\perp}$.
Proof.
Let $\mathcal{F}$ be an object of $\langle\mathcal{O}_{X}\rangle^{\perp}$, i.e. $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},\mathcal{F})=0$. Let $w\in W$ and $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$. We need to show that $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},{\sf D}_{w}\mathcal{F})=0$. We argue by induction on $l(w)$. The case $l(w)=0$ holds by the assumption. Let $l(w)>0$. Setting $\mathcal{E}=\mathcal{L}(0)=\mathcal{O}_{X}$ in the equation (46) of Proposition 14.7, we obtain an isomorphism $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},{\sf D}_{w}\mathcal{F})=\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},{\sf D}_{s_{\alpha_{1}}w}\mathcal{F})$. Now $l(s_{\alpha_{1}}w)<l(w)$ and we conclude by induction on $l(w)$. ∎
By Theorems 14.1 and 14.3, the collection of line bundles $\mathcal{L}(\omega_{1}),\dots,\mathcal{L}(\omega_{k})$ is an exceptional collection on $X$ the terms of which are completely orthogonal to each other. By Theorem 14.1, $\mathop{\mathrm{RHom}}\nolimits_{X}(\bigoplus_{i=1}^{i=r}\mathcal{L}(\omega_{i}),\mathcal{O}_{X})=0$ and the admissible subcategories $\langle\mathcal{O}_{X}\rangle$ and $\langle\mathcal{L}(\omega_{1}),\dots,\mathcal{L}(\omega_{k})\rangle$ form a semiorthogonal sequence in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$.
Proposition 14.19.
The functors ${\sf D}_{w}$ preserve the subcategory $\langle\mathcal{O}_{X},\bigoplus_{i=1}^{i=r}\mathcal{L}(\omega_{i})\rangle^{\perp}$.
Proof.
Let $\mathcal{E}$ be an object of $\langle\mathcal{O}_{X},\bigoplus_{i=1}^{i=r}\mathcal{L}(\omega_{i})\rangle^{\perp}$, i.e. $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},\mathcal{E})=0$ and $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{i}),\mathcal{E})=0$ for $i=1,\dots,r$. It is sufficient to prove the statement for a simple reflection $w=s_{k}$. Proposition 14.18 reduces the statement to verifying $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{i}),{\sf D}_{k}\mathcal{E})=0$ for $i=1,\dots,r$. Let first $k=i$; then $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{k}),{\sf D}_{k}\mathcal{E})=\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{O}_{X},{\sf D}_{k}\mathcal{E}\otimes\mathcal{L}(-\omega_{k}))=0$ as ${\pi_{k}}_{\ast}\mathcal{L}(-\omega_{k})=0$. By Proposition 14.12, for $k\neq i$ consider the exact triangle
$\dots\rightarrow{\sf D}_{k}(\mathcal{E}\otimes\mathcal{L}(-\omega_{k}))\otimes\mathcal{L}(\omega_{k}-\alpha_{k})\rightarrow{\sf D}_{k}\mathcal{E}\rightarrow\mathcal{E}\rightarrow\dots$
By the assumption, the required orthogonality will follow from $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{i}),{\sf D}_{k}(\mathcal{E}\otimes\mathcal{L}(-\omega_{k}))\otimes\mathcal{L}(\omega_{k}-\alpha_{k}))=0$. The latter equality follows from ${\pi_{k}}_{\ast}\mathcal{L}(\omega_{k}-\alpha_{k}-\omega_{i})=0$ as $\langle\omega_{k}-\alpha_{k}-\omega_{i},\alpha_{k}^{\vee}\rangle=-1$. We conclude by induction on $l(w)$ as in Proposition 14.18. ∎
Corollary 14.20.
The functors ${\sf D}_{w}$ preserve the subcategory $\langle\bigoplus_{i=1}^{i=r}\Psi_{1}^{\omega_{i}},\mathcal{O}_{X}\rangle^{\perp}$
Proof.
The categories $\langle\mathcal{O}_{X},\bigoplus_{i=1}^{i=r}\mathcal{L}(\omega_{i})\rangle$ and $\langle\bigoplus_{i=1}^{i=r}\Psi_{1}^{\omega_{i}},\mathcal{O}_{X}\rangle$ coincide, hence so do their right orthogonals. ∎
Proposition 14.21.
Let $w\in W$ and ${\sf D}_{w}$ the associated Demazure functor. Then ${\sf D}_{w}\Psi_{1}^{\omega_{i}}$ is a vector bundle for any $i=1,\dots,r$.
Proof.
Applying ${\sf D}_{w}$ to (57), we obtain the exact triangle
$\dots\rightarrow{\sf D}_{w}\Psi_{1}^{\omega_{i}}\rightarrow\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{w}\mathcal{L}(\omega_{i})\rightarrow\dots$ (64)
Consider ${\sf D}_{w}\mathcal{L}(\omega_{i})$ which a priori belongs to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{\geq 0}(X)$, the functor ${\sf D}_{w}$ being a right derived functor. In fact, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is a coherent sheaf, i.e. a pure object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$: this follows from Definition 1, Proposition 14.5 and the higher cohomology vanishing ${\rm H}^{k}(X_{w},\mathcal{L}(\omega_{i}))=0$ for $k>0$. Moreover, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is locally free, being a ${G}$-equivariant coherent sheaf on $X$.
The space of global sections of $\mathcal{L}(\omega_{i})$ is also isomorphic to $\nabla_{\omega_{i}}$ and remains to show that the morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{w}\mathcal{L}(\omega_{i})$ is surjective; this also follows from the surjectivity of the above restriction map on the global sections $\nabla_{\omega_{i}}\rightarrow{\rm H}^{0}(X_{w},\mathcal{L}(\omega_{i}))$ (cf. [47, Proposition A.2.6]) and applying Proposition 14.5 to that surjection. Now ${\sf D}_{w}\Psi_{1}^{\omega_{i}}$ is the kernel of a surjective map of vector bundles, hence is also a vector bundle. ∎
Proposition 14.22.
The vector bundle ${\sf D}_{i}\Psi_{1}^{\omega_{i}}$ fits into the following short exact sequence:
$0\rightarrow{\sf D}_{i}\Psi_{1}^{\omega_{i}}\rightarrow\Psi_{1}^{\omega_{i}}\rightarrow\mathcal{L}(\omega_{i}-\alpha_{i})\rightarrow 0$ (65)
Proof.
Writing the exact triangle (54) from Proposition 14.12 for $\Psi_{1}^{\omega_{i}}$, we obtain
$\dots\rightarrow{\sf D}_{i}(\Psi_{1}^{\omega_{i}}\otimes\mathcal{L}(-\omega_{i}))\otimes\mathcal{L}(\omega_{i}-\alpha_{i})\rightarrow{\sf D}_{i}\Psi_{1}^{\omega_{i}}\rightarrow\Psi_{1}^{\omega_{i}}\rightarrow\dots$ (66)
Proposition 14.17 computes ${\sf D}_{i}(\Psi_{1}^{\omega_{i}}\otimes\mathcal{L}(-\omega_{i}))\otimes\mathcal{L}(\omega_{i}-\alpha_{i})$ to be isomorphic to $\mathcal{L}(\omega_{i}-\alpha_{i})[-1]$, so the above exact triangle gives a short exact sequence.
$0\rightarrow{\sf D}_{i}\Psi_{1}^{\omega_{i}}\rightarrow\Psi_{1}^{\omega_{i}}\rightarrow\mathcal{L}(\omega_{i}-\alpha_{i})\rightarrow 0$ (67)
The map $\Psi_{1}^{\omega_{i}}\rightarrow\mathcal{L}(\omega_{i}-\alpha_{i})$ is given by an isomorphism $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},\mathcal{L}(\omega_{i}-\alpha_{i}))=\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-\alpha_{i})[1])={\sf k}$ as can be seen from the long exact cohomology sequence associated to the tensor product of sequence (57) with $\mathcal{L}(\omega_{i}-\alpha_{i})$. ∎
Proposition 14.23.
Let $\omega_{i},\omega_{j}$ be two fundamental weights and $w\in W$. Then $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$ if $l(w)>0$ and a reduced decomposition of $w$ contains $s_{\alpha_{j}}$ or $i\neq j$ and $l(w)\geq 0$.
Proof.
We proceed by induction on the length $l(w)$ of $w\in W$. We assume first $i\neq j$, so the base of induction is $l(w)=0$. Then ${\sf D}_{w}={\sf id}$ and we need to show $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},\Psi_{1}^{\omega_{j}})=0$. The pair $\Psi_{1}^{\omega_{i}},\Psi_{1}^{\omega_{j}}$ is the left mutation of the pair $\mathcal{L}(\omega_{i}),\mathcal{L}(\omega_{j})$ through $\mathcal{L}(0)$ and the line bundles in the latter pair are mutually orthogonal, hence the bundles $\Psi_{1}^{\omega_{i}},\Psi_{1}^{\omega_{j}}$ are also mutually orthogonal.
Let $l(w)>0$. There is an isomorphism
$\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},{\sf D}_{w}\Psi_{1}^{\omega_{j}})=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{w}\Psi_{1}^{\omega_{j}}\otimes(\Psi_{1}^{\omega_{i}})^{\ast})$ (68)
The base of induction is $l(w)=1$, thus $w=s_{\alpha_{l}}$ for a simple root $\alpha_{l}$ and by the assumption on $w$ we must have $l=j$. If $i=j$ then by projection formula we obtain $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{i}\Psi_{1}^{\omega_{i}}\otimes(\Psi_{1}^{\omega_{i}})^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{i},{\pi_{\alpha_{i}}}_{\ast}\Psi_{1}^{\omega_{i}}\otimes{\pi_{\alpha_{i}}}_{\ast}(\Psi_{1}^{\omega_{i}})^{\ast})$ and by Proposition 14.14 the last group is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{i}}}_{\ast}\Psi_{1}^{\omega_{i}})\otimes(\nabla_{\omega_{i}})^{\ast}=\mathop{\mathrm{R\Gamma}}\nolimits(X,\Psi_{1}^{\omega_{i}})\otimes(\nabla_{\omega_{i}})^{\ast}=0$. If $i\neq j$ then $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},{\sf D}_{j}\Psi_{1}^{\omega_{j}})=\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},\Psi_{1}^{\omega_{j}})=0$ where the last isomorphism follows from the previous paragraph and the first isomorphism is explained in detail below in the equation (71).
Let $w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{k}},k=l(w)>0$ be a reduced expression of $w\in W$, so ${\sf D}_{w}={\sf D}_{1}{\sf D}_{s_{\alpha_{1}}w}$. Then by projection formula
$\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{w}\Psi_{1}^{\omega_{j}}\otimes(\Psi_{1}^{\omega_{i}})^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}}\otimes{\pi_{\alpha_{1}}}_{\ast}(\Psi_{1}^{\omega_{i}})^{\ast}).$ (69)
Assume first $i=1$. Then using Proposition 14.14, we obtain
$\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}}\otimes{\pi_{\alpha_{1}}}_{\ast}(\Psi_{1}^{\omega_{1}})^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}})\otimes(\nabla_{\omega_{i}})^{\ast}.$ (70)
Now $\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}})=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$ where the last isomorphism follows from Proposition 14.18.
Assume $i\neq 1$, then using again Proposition 14.14, we obtain
$\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}}\otimes{\pi_{\alpha_{1}}}_{\ast}(\Psi_{1}^{\omega_{i}})^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{1},{\pi_{\alpha_{1}}}_{\ast}{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}}\otimes(\Psi_{1}^{\omega_{i}})^{\ast})=$ (71)
$\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}}).$
Now $l(s_{\alpha_{1}}w)<l(w)$ and by the induction assumption $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}},{\sf D}_{s_{\alpha_{1}}w}\Psi_{1}^{\omega_{j}})=0$. ∎
Proposition 14.24.
Let $\omega_{i},\omega_{j}$ be two distinct fundamental weights of $G$. Then
$\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{i}),{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$
Proof.
Consider the sequence (56) and apply to it $\mathop{\mathrm{RHom}}\nolimits_{X}(-,{\sf D}_{w}\Psi_{1}^{\omega_{j}})$. We have $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$ by Proposition 14.18 and $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{i}},{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$ by the previous Proposition 14.23. Hence, $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{i}),{\sf D}_{w}\Psi_{1}^{\omega_{j}})=0$. ∎
Proposition 14.25.
The rank two vector bundles ${\sf D}_{i}\mathcal{L}(\omega_{i}),i=1,\dots,r$ are exceptional, except for the two exceptional cases (see the footnote).
Proof.
[…]
Corollary 14.26.
In the assumption of the previous proposition, the bundles ${\sf D}_{i}\Psi_{1}^{\omega_{i}},i=1,\dots,r$ are exceptional.
Proof.
Recall that $\Psi_{1}^{\omega_{i}},i=1,\dots,r$ are defined to be the left mutations of $\mathcal{L}(\omega_{i})$ through $\mathcal{O}_{X}$. Hence ${\sf D}_{i}\Psi_{1}^{\omega_{i}}$ is the cokernel of the evaluation morphism $\nabla_{\omega_{i}}\otimes\mathcal{O}_{X}\rightarrow{\sf D}_{i}\mathcal{L}(\omega_{i})$. By the previous Proposition 14.25, the bundles ${\sf D}_{i}\mathcal{L}(\omega_{i}),i=1,\dots,r$ are exceptional, hence so are ${\sf D}_{i}\Psi_{1}^{\omega_{i}}$. ∎
Proposition 14.27.
Let the rank of group $G$ be equal to 2, and let ${P}_{\alpha_{i}}$ denote the minimal parabolic generated by ${B}$ and $U_{\alpha_{i}}$ for a simple root $\alpha_{i},i=1,2$. Consider the subcategory $\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1})\diamond\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2})$, i.e. the minimal thick strict full triangulated subcategory containing both subcategories $\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$ for $i=1,2$. Then
$\langle\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1})\diamond\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2})\rangle={{}^{\perp}\langle\mathcal{L}(-\rho)\rangle}.$ (72)
Proof.
Note that the subcategory $\langle\mathcal{L}(-\rho)\rangle$ is admissible by Theorem 14.3 and Section 7.6. We will prove an equivalent statement:
$^{\perp}\langle(\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1}))\otimes\mathcal{L}(-\rho)\diamond(\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2}))\otimes\mathcal{L}(-\rho)\rangle=\langle\mathcal{L}(0)\rangle.$ (73)
Indeed, having established the above equality, we obtain (72) by tensoring (73) with $\mathcal{L}(\rho)$ and mutating the subcategory $\langle\mathcal{L}(\rho)\rangle$ to the left past its right orthogonal. By [32, Lemma 2.11], the mutation is given by the Serre functor which is isomorphic to tensoring with $\omega_{X}=\mathcal{L}(-2\rho)$.
[…]
$\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{G},\pi_{\alpha_{i}}^{\ast}{\pi_{\alpha_{i}}}_{\ast}\mathcal{F}\otimes\mathcal{L}({-\rho}))=\mathop{\mathrm{RHom}}\nolimits_{Y_{i}}({\pi_{\alpha_{i}}}_{\ast}\mathcal{F},{\pi_{\alpha_{i}}}_{\ast}(\mathcal{G}\otimes\mathcal{L}({-\rho}))[{\rm dim}(X)])^{\ast}=0.$ (74)
Since $\mathcal{F}$ is arbitrary and the functor ${\pi_{\alpha_{i}}}_{\ast}$ is surjective, it follows that ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{G}\otimes\mathcal{L}({-\rho}))=0$. This implies $\mathcal{G}\in\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$, the line bundle $\mathcal{L}({-\rho})$ having degree $-1$ along $\pi_{\alpha}$. It follows that $\sf D_{\alpha_{i}}\mathcal{G}=\mathcal{G}$ for $i=1,2$ for such an object $\mathcal{G}$. Choosing a reduced expression for the longest element $w_{0}$, we obtain $\sf D_{w_{0}}\mathcal{G}=\mathcal{G}$, and hence $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{G}$, since $\sf D_{w_{0}}=\pi^{\ast}\pi_{\ast}$ by Theorem 14.6. Now $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{G})\in\langle\mathcal{L}(0)\rangle$, and the statement follows. ∎
Corollary 14.28.
Let the rank of group $G$ be equal to 2, and $X$ the flag variety of $G$. Then $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ has a semiorthogonal decomposition:
$\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)=\langle\langle\mathcal{L}(-\rho)\rangle,\langle\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1})\diamond\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2})\rangle\rangle.$ (75)
Remark 14.29.
Proposition 14.27 and Corollary 14.28 hold more generally for groups of arbitrary rank.
14.10 Type ${\bf A}_{2}$
Let $\omega_{1}$ and $\omega_{2}$ be the two fundamental weights for ${\bf SL}_{3}$ and $\alpha_{1},\alpha_{2}$ the simple roots. As before, set $X:={\bf SL}_{3}/{B}$ and $Y_{i}:={\bf SL}_{3}/{P}_{\alpha_{i}},i=1,2$.
Lemma 14.30.
The following block collection is a full strong exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$:
[…]
$\mathcal{L}(-\rho)$
${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$
${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})$
$\Psi_{1}^{\omega_{1}}$
$\Psi_{1}^{\omega_{2}}$
$\mathcal{L}(0)$
That is, ${\sf A}_{i}\subset{{}^{\perp}{\sf A}_{j}}$ for $-3\leq j<i\leq 0$, each of the vector bundles is exceptional, the vector bundles in a box are completely orthogonal, and the triangulated hull of $\langle{\sf A}_{-3},\dots,{\sf A}_{0}\rangle$ is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$.
Proof.
Proposition 14.18 and Theorem 14.1 give ${\sf A}_{0}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=1,2,3$. Proposition 14.23 gives ${\sf A}_{-1}\subset{{}^{\perp}{\sf A}_{-2}}$. Finally, Proposition 14.10 gives ${\sf A}_{-i}\subset{{}^{\perp}{\sf A}_{-3}}$ for $i=0,1,2$. This proves that the sequence $\langle{\sf A}_{-3},\dots,{\sf A}_{0}\rangle$ is semiorthogonal.
By Corollary 14.4, the terms outside the block ${\sf A}_{-2}$ are exceptional: these are either line bundles or the bundles in the block ${\sf A}_{-1}$ which are obtained by the left mutation of the block $\mathcal{L}(\omega_{1}),\mathcal{L}(\omega_{2})$ through $\mathcal{L}(0)$; since the latter block is exceptional, then so is the mutated block ${\sf A}_{-1}$. Proposition 14.25 states that the bundles ${\sf D}_{\alpha_{i}}\mathcal{L}(\omega_{i})$ are exceptional for $i=1,2$ and so are the bundles ${\sf D}_{\alpha_{i}}(\Psi_{1}^{\omega_{i}})$, their left mutations through the exceptional bundle $\mathcal{L}(0)$. Proposition 14.22 allows to identify the bundles ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ and ${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})$; it is sufficient to compute one of these, say ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ that is the kernel of the map $\Psi_{1}^{\omega_{1}}\rightarrow\mathcal{L}(\omega_{1}-\alpha_{1})$. The weights of $\Psi_{1}^{\omega_{1}}$ are $\omega_{1}-\alpha_{1}$ and $-\omega_{2}$, thus the kernel ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ is isomorphic to $\mathcal{L}(-\omega_{2})$. Similarly, ${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})=\mathcal{L}(-\omega_{1})$. This implies $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}),{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}),{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))=0$. Thus, ${\sf A}_{-2}$ is an exceptional block. Finally, Proposition 14.27 implies that the minimal full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ that contains all ${\sf A}_{-i},i=0,\dots,3$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$. ∎
Lemma 14.31.
Consider the following collection:
[…]
$\mathcal{L}(0)$
${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)$
$\mathcal{L}(\rho)$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,3$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$. In other words, the above collection is obtained from the collection (76) by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection (77) is left dual to (76).
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-3],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{1}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{2}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{2}}))={\sf k}$; $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{1}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
Remark 14.32.
The Steinberg weights for ${\bf A}_{2}$ are given by (see Section 3.2)
[…]
these coincide with the weights of the decomposition factors of the collection (76).
14.11 The quiver for ${\bf A}_{2}$
The picture below depicts the path quiver algebra of the above full exceptional collection. There are only zero-degree $\mathop{\mathrm{Hom}}\nolimits$-spaces in the non-trivial direction in the collection (76). The underlying graph of the quiver is the Bruhat graph for $SL_{3}/{B}$; the quiver path algebra is generated in degree one and the relations are quadratic that one reads off the $\mathop{\mathrm{Hom}}\nolimits$–spaces next to the arrows of the quiver. The labels assume $p$ is not too small. For small $p$ the characters are still correct and the $\mathop{\mathrm{Hom}}\nolimits$–spaces have a good filtration.
${{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\Psi_{1}^{\omega_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$${\mathcal{L}(-\rho)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\quad}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{1}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{1}+\rho})}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{2}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{2}+\rho})}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\quad\mathcal{L}(0)}$${{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$$\scriptstyle{\nabla_{\omega_{1}}}$${\Psi_{1}^{\omega_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$ (78)
The Weyl group elements that we associate to the nodes are as follows: The leftmost node belongs with the longest element $w_{0}$, the rightmost with the identity element. In the top row we have from left to right $s_{2}s_{1}$, $s_{2}$, and in the bottom row from left to right $s_{1}s_{2}$, $s_{1}$.
We list some of the arrows:
  • (1)
    $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(-\rho),\Psi_{1}^{\omega_{1}})={\sf Ker}(\nabla_{\omega_{1}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{1}+\rho})$. Note that $\nabla$-filtration on both ${\sf Ker}(\nabla_{\omega_{1}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{1}+\rho})$ and on $\nabla_{\omega_{2}}\otimes\nabla_{\omega_{2}}$ coincide and the associated graded is isomorphic to $\nabla_{\omega_{1}}\oplus\nabla_{2\omega_{2}}$.
  • (2)
    $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(-\rho),\Psi_{1}^{\omega_{2}})={\sf Ker}(\nabla_{\omega_{2}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{2}+\rho})$. Note that $\nabla$-filtration on both ${\sf Ker}(\nabla_{\omega_{2}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{2}+\rho})$ and on $\nabla_{\omega_{1}}\otimes\nabla_{\omega_{1}}$ coincide and the associated graded is isomorphic to $\nabla_{\omega_{2}}\oplus\nabla_{2\omega_{1}}$.
  • (3)
    $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(-\omega_{2}),\Psi_{1}^{\omega_{1}})={\sf k}$
  • (4)
    $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(-\omega_{1}),\Psi_{1}^{\omega_{2}})={\sf k}$
The $\nabla$-filtrations on some tensor products are given by:
  • (1)
    $\nabla_{\omega_{1}}\otimes\nabla_{\omega_{1}}$ is filtered by $\nabla_{2\omega_{1}}$ and $\nabla_{\omega_{2}}$;
  • (2)
    $\nabla_{\omega_{2}}\otimes\nabla_{\omega_{2}}$ is filtered by $\nabla_{\omega_{1}}$ and $\nabla_{2\omega_{2}}$
  • (3)
    $\nabla_{\omega_{1}}\otimes\nabla_{2\omega_{1}}$ is filtered by $\nabla_{3\omega_{1}}$ and $\nabla_{\rho}$;
  • (4)
    $\nabla_{\omega_{2}}\otimes\nabla_{2\omega_{2}}$ is filtered by $\nabla_{3\omega_{2}}$ and $\nabla_{\rho}$;
  • (5)
    $\nabla_{\omega_{1}}\otimes\nabla_{\omega_{2}}$ is filtered by $\nabla_{\rho}$ and ${\sf k}$;
  • (6)
    $\nabla_{\omega_{1}}\otimes\nabla_{\rho}$ is filtered by $\nabla_{\omega_{1}+\rho},\nabla_{\omega_{1}}$ and $\nabla_{2\omega_{2}}$.
14.12 Type ${\bf B}_{2}$
We assume here that the prime $2$ is invertible in ${\sf k}$. Let $G$ be a simple group of type ${B}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${B}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${P}_{\alpha}$ (resp., ${P}_{\beta}$) be the minimal parabolic generated by ${B}$ and $U_{\alpha}$ (resp., ${B}$ and $U_{\beta}$). The Grassmannian ${G}/{P}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=5$. The Grassmannian ${G}/{P}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=4$ (it is in fact isomorphic to $\mathbb{P}^{3}=\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast}))$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (56).
Lemma 14.33.
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/{B})$:
[…]
$\mathcal{L}(-\rho)$
${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$
${\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})$
${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})$
${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})$
$\Psi_{1}^{\omega_{\alpha}}$
$\Psi_{1}^{\omega_{\beta}}$
$\mathcal{L}(0)$
That is, ${\sf A}_{i}\subset{{}^{\perp}{\sf A}_{j}}$ for $-4\leq j<i\leq 0$, each of the vector bundles is exceptional, the vector bundles in a box are completely orthogonal, and the triangulated hull of $\langle{\sf A}_{-4},\dots,{\sf A}_{0}\rangle$ is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$.
Remark 14.34.
The $G$-equivariant vector bundles in the above collection can be explicitly identified:
  • [(1)]
  • (1)
    ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{U}_{2}$
  • (2)
    ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})=\Psi_{2}^{\omega_{\alpha}}$
  • (3)
    ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{L}(-\omega_{\alpha})$
  • (4)
    ${\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})=\mathcal{L}(-\omega_{\beta})$.
Here the rank two bundle $\mathcal{U}_{2}$ fits into a short exact sequence
$0\rightarrow\mathcal{L}(-\omega_{\alpha})\rightarrow\mathcal{U}_{2}\rightarrow\mathcal{L}(\omega_{\alpha}-\omega_{\beta})\rightarrow 0;$ (80)
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{G/B}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(G/B,\mathcal{L}(-\beta))=k$. The bundle $\Psi_{2}^{\omega_{\alpha}}$ is the left mutation of line bundle $\mathcal{L}(2\omega_{\alpha})$ through the admissible subcategory $\langle\mathcal{L}(0),\mathcal{L}(\omega_{\alpha})\rangle\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$ and fits into a short exact sequence
$0\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow\Lambda^{2}\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\alpha})\rightarrow 0$ (81)
The isomorphism (1) follows from Proposition 14.22. Let us show how (2) is obtained as Proposition 14.25 doesn’t directly lead to conclusion that the bundle ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})$ is exceptional in the non simply-laced case which is ${\bf B}_{2}$. To this end, observe that the collection of bundles
$\mathcal{L}(-\omega_{\alpha}),\Psi_{2}^{\omega_{\alpha}},\Psi_{1}^{\omega_{\alpha}},\mathcal{L}(0);$ (82)
is a full exceptional collection on $G/P_{\beta}$. To ensure that ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}=\Psi_{2}^{\omega_{\alpha}}$ it is sufficient to check that $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(0),{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$, $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$, and the ranks of bundles in Remark 14.34, (2) are the same. That the ranks are the same is immediate from Proposition 14.21. As for the orthogonality, the first equality holds by Proposition 14.18, the second one by Proposition 14.23, and the third one is a consequence of $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$. The latter equality is obtained by applying $\mathop{\mathrm{Hom}}\nolimits_{X}(-,\mathcal{L}(-\omega_{\alpha}))$ to the sequence (64) and then using Proposition 14.18, the exact sequence (47), and Theorem 14.1: all weights that are needed to compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\mathcal{L}(\omega_{\beta})),\mathcal{L}(-\omega_{\alpha}))$ are singular.
Proof.
Theorem 14.1 and Proposition 14.18 give ${\sf A}_{0}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=1,\dots,4$. Proposition 14.23 and Proposition 14.10 give ${\sf A}_{-1}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=2,\dots,4$. By Proposition 14.10 ${\sf A}_{-i}\subset{{}^{\perp}{\sf A}_{-4}}$ for $i=2,3$. Showing ${\sf A}_{-2}\subset{{}^{\perp}{\sf A}_{-3}}$ amounts to the vanishing
$\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0,\ \mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$ (83)
$\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0,\ \mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$ (84)
By Proposition 14.8, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(-\alpha)[1])$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(-\beta)[1])$. It follows from Remark 14.34 that ${\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=0$ and ${\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\beta}})=0$ as $\langle-\omega_{\alpha},\alpha^{\vee}\rangle=\langle-\omega_{\beta},\beta^{\vee}\rangle=-1$. Hence, the groups from (83) are trivial.
By Proposition 14.7, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))$. Let us show that both groups are trivial; it proves at the same time that ${\sf A}_{-2}$ is a block. By Proposition 14.8, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(-\beta)[1])$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))\otimes\mathcal{L}(-\alpha)[1])$. By Remark 14.34, ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$ and ${\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))$ are identified with line bundles $\mathcal{L}(-\omega_{\alpha})$ and $\mathcal{L}(-\omega_{\beta})$, respectively. Thus, we need to compute the groups $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}-\beta)[1])$ and $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},\mathcal{L}(-\omega_{\beta}-\alpha)[1])$. Using the sequence (56), this amounts to computing cohomology of line bundles; one easily checks that all the weights occurring are singular and thus all the groups vanish by Theorem 14.1. This proves the vanishing of groups in (84) and ${\sf A}_{-2}$ being a block.
Block structure in blocks ${\sf A}_{-1},{\sf A}_{-3}$ follows easily again from Theorem 14.1. Note that by [8], the collection of bundles $\langle\mathcal{L}(0),\Psi_{1}^{\omega_{\beta}},{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\rangle$ generates the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P_{\alpha})$ and the collection of bundles $\langle\mathcal{L}(0),\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\rangle$ generates the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P_{\beta})$. By Proposition 14.27, the collection of subcategories ${\sf A}_{-i},i=0,\dots 4$ is a semiorthogonal decomposition. Finally, all the line bundles in (79) are exceptional by Corollary 14.4, while ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})$ is exceptional by Remark 14.34, (2) and ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})$ is exceptional by Corollary 14.26. ∎
Lemma 14.35.
Consider the following collection:
[…]
$\mathcal{L}(0)$
${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)$
$\mathcal{L}(\rho)$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,4$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\bf Sp}_{4}/{B})$. In other words, the above collection is obtained from the collection (79) by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection (85) is right dual to (79).
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-4],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  8. 7.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  9. 8.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
2)2)2 ): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$. We have
$\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\rho)[3]).$ (86)
By Remark 14.34, (3), the latter group is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-\omega_{\beta})\otimes(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\rho)[3])$. Tensoring the sequence (57) with $\mathcal{L}(-\omega_{\beta}-\rho)$ and using Theorem 14.1, wee see from the long cohomology sequence that the required isomorphism follows from $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))={\sf k}[-4]$.
[…]
4)4)4 ): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$. By Remark 14.34, (1) and (2), ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{U}_{2}$ and ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})=\Psi_{2}^{\omega_{\alpha}}$. The bundle $\Psi_{2}^{\omega_{\alpha}}$ is the cokernel in the short exact sequence
$0\rightarrow\mathcal{L}(-2\omega_{\alpha})\rightarrow\nabla_{\omega_{\alpha}}^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow 0;$ (87)
applying $\mathop{\mathrm{Hom}}\nolimits_{X}(\mathcal{U}_{2}\otimes\mathcal{L}(-\rho),-)$ to this sequence and using Theorem 14.1, we see that the only non-vanishing cohomology group that appears in the long cohomology sequences is $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-4\omega_{\alpha}))={\sf k}[-3]$ (note that this isomorphism holds for $p=2$, despite the restriction on $p$ in Theorem 14.1: there is an isomorphism $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-4\omega_{\alpha}))=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{\beta},\mathcal{L}(-4\omega_{\alpha}))$ and $Y_{\beta}$ is isomorphic to ${\mathbb{P}}(\nabla_{\omega_{\alpha}})$, while the sheaf $\mathcal{L}(-4\omega_{\alpha})$ is the canonical sheaf on $Y_{\beta}$. Therefore, the required isomorphism holds by Serre duality on $Y_{\beta}$). The coboundary homomorphism makes it isomorphic to the group $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}[-2]$. This cancels out with the shift by $[2]$ giving the required isomorphism. ∎
Remark 14.36.
The Steinberg weights for ${\bf B}_{2}$ are given by (see Section 3.2)
[…]
these coincide with the weights of the decomposition factors of the collection (79).
14.13 The quiver for ${\bf B}_{2}$
[…]
${{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\Lambda^{2}\nabla_{\omega_{\alpha}}=\nabla_{\omega_{\beta}}\oplus{\sf k}}$${{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\Lambda^{2}\nabla_{\omega_{\alpha}}=\nabla_{\omega_{\beta}}\oplus{\sf k}}$${\Psi_{1}^{\omega_{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$${\mathcal{L}(-\rho)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\quad}$$\scriptstyle{\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\beta}}=\nabla_{\rho}\oplus\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{2\omega_{\alpha}}\oplus\nabla_{\omega_{\beta}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\beta}}}$${\quad\mathcal{L}(0)}$${{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{2\omega_{\alpha}}\oplus{\sf k}}$${{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$${\Psi_{1}^{\omega_{\beta}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\beta}}}$
The Weyl group elements that we associate to the nodes are as follows: The leftmost node belongs with the longest element $w_{0}$, the rightmost with the identity element. In the top row we have from left to right $s_{\alpha}s_{\beta}s_{\alpha}$, $s_{\beta}s_{\alpha}$, $s_{\alpha}$, and in the bottom row from left to right $s_{\beta}s_{\alpha}s_{\beta}$, $s_{\alpha}s_{\beta}$, $s_{\beta}$.
The $\nabla$-filtrations on the tensor products of fundamental representations are given by:
  1. 1.
    $\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\alpha}}$ is filtered by $\nabla_{2\omega_{\alpha}},\nabla_{\omega_{\beta}}$, and ${\sf k}$.
  2. 2.
    $\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\beta}}$ is filtered by $\nabla_{\rho}$ and $\nabla_{\omega_{\alpha}}$.
14.14 Type ${\bf G}_{2}$
Let $p\geq 5$ be a prime. Let $G$ be a simple group of type ${G}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${G}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${P}_{\alpha}$ (resp., ${P}_{\beta}$) be the minimal parabolic generated by ${B}$ and ${\bf U}_{\alpha}$ (resp., ${B}$ and ${\bf U}_{\beta}$). The Grassmannian ${G}/{P}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=7$. The Grassmannian ${G}/{P}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=14$. As usual, denote $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ (resp., $\pi_{\beta}:{G}/{B}\rightarrow{G}/{P}_{\beta}$) the projection. Associated to $\pi_{\alpha}$ and $\pi_{\beta}$ are Demazure operators ${\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\pi_{\alpha}}_{\ast}$ and ${\sf D}_{\beta}:=\pi_{\beta}^{\ast}{\pi_{\beta}}_{\ast}$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (56).
Lemma 14.37.
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({G}_{2}/{B})$:
[…]
$\mathcal{L}(-\rho)$
${\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$
${\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}(-\omega_{\alpha}))$
$(\Psi_{1}^{\omega_{\alpha}})^{\ast}(-\omega_{\alpha})$
${\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$
${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$
${\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})$
$\Psi_{2}^{\omega_{\alpha}}$
${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})$
$\Psi_{1}^{\omega_{\alpha}}$
$\Psi_{1}^{\omega_{\beta}}$
$\mathcal{L}(0)$
That is, ${\sf A}_{i}\subset{{}^{\perp}{\sf A}_{j}}$ for $-6\leq j<i\leq 0$, each of the vector bundles is exceptional, the vector bundles in a box are completely orthogonal, and the triangulated hull of $\langle{\sf A}_{-6},\dots,{\sf A}_{0}\rangle$ is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$.
Here $\Psi_{2}^{\omega_{\alpha}}$ is the left mutation of $\mathcal{L}(2\omega_{\alpha})$ through the admissible subcategory $\langle\mathcal{O},\mathcal{L}(\omega_{\alpha})\rangle$ and $=(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})$ is the right mutation of $\mathcal{L}(-2\omega_{\alpha})$ through $\mathcal{L}(2\omega_{\alpha})$. Specifically, $\Psi_{2}^{\omega_{\alpha}}$ is defined by the following exact sequence:
$0\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{G}_{2}/{B}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\alpha})\rightarrow\mathcal{L}(2\omega_{\alpha})\rightarrow 0$ (89)
The bundle $(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})$ is defined by the following short exact sequence:
$0\rightarrow\mathcal{L}(-2\omega_{\alpha})\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(-\omega_{\alpha})\rightarrow(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})\rightarrow 0$ (90)
The above sequence is obtained by dualizing (57) for $\omega_{\alpha}$ and tensoring the result with $\mathcal{L}(-\omega_{\alpha})$ (note that $\nabla_{\omega_{\alpha}}$ is self-dual).
Remark 14.38.
The case of ${G}_{2}$ differs from other types in rank two in that the bundle $\Psi_{2}^{\omega_{\alpha}}$ appears in the second block of the exceptional collection for ${G}_{2}$ in Lemma 14.37 (as opposed to ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ suggested by the cases of $A_{2}$ and ${B}_{2}$). The reason is that the bundle ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional (one of the two deviations in Proposition 14.25): ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is the kernel of the surjective map $\nabla_{\omega_{\beta}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ and the latter vector bundle ${\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ is not exceptional. In turn, ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional either.
Remark 14.39.
The $G$-equivariant vector bundles in the above collection can be explicitly identified:
  • [(1)]
  • (1)
    ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{U}_{2}^{\perp}$ which is the kernel of canonical surjective morphism $\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\mathcal{U}_{2}^{\ast}$; the bundle $\mathcal{U}_{2}$ is defined by the short exact sequence (91) below.
  • (2)
    ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})={\sf S}$ where ${\sf S}$ is a rank 4 vector bundle called the spinor bundle defined by the short exact sequence (92) below.
  • (3)
    ${\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{U}_{2}$ which is defined by the short exact sequence (91) below.
  • (4)
    ${\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{L}(-\omega_{\alpha})$.
  • (5)
    ${\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=\mathcal{E}$ is a rank 11 vector bundle defined by the short exact sequence (93) below.
  • (6)
    ${\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))=\mathcal{L}(-\omega_{\beta})$, by Remark 14.16, (62).
Let us give concrete descriptions of the high rank $G$-equivariant vector bundles in the above list. The rank two bundle $\mathcal{U}_{2}$ fits into a short exact sequence
$0\rightarrow\mathcal{L}(-\omega_{\alpha})\rightarrow\mathcal{U}_{2}\rightarrow\mathcal{L}(\omega_{\alpha}-\omega_{\beta})\rightarrow 0;$ (91)
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{G/B}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(G/B,\mathcal{L}(-\beta))=k$.
The spinor bundle $\sf S$ fits into the short exact sequence:
$0\rightarrow\mathcal{U}_{2}\rightarrow{\sf S}\rightarrow\mathcal{U}_{2}^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})\rightarrow 0$ (92)
Applying $\sf D_{\alpha}$ to (89), and taking into account the isomorphisms $\sf D_{\alpha}(\mathcal{L}(\omega_{\alpha}))=\mathcal{U}_{2}^{\ast},\sf D_{\alpha}(\mathcal{L}(2\omega_{\alpha}))={\sf S}^{2}\mathcal{U}_{2}^{\ast}$, we obtain:
$0\rightarrow\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{G}_{2}/{B}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{U}_{2}^{\ast}\rightarrow{\sf S}^{2}\mathcal{U}_{2}^{\ast}\rightarrow 0$ (93)
Now $\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=:\mathcal{E}$ is a vector bundle on $G_{2}/B$ since ${\rm R}^{1}{\pi_{\alpha}}_{\ast}\Psi_{2}^{\omega_{\alpha}}=0$; the latter fact follows from the short sequence
$0\rightarrow\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{G/P_{\beta}}\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow\mathcal{L}(0)\rightarrow 0,$ (94)
and from the splitting type of $\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{G/P_{\beta}}$ to a line $l$ on $G/P_{\beta}$ (cf. [45, Section 4]) as $(\underbrace{-1,\dots,-1,}_{5}\underbrace{0,\dots,0}_{10})$. It also follows from the above that $\mathcal{E}$ fits into the short exact sequence
$0\rightarrow\Lambda^{2}\mathcal{U}_{2}^{\perp}\rightarrow\mathcal{E}\rightarrow\mathcal{L}(0)\rightarrow 0;$ (95)
These exact sequences will be of help in cohomological computations related to the collection of Lemma 14.37.
Lemma 14.40.
Consider the following collection:
[…]
$\mathcal{L}(0)$
${\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}(-\omega_{\alpha}))\otimes\mathcal{L}(\rho)$
$(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{\beta})$
${\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
$\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)$
${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)$
$\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)$
$\mathcal{L}(\rho)$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,6$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G_{2}/{B})$. In other words, the above collection is obtained from the collection Lemma 14.37 by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection above is right dual to that of Lemma 14.37.
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-6],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-6],{{\sf A}_{i}})=0$ for $i\neq-6$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{{\sf A}_{i}})=0$ for $i\neq-5$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{{\sf A}_{i}})=0$ for $i\neq-5$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  8. 7.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  9. 8.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],\Psi_{2}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  10. 9.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],\Psi_{2}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  11. 10.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  12. 11.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],,\Psi_{1}^{\omega_{\alpha}})={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],,{{\sf A}_{i}})=0$ for $i\neq-1$
  13. 12.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
2): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))={\sf k}$. By Remark 14.39, (6), the group in question is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\beta}-\rho)[5])$; tensoring the sequence (57) for $\omega_{\alpha}$ with $\mathcal{L}(-\omega_{\beta}-\rho)$, the coboundary homomorphism of the associated long exact cohomology sequence gives the required isomorphism as $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))={\sf k}[-6]$.
3): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, (4), the group in question is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,(\Psi_{1}^{\omega_{\beta}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}-\rho)[5])$ and we obtain the required isomorphism as in the previous item 2).
4): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, (3), the group in question is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}},\mathcal{U}_{2}\otimes\mathcal{L}(-\rho)[4])$. Applying $\mathop{\mathrm{Hom}}\nolimits_{X}(-,\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))$ to (89) and using isomorphisms $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=0$, $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{\alpha}),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=0$, and $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(2\omega_{\alpha}),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))=\sf k[-6]$ (use (91) for the last isomorphism), we obtain the required isomorphism.
5): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))={\sf k}$. By Serre duality, the group in question is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}),{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\rho)[2])^{\ast}=\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast},{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\omega_{\beta})[2])^{\ast}$. Now using the sequences (57) and Proposition 14.22, one calculates the group $\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast},{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\omega_{\beta}))$ to be isomorphic to $\sf k[-2]$; this gives the required isomorphism.
6): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=\sf S$ and ${\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=\mathcal{E}$. To compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})$, use the sequence (95) from which one obtains $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\Lambda^{2}\mathcal{U}_{2}^{\perp})$ as $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{L}(0)=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}^{\ast}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{\beta}{\sf S}^{\ast}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{\beta},{\pi_{\beta}}_{\ast}{\sf S}^{\ast}\otimes{\pi_{\beta}}_{\ast}\mathcal{L}(-\rho))=0$. Now $\Lambda^{2}\mathcal{U}_{2}^{\perp}$ has a resolution which is the second exterior power of the tautological short exact sequence $0\rightarrow\mathcal{U}_{2}^{\ast}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\mathcal{U}_{2}^{\perp}\rightarrow 0$:
$0\rightarrow{\sf S}^{2}\mathcal{U}_{2}^{\ast}\rightarrow\mathcal{U}_{2}^{\ast}\otimes\nabla_{\omega_{\alpha}}\rightarrow\Lambda^{2}\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\Lambda^{2}\mathcal{U}_{2}^{\perp}\rightarrow 0$ (96)
Apply $\mathop{\mathrm{Hom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],-)$ to the above sequence. We have already obtained $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{L}(0))=0$. A straightforward calculation with the help of the dual to sequence (91) and Theorem 14.1 shows that $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{U}_{2}^{\ast})=0$. We have to compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],{\sf S}^{2}\mathcal{U}_{2}^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}^{2}\mathcal{U}_{2}^{\ast}\otimes{\sf S}^{\ast}\otimes\mathcal{L}(-\rho)[3])$. Now the weights of the $G$-equivariant bundle ${\sf S}^{2}\mathcal{U}_{2}^{\ast}$ are equal to $2\omega_{\beta}-2\omega_{\alpha},\omega_{\beta},2\omega_{\alpha}$ and using also the isomorphism ${\sf S}={\sf S}^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})$, we obtain three cohomology groups (ignoring the shift by $[3]$):
$\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(\omega_{\beta}-2\omega_{\alpha})),\ \mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}),\ \mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(2\omega_{\alpha}-\omega_{\beta}))$
Now both groups $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S})$ and $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(2\omega_{\alpha}-\omega_{\beta}))$ are trivial. Calculating the weights of the remaining $G$-equivariant bundle ${\sf S}\otimes\mathcal{L}(\omega_{\beta}-2\omega_{\alpha})$, we see that the only weight having a non-trivial cohomology is equal to $\omega_{\beta}-2\omega_{\alpha}=-\alpha$ and $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-\alpha))=\sf k[-1]$. Coming back to the sequence (96) and remembering the shift by $[3]$, we obtain the isomorphism $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\Lambda^{2}\mathcal{U}_{2}^{\perp})=\sf k$ and hence the required isomorphism in 6). ∎
Remark 14.41.
The Steinberg weights for ${\bf G}_{2}$ are given by (see Section 3.2)
[…]
the weights of the decomposition factors of the collection of Lemma 14.37 are the Steinberg weights and the dominant weight $\omega_{\alpha}$.
14.15 The quiver for ${\bf G}_{2}$
[…]
${{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$${(\Psi_{1}^{\omega_{\alpha}})^{\ast}(-\omega_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\alpha}}\oplus\nabla_{\omega_{\alpha}}\oplus\sf k}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$${{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$${\Psi_{2}^{\omega_{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\Lambda^{2}\nabla_{\omega_{\alpha}}\oplus{\sf k}}$${\Psi_{1}^{\omega_{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$${\mathcal{L}(-\rho)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\quad}$$\scriptstyle{\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\beta}}}$$\scriptstyle{\nabla_{2\omega_{\alpha}}\oplus\nabla_{\omega_{\beta}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\beta}}}$${\quad\mathcal{L}(0)}$${{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}(-\omega_{\alpha}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\alpha}}}$${{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$${{\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}\oplus\sf k}$$\scriptstyle{\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\alpha}}}$${{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$$\scriptstyle{\nabla_{\omega_{\alpha}}}$${\Psi_{1}^{\omega_{\beta}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{\beta}}}$
The $\nabla$-filtrations on the tensor products of fundamental representations are given by:
  • (1)
    $\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\beta}}$ is filtered by $\nabla_{\rho},\nabla_{2\omega_{\alpha}}$, and $\nabla_{\omega_{\alpha}}$.
  • (2)
    $\nabla_{\omega_{\alpha}}\otimes\nabla_{\omega_{\alpha}}$ is filtered by $\nabla_{2\omega_{\alpha}},\nabla_{\omega_{\beta}},\nabla_{\omega_{\alpha}}$, and ${\sf k}$.
The Weyl group elements that we associate to the nodes are as follows: The leftmost node belongs with the longest element $w_{0}$, the rightmost with the identity element. In the top row we have from left to right $s_{\alpha}s_{\beta}s_{\alpha}s_{\beta}s_{\alpha}$, $s_{\beta}s_{\alpha}s_{\beta}s_{\alpha}$, $s_{\alpha}s_{\beta}s_{\alpha}$, $s_{\beta}s_{\alpha}$, $s_{\alpha}$, and in the bottom row from left to right $s_{\beta}s_{\alpha}s_{\beta}s_{\alpha}s_{\beta}$, $s_{\alpha}s_{\beta}s_{\alpha}s_{\beta}$, $s_{\beta}s_{\alpha}s_{\beta}$, $s_{\alpha}s_{\beta}$, $s_{\beta}$.
14.16 Type ${\bf G}_{2}$ over ${\mathbb{Z}}$
The non-vanishing statements of Lemmas 14.31, 14.35, and 14.40 boiled down to non-vanishing cohomology groups of line bundles either in degree zero, or in the top degree, or in degree one. By [3, Corollary 2.6] and [29, II, Section 8.8, (6)], those cohomology groups are free $\mathbb{Z}$-modules considered as cohomology groups of line bundles over $\mathbb{G}/\mathbb{B}$. Thus, the results of the above lemmas hold over ${\mathbb{Z}}$.
14.16.1 Type ${\bf G}_{2}$ over ${\mathbb{Z}}$, without proofs
Let $\alpha$ be the short simple root, $\beta$ the long one. Put $a=s_{\alpha}$, $b=s_{\beta}$. The exceptional $\mathbb{B}$-modules $X_{v}$ have no ${\mathbb{Z}}$ torsion. Over ${\mathbb{Z}}[1/2]$ each $X_{v}$ has a relative Schubert filtration and over ${\mathbb{Z}}[1/6]$ each $X_{v}^{*}$ has an excellent filtration. If $v,w\in W$ are distinct and $\ell(v)\leq\ell(w)$, then $\mathop{\mathrm{RHom}}\nolimits^{\bullet}(\mathcal{L}(X_{v}),\mathcal{L}(X_{w}))=0$.
The natural map $\phi_{w}:P(-e_{w})^{*}\to X_{w}$ is injective for $w\in W$.
Some properties of ${\mathsf{coker}}(\phi_{w})$:
${\mathsf{coker}}(\phi_{\mathrm{id}})=0$.
${\mathsf{coker}}(\phi_{a})\cong{\mathbb{Z}}/2{\mathbb{Z}}$.
${\mathsf{coker}}(\phi_{b})\cong\nabla(\omega_{\alpha})/3\nabla(\omega_{\alpha})$.
$0\to\Delta(\omega_{\alpha})\to{\mathsf{coker}}(\phi_{ab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{ba})\cong{\mathbb{Z}}/2{\mathbb{Z}}$.
$0\to P(-e_{ba})^{*}\to{\mathsf{coker}}(\phi_{bab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{aba})\cong{\mathbb{Z}}$.
$0\to P(-e_{aba})^{*}\to{\mathsf{coker}}(\phi_{abab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{baba})={\mathsf{coker}}(\phi_{babab})={\mathsf{coker}}(\phi_{ababa})={\mathsf{coker}}(\phi_{bababa})=0$.